### v1

# The Riemannian geometry of Sinkhorn divergences

Abstract
We propose a new metric between probability measures on a compact metric space that mirrors the Riemannian manifold-like structure of quadratic optimal transport but includes entropic regularization. Its metric tensor is given by the Hessian of the Sinkhorn divergence, a debiased variant of entropic optimal transport. We precisely identify the tangent space it induces, which turns out to be related to a Reproducing Kernel Hilbert Space (RKHS). As usual in Riemannian geometry, the distance is built by looking for shortest paths. We prove that our distance is geodesic, metrizes the weak-$\ast$ topology, and is equivalent to a RKHS norm. Still it retains the geometric flavor of optimal transport: as a paradigmatic example, translations are geodesics for the quadratic cost on $\mathbb{R}^{d}$. We also show two negative results on the Sinkhorn divergence that may be of independent interest: that it is not jointly convex, and that its square root is not a distance because it fails to satisfy the triangle inequality.
1 Introduction
Optimal transport is a mathematical optimization problem which allows to lift a metric $d$ on a space $X$ to the space $\mathcal{P}(X)$ of probability measures over $X$, interpreted as the average distance that infinitesimal units of mass have to travel when efficiently transforming one measure into the other. When the base space $X$ is a Riemannian manifold, then the space of probability measures can be formally interpreted as a Riemannian manifold as well.
In recent years there was an increasing interest in entropic regularization of optimal transport, which adds some stochasticity to the coupling of probability measures. This variant is easier to compute numerically, has better statistical properties, and is more regular as a function of its input measures and ground cost. Its debiased version defines a smooth positive definite loss function on $\mathcal{P}(X)$. However, it does not define a distance, or even a Riemannian geometry, on $\mathcal{P}(X)$.
In this work we aim at combining both worlds: having a Riemannian metric on $\mathcal{P}(X)$ which lifts the geometry of $X$ in a similar way to what optimal transport does, while retaining the smoothness of the Sinkhorn divergence.
To that end, we consider the second-order expansion of the Sinkhorn divergence on the diagonal: this yields a positive definite quadratic form on a meaningful tangent space to the space of probability measures, in other words a Riemannian metric tensor. This metric tensor induces a Riemannian distance on $\mathcal{P}(X)$ via minimization over lengths of curves, which is aware of the geometry of $X$ while being “smoother” than the optimal transport geometry. Investigation of the additional properties of our distance, in particular regarding computational and statistical aspects, is postponed to future works. In the rest of this introduction we recall what is meant by the “geometry of optimal transport” before summarizing the main results of the present work. We emphasize that geodesics for our distance are different from Schrödinger bridges: a more detailed discussion is given on page 1.
Optimal transport and its geometry.
On a (compact) space $X$, the value of the optimal transport problem with ground cost $c:X\times X\to[0,+\infty)$ between two probability measures $\mu,\nu\in\mathcal{P}(X)$ will be denoted by $\mathrm{OT}_{0}(\mu,\nu)$. It is defined as [62, 57, 2]
$\mathrm{OT}_{0}(\mu,\nu)\coloneqq\inf_{\pi\in\Pi(\mu,\nu)}\iint_{X\times X}c(x,y)\,\mathrm{d}\pi(x,y),$ (1)
where $\Pi(\mu,\nu)=\left\{\pi\in\mathcal{P}(X\times X)\,\middle|\,(\operatorname{P}_{1})_{\#}\pi=\mu,~{}(\operatorname{P}_{2})_{\#}\pi=\nu\right\}$ is the set of couplings having marginals $\mu,\nu$ and $(P_{i})_{\#}\pi$ is the push-forward of the measure $\pi$ by the projection $P_{i}$ on the $i$-th factor. If $c=d^{2}$ is the square of a distance $d$ on $X$, then $(\mu,\nu)\mapsto\sqrt{\mathrm{OT}_{0}(\mu,\nu)}$ defines a distance on $\mathcal{P}(X)$, the so-called (quadratic) Wasserstein or Monge-Kantorovich distance [2, Sec. 7].
To explain the geometry it induces, we restrict to $X$ being a convex subset of a Euclidean space $\mathbb{R}^{d}$, though everything could be extended to Riemannian manifolds. Historically this geometric structure was introduced in [51, Sec. 4] as the quotient of the space of maps by measure preserving maps with applications to dissipative evolution equations, and can be traced back to hydrodynamics [3, 10]. In this presentation we will rather focus on the “metric tensor” that optimal transport defines. If $(\mu_{t})_{t}$ is a curve of probability measures obtained by following the flow of a potential velocity field $v=\nabla\psi$, that is, $\mu_{t}=\Psi(t,\cdot)_{\#}\mu_{0}$ is the pushforward of $\mu_{0}$ by the flow $\Psi(t,\cdot)$ of the ODE $\dot{x}_{t}=\nabla\psi(x_{t})$, then by [2, Prop. 8.5.6]
$\lim_{t\to 0}\frac{\mathrm{OT}_{0}(\mu_{0},\mu_{t})}{t^{2}}=\int_{X}|\nabla\psi(x)|^{2}\,\mathrm{d}\mu_{0}(x).$ (2)
That is, on small distances, the squared Wasserstein distance is quadratic in $v=\nabla\psi$ which parametrizes how the curve moves on $\mathcal{P}(X)$. As the curve $(\mu_{t})_{t}$ satisfies the continuity equation $\dot{\mu}_{t}=-\operatorname{div}(\mu_{t}\nabla\psi)$, we can think of the first-order distribution $\dot{\mu}_{t}$ as the tangent vector whose squared length is given by the right-hand side of (2). That is, for $\mu\in\mathcal{P}(X)$ the formal metric tensor $\mathbf{g}^{0}$ is
$\mathbf{g}^{0}_{\mu}(\dot{\mu},\dot{\mu})=\int_{X}|\nabla\psi(x)|^{2}\,\mathrm{d}\mu(x),\qquad\dot{\mu}=-\operatorname{div}(\mu\nabla\psi).$ (3)
That $\mathcal{P}(X)$ has such a Riemannian structure means that the squared distance $\mathrm{OT}_{0}$ can be recovered by minimizing the energy of paths joining $\mu$ to $\nu$:
$\mathrm{OT}_{0}(\mu,\nu)=\inf_{(\mu_{t})_{t\in[0,1]}}\int_{0}^{1}\mathbf{g}^{0}_{\mu_{t}}(\dot{\mu}_{t},\dot{\mu}_{t})\,\mathrm{d}t$ (4)
where the infimum is taken over sufficiently regular curves $(\mu_{t})_{t\in[0,1]}$ of measures joining $\mu$ to $\nu$: this is the dynamical formulation of optimal transport, the so-called Benamou-Brenier formula [4][57, Sec. 6.1].
In summary: for $\mu\in\mathcal{P}(X)$ a “point” in the manifold, the tangent vectors at $\mu$ read $\dot{\mu}=-\operatorname{div}(\mu\nabla\psi)$ for $\psi:X\to\mathbb{R}$ , and $\mathbf{g}^{0}_{\mu}(\dot{\mu},\dot{\mu})$ defined in (3) is the squared length of the tangent vector $\dot{\mu}$. The proper tangent space is actually the completion of $\left\{-\operatorname{div}(\mu\nabla\psi)\,\middle|\,\psi\in\mathcal{C}^{\infty}(X)\right\}$ with respect to $\mathbf{g}^{0}_{\mu}$. It is canonically identified with vector fields $v\in L^{2}(X,\mu;\mathbb{R}^{d})$ which can be approximated in $L^{2}(X,\mu;\mathbb{R}^{d})$ by smooth gradients. Even though this does not define an infinite-dimensional Riemannian manifold in the sense of [38], a precise meaning can still be given thanks to the tools of analysis on non-smooth spaces [2, Sec. 8].
Once viewing the problem through this lens, we can import many concepts of Riemannian geometry into $\mathcal{P}(X)$. Geodesics are defined as the minimizers of the right-hand side of (4) and correspond to McCann’s displacement interpolation [45]. For two probability measures $\mu,\nu\in\mathcal{P}(X)$, the Riemannian logarithmic map $\log_{\mu}(\nu)$ is the initial velocity of the geodesic from $\mu$ to $\nu$. Conversely, given a tangent vector $\dot{\mu}=-\operatorname{div}(\mu\nabla\psi)$, the exponential map is $\exp_{\mu}(\dot{\mu})=(\operatorname{id}+\nabla\psi)_{\#}\mu$, but the extension to the whole tangent space and the injectivity radius can be tricky to characterize. With logarithmic and exponential maps, one can for instance define a linearized version of optimal transport [63, 58] and subsequently perform principal component analysis in tangent space [9]. Several extensions of optimal transport rely on its Riemannian structure, such as barycenters [1], gradient flows [2], harmonic maps [39], parallel transport [31], to cite a few examples. As the Riemannian interpretation is mostly formal, giving a rigorous meaning to these concepts usually requires some (hard) work of analysis on non-smooth spaces.
The Sinkhorn divergence.
For a regularization parameter $\varepsilon>0$, entropic optimal transport [22, 54] is given by the minimization problem
$\mathrm{OT}_{\varepsilon}(\mu,\nu)\coloneqq\inf_{\pi\in\Pi(\mu,\nu)}\left\{\iint_{X\times X}c\,\mathrm{d}\pi+\varepsilon\operatorname{KL}(\pi\,|\,\mu\otimes\nu)\right\},$ (5)
where $\operatorname{KL}$ is the Kullback–Leibler divergence
$\operatorname{KL}(\pi\,|\,\rho)\coloneqq\begin{cases}\displaystyle{\int_{X}h\left(\frac{\mathrm{d}\pi}{\mathrm{d}\rho}\right)\,\mathrm{d}\rho}&\textnormal{if }\pi\ll\rho,\,\rho\geq 0,\\ +\infty&\textnormal{else}\end{cases}$
with $h(s)\coloneqq s\log(s)-s+1$ for $s>0$, $h(0)\coloneqq 1$, $h(s)\coloneqq+\infty$ for $s<0$. As $\varepsilon\to 0$, $\mathrm{OT}_{\varepsilon}$ converges to $\mathrm{OT}_{0}$ [15, 40, 48]. Compared to the original transport problem it can be computed faster [54] and the statistical complexity suffers much less from the curse of dimensionality [28, 46]. Moreover, as a function of $\mu$ and $\nu$, the cost $\mathrm{OT}_{\varepsilon}$ is more regular and its gradients can be computed efficiently [29].
However, even if $c=d^{2}$ for $d$ a distance on $X$, the function $\sqrt{\mathrm{OT}_{\varepsilon}}$ is not a distance, as generically $\mathrm{OT}_{\varepsilon}(\mu,\mu)>0$ and $\operatorname*{argmin}_{\nu\in\mathcal{P}(X)}\mathrm{OT}_{\varepsilon}(\mu,\nu)\neq\mu$. For this reason [29] introduced the Sinkhorn divergence, or debiased entropic optimal transport:
$S_{\varepsilon}(\mu,\nu)\coloneqq\mathrm{OT}_{\varepsilon}(\mu,\nu)-\frac{1}{2}\mathrm{OT}_{\varepsilon}(\mu,\mu)-\frac{1}{2}\mathrm{OT}_{\varepsilon}(\nu,\nu).$ (6)
By construction it satisfies $S_{\varepsilon}(\mu,\mu)=0$, in fact more generally $S_{\varepsilon}(\mu,\nu)\geq 0$ with equality if and only if $\mu=\nu$, and it also converges to $\mathrm{OT}_{0}$ as $\varepsilon\to 0$. In [26] the authors prove the following theorem which justify its use in machine learning as a loss function:
Theorem 1.1 ([26, Thm. 1]).
When $k_{c}=\exp(-c/\varepsilon)$ is a positive definite, universal, jointly Lipschitz continuous kernel over a compact space $X$, then $S_{\varepsilon}$ defines a symmetric, positive definite and smooth loss function that is convex in each of its input variables. It also metrizes the convergence in law.
The assumption of positivity of the kernel is satisfied in particular for the quadratic cost on a Euclidean domain. Nevertheless, $\sqrt{S_{\varepsilon}}$ is not a distance, because it does not satisfy the triangle inequality. Although this is well known in the community, we have not been able to find an explicit statement of this property and consequently provide an elementary proof in Section 7.1 of this article.
Outline of the contributions.
As mentioned earlier we seek to build a Riemannian distance out of the loss function $S_{\varepsilon}$. We restrict to $X$ being a compact metric space with a continuous cost function $c\in\mathcal{C}(X\times X)$. Similarly to [26], we assume that it induces a positive definite universal kernel $k_{c}:=\exp(-c/\varepsilon)$ on $X\times X$ (see [47, 60]). Throughout this article we consider some fixed $\varepsilon>0$ and we will not consider any limit $\varepsilon\to 0$ or $\varepsilon\to+\infty$ here.
The Hessian of the Sinkhorn divergence. To construct a Riemannian structure based on the Sinkhorn divergence, we follow an approach from information geometry. For each point $\mu\in\mathcal{P}(X)$ we build a metric tensor, that is, a positive definite quadratic form $\mathbf{g}_{\mu}$ on some appropriate tangent space to $\mathcal{P}(X)$ at $\mu$, by considering the Hessian of $S_{\varepsilon}(\mu,\nu)$ on the diagonal $\mu=\nu$. Applying this strategy to the Kullback–Leibler divergence in place of the Sinkhorn divergence, as in the pioneering works of Hotelling [36] and Rao [55], leads to the Fisher information metric. In Theorem 3.3 we derive the formula for the Hessian of $S_{\varepsilon}$ for simple paths following a “vertical” perturbation $\mu_{t}=\mu+t\delta\mu$. Writing $(f_{\mu_{t},\mu_{s}},g_{\mu_{t},\mu_{s}})$ for the entropic dual variables between $\mu_{t}$ and $\mu_{s}$ (i.e. the entropic analogues of Kantorovich potentials, see Section 2), the formula reads
$\lim_{t\to 0}\frac{S_{\varepsilon}(\mu,\mu_{t})}{t^{2}}=\frac{1}{2}\left\langle\delta\mu,\frac{\partial f_{\mu,\mu_{s}}}{\partial s}\Big{|}_{s=0}\right\rangle=\frac{\varepsilon}{2}\left\langle\delta\mu,(\operatorname{id}-K_{\mu}^{2})^{-1}H_{\mu}\left[\delta\mu\right]\right\rangle.$ (7)
Here $K_{\mu}$ and $H_{\mu}$, introduced in Definition 3.2, are integral operators induced by the self- transport kernel
$k_{\mu}:=\exp\left(\frac{1}{\varepsilon}({f_{\mu,\mu}}\oplus{f_{\mu,\mu}}-c)\right).$
The result is not completely new and can be related to works studying the central limit theorem for the Sinkhorn divergence [34, 35], where the “perturbations” $\mu_{t}$ are estimations of the measure $\mu$ by i.i.d. samples.
The metric tensor and its canonical tangent space. We define the metric tensor $\mathbf{g}_{\mu}$ as the right-hand side of (7), that is, at least for a “balanced” signed measure $\dot{\mu}$ with $\langle\dot{\mu},\mathbb{1}_{X}\rangle=0$,
$\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})=\frac{\varepsilon}{2}\left\langle\dot{\mu},(\operatorname{id}-K_{\mu}^{2})^{-1}H_{\mu}\left[\dot{\mu}\right]\right\rangle.$
We then want to understand for which class of “tangent vectors” $\dot{\mu}$ this can be extended. This leads us to the most technical part of the paper in Section 4. It turns out that the right regularity condition of the perturbation is expressed in terms of the Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}_{\mu}$ induced by the self-transport kernel $k_{\mu}$. Specifically, $\mathbf{g}_{\mu}$ defines a positive definite quadratic form on the dual $\mathcal{H}_{\mu}^{*}$ of the RKHS $\mathcal{H}_{\mu}$, restricted to functionals $\dot{\mu}$ such that $\langle\dot{\mu},\mathbb{1}_{X}\rangle=0$. On this subspace of $\mathcal{H}_{\mu}^{*}$ the norm induced by the metric tensor $\mathbf{g}_{\mu}$ is equivalent to the operator norm. Interestingly enough, for a smooth cost function $c$ this space contains both the “vertical” perturbations $\mu_{t}=\mu+t\delta\mu$, as well as the “horizontal” perturbations $\mu_{t}=\Psi(t,\cdot)_{\#}\mu$ when $\Psi$ is the flow of a velocity field. In contrast, the tangent space for the optimal transport metric tensor (3) only allows for horizontal perturbations.
A useful change of variables. In order to build a distance from the metric tensor $\mathbf{g}_{\mu}$ and to prove the existence of geodesics we need some uniform coercivity. To that end we exploit the change of variables $\mu\leftrightarrow\beta:=\exp(-{f_{\mu,\mu}}/\varepsilon)$ which maps $\mathcal{P}(X)$ into a subset of the unit sphere of the Reproducing Kernel Hilbert Spaces $\mathcal{H}_{c}$ induced by $k_{c}:=\exp(-c/\varepsilon)$ (Theorem 4.8). This change of variables is not completely new and can be seen implicitly in [26, Prop. 3] when the authors derive lower bounds on the Sinkhorn divergence. For a perturbation $\dot{\beta}$ of the variable $\beta=\exp(-{f_{\mu,\mu}}/\varepsilon)$ and its associated perturbation $\dot{\mu}$ of the measure $\mu$, Theorem 4.11 expresses $\mathbf{g}_{\mu}$ as
$\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})=\tilde{\mathbf{g}}_{\mu}(\dot{\beta},\dot{\beta}):=\frac{\varepsilon}{2}\left(\lVert\dot{\beta}\rVert_{\mathcal{H}_{c}}^{2}+2\left\langle\exp({f_{\mu,\mu}}/\varepsilon)\cdot\dot{\beta},(\operatorname{id}-K_{\mu})^{-1}[\exp({f_{\mu,\mu}}/\varepsilon)\cdot\dot{\beta}]\right\rangle_{\mu}\right).$ (8)
Expressed in this space $\mathcal{H}_{c}$, independent of $\mu$, the metric tensor $\tilde{\mathbf{g}}_{\mu}(\dot{\beta},\dot{\beta})$ is jointly continuous in $\mu$ and $\dot{\beta}$ (Proposition 4.13) and equivalent to $\|\dot{\beta}\|^{2}_{\mathcal{H}_{c}}$ the squared norm on $\mathcal{H}_{c}$ uniformly in $\mu$ (Proposition 4.14). This is in stark contrast with the metric tensor $\mathbf{g}^{0}_{\mu}$ of the optimal transport problem (3), which depends in a non-smooth way on $\mu$ and which does not compare well to any other functional space uniformly on $\mathcal{P}(X)$.
Building a new distance. Once we understand well our metric tensor $\mathbf{g}_{\mu}$, we build our new metric on the space of probability measures in Section 5. While in the classical case the “dynamic” formula (4) is a consequence of the “static” definition (1), here we take it as definition of the distance. Specifically, in Definition 5.1 we set
$\mathsf{d}_{S}(\mu_{0},\mu_{1}):=\left(\inf\int_{0}^{1}\mathbf{g}_{\mu_{t}}(\dot{\mu}_{t},\dot{\mu}_{t})\,\mathrm{d}t\right)^{\frac{1}{2}}$ (9)
(for ease of notation we omit the dependence on $\varepsilon$), where the infimum is taken over a suitable set of admissible paths $(\mu_{t})_{t\in[0,1]}$ between $\mu_{0}$ and $\mu_{1}$. Theorem 5.2 shows that $\mathsf{d}_{S}$ is a metric which metrizes the weak-$\ast$ topology on $\mathcal{P}(X)$. We actually prove more: $\mathsf{d}_{S}(\mu_{0},\mu_{1})$ is uniformly equivalent to $\lVert\beta_{0}-\beta_{1}\rVert_{\mathcal{H}_{c}}$. This is to be compared to optimal transport, where $\sqrt{\mathrm{OT}_{0}}$ can be compared to an $H^{-1}$ norm as in [53] and [57, Sec. 5.5.2], but these results require some regularity assumptions on the measures involved and are not valid uniformly on $\mathcal{P}(X)$. Estimates can also be made with linearized optimal transport [23], though this only leads to a bi-Hölder equivalence. The price to pay is that our constants depend on $\varepsilon$ and blow up as $\varepsilon\to 0$.
Existence of geodesics. In Theorem 5.3 we prove existence of geodesics, that is, that the infimum in (9) is attained. In most of the problems which are variations of the Benamou-Brenier formula, such as unbalanced optimal transport [18], variational mean field games [6] or harmonic maps [39], the existence of solutions to problems of the type (9) uses the property that $\mathbf{g}^{0}_{\mu}(\dot{\mu},\dot{\mu})$ (or its suitable counterpart) is jointly convex in $\mu$ and $\dot{\mu}$. Thus tools of convex analysis can be used and lower semicontinuity is easy to establish. In our case, however, neither the Sinkhorn divergence $S_{\varepsilon}(\mu,\nu)$ nor the associated metric tensor $\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})$ are jointly convex: we provide counterexamples that we defer to Section 7.2. To prove that the infimum in (9) is attained we still rely on the direct method of calculus of variations, but we have to work a little bit harder: coercivity comes from the change of variables $\mu\leftrightarrow\beta$ studied in Section 4, and lower semicontinuity is proved thanks to the smoothness of the metric tensor in its inputs.
The quadratic case. Our theory applies to any compact metric space with any continuous cost function as soon as $k_{c}=\exp(-c/\varepsilon)$ defines a positive definite universal kernel. The special case of the squared Euclidean distance as the cost function on a subset of $\mathbb{R}^{d}$ is studied in Section 5.3. For measures $\mu_{0}$, $\mu_{1}$ it has been known for a while (see e.g. [8, Lemma 8.8] or the proof of Theorem 2.1 in [27]) that the following decomposition holds:
$\mathrm{OT}_{0}(\mu_{0},\mu_{1})=\lVert m_{1}-m_{0}\rVert^{2}+\mathrm{OT}_{0}(\bar{\mu}_{0},\bar{\mu}_{1}),$
where $m_{i}$ denotes the mean of the measure $\mu_{i}$ and $\bar{\mu}_{i}=(\operatorname{id}-m_{i})_{\#}\mu_{i}$ is the centered version of $\mu_{i}$, for $i=0,1$. In other words, the transport of the means and the transport of the centered parts decouple. We show that this extends to Sinkhorn divergences and that it can be read as an orthogonal decomposition of the tangent space for the geometry of $\mathbf{g}_{\mu}$ in Proposition 5.7. We then integrate this inequality and show in Theorem 5.9 that our novel distance satisfies
$\mathsf{d}_{S}(\mu_{0},\mu_{1})^{2}=\lVert m_{1}-m_{0}\rVert^{2}+\mathsf{d}_{S}(\bar{\mu}_{0},\bar{\mu}_{1})^{2}.$
In particular, uniform translations are geodesics, as in the case of classical optimal transport.
Some special cases. Section 6 covers two examples where we can explicitly compute the metric tensor. We start with Gaussians: it is one of the few examples in optimal transport for which there are closed formulas and for which the geometry is well understood [61]; and it is one of the paradigmatic examples in information geometry which relates Gaussian measures to the Poincaré half-plane, see e.g. [21]. Building on an explicit formula for $\mathrm{OT}_{\varepsilon}$ between Gaussians from [44], Theorem 6.1 shows that
$\mathbf{g}_{v}(\dot{v},\dot{v})=\frac{1}{4\sqrt{\frac{\varepsilon^{2}}{16}+v^{2}}}\dot{v}^{2}.$
Here we identify a one-dimensional Gaussian measure $\mathcal{N}(0,v)$ with its variance $v$. Theorem 6.2 then provides a formula for the geodesic restriction of $\mathsf{d}_{S}$ to Gaussian measures. Note that strictly speaking we are not in our framework since Gaussians are not supported on a compact space. As a second example, we consider the two-point space $X=\{x_{1},x_{2}\}$, where we can compute the metric tensor. In particular, in Proposition 6.3 we study the asymptotic of the distance in two different regimes, namely $c(x_{1},x_{2})\ll\varepsilon$ and $c(x_{1},x_{2})\gg\varepsilon$. The formula works for the squared distance cost function and asymptotically recovers the cost of horizontal movement on $\mathbb{R}^{d}$ as $\lVert x_{1}-x_{2}\rVert\to 0$.
Negative results. Section 7 collects some negative results on the Sinkhorn divergence and the metric tensor that were already mentioned throughout the introduction. Specifically, we provide examples showing that $\sqrt{S_{\varepsilon}}$ does not satisfy the triangle inequality, that $(\mu,\nu)\mapsto S_{\varepsilon}(\mu,\nu)$ is not jointly convex, and that the metric tensor $\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})$ is also not jointly convex in $\mu$ and $\dot{\mu}$.
Comparison with Schrödinger bridges.
We emphasize that our geometry and its geodesics are fundamentally different from Schrödinger bridges, a notion of displacement interpolation associated with entropic optimal transport [30, 41]. In our construction, given two probability measures $\mu_{0},\mu_{1}$ and $\varepsilon>0$ there exists $(\mu^{\text{geo},\varepsilon}_{t})_{t\in[0,1]}$ which is a geodesic in our metric $\mathsf{d}_{S}$, i.e. a minimizer in (9). In contrast, when $c=\lVert x-y\rVert^{2}$ is the quadratic cost on $\mathbb{R}^{d}$, the Schrödinger bridge $(\mu^{\text{Sbr},\varepsilon}_{t})_{t\in[0,1]}$ is constructed as follows [41]: Let $\pi_{\varepsilon}$ be the optimal transport plan, that is, the minimizer in (5). Let $q_{t,\varepsilon}(x,y)$ denote the law of the Brownian bridge at time $t$ starting at $t=0$ in $x$ and ending at $t=1$ in $y$ with diffusivity $\sqrt{\varepsilon/2}$. That is, $q_{t,\varepsilon}(x,y)\sim\mathcal{N}((1-t)x+ty,t(1-t)\varepsilon/2)$. Then $\mu^{\text{Sbr},\varepsilon}_{t}$ is given by
$\mu^{\text{Sbr},\varepsilon}_{t}=\iint_{X\times X}q_{t,\varepsilon}(x,y)\,\mathrm{d}\pi_{\varepsilon}(x,y).$ (10)
It is also a curve which satisfies $\mu^{\text{Sbr},\varepsilon}_{0}=\mu_{0}$, $\mu^{\text{Sbr},\varepsilon}_{1}=\mu_{1}$ and which interpolates smoothly in between. Nevertheless, the two notions of interpolation are conceptually very different. Let us point out three major differences between $\mu^{\text{geo},\varepsilon}$ and $\mu^{\text{Sbr},\varepsilon}$:
  1. 1.
    The geodesic interpolation $\mu^{\text{geo},\varepsilon}$ is defined on any compact space $X$ if the continuous cost function $c$ defines a universal positive definite kernel $k_{c}=\exp(-c/\varepsilon)$ whereas Schrödinger bridges need the quadratic cost on $\mathbb{R}^{d}$, or more generally a reference continuous-time Markov process (the Brownian motion in the case of quadratic cost) to be defined [41].
  2. 2.
    If $\mu_{0}=\mu_{1}=\mu$, then the geodesic interpolation $\mu^{\text{geo},\varepsilon}$ is constant in time and coincides with $\mu$. In the same case $\mu^{\text{Sbr},\varepsilon}$ is not constant in time, and typically has larger variance for intermediate $t\in(0,1)$ than for $t=0,1$. This is linked to the phenomenon that generically $\mathrm{OT}_{\varepsilon}(\mu,\mu)>0$; see also the discussion in [42], after Theorem 1.7.
  3. 3.
    In both cases for the quadratic cost the parameter $\sqrt{\varepsilon}$ acts as a typical length. In the case of Schrödinger bridges, it is also related to a typical time, whereas for geodesics there is no typical time. To understand this, for $\tau\in(0,1)$, let us look at the curve $(\mu_{t})_{t\in[0,\tau]}$ but rescaled to be over $[0,1]$. That is, let us consider $(\mu^{\text{geo},\varepsilon}_{t\tau})_{t\in[0,1]}$ and $(\mu^{\text{Sbr},\varepsilon}_{t\tau})_{t\in[0,1]}$. Then $(\mu^{\text{geo},\varepsilon}_{t\tau})_{t\in[0,1]}$ is a geodesic from $\mu_{0}$ to $\mu_{\tau}$ for any value of $\tau\in(0,1)$, as usual in Riemannian geometry. On the other hand, $(\mu^{\text{Sbr},\varepsilon}_{t\tau})_{t\in[0,1]}$ is a Schrödinger bridge between its endpoints $\mu^{\text{Sbr},\varepsilon}_{0}$ and $\mu^{\text{Sbr},\varepsilon}_{\tau}$, but with a regularization parameter $\varepsilon\tau$. This can be seen by using that the Schrödinger bridge is the marginal of a process which minimizes an entropy over the space of measures on paths [41], and to use the concatenation property of the entropy for Markov processes in this space [5, Lemma 3.4]. The key insight is that, when rescaled from time $[0,\tau]$ to time $[0,1]$, a Brownian motion with diffusivity $\sqrt{\varepsilon/2}$ becomes a Brownian motion with diffusivity $\sqrt{\tau\varepsilon/2}$. So in particular as $\tau\to 0$, that is, as we “zoom in” in time on a Schrödinger bridge, it increasingly resembles a Schrödinger bridge with $\varepsilon=0$, that is, a Wasserstein geodesic [40, 48].
We also refer to Figure 1 with Gaussian measures where we can do explicit computations for our geodesics and Schrödinger bridges.
Setting and notation.
We consider a compact metric space $(X,d)$. We fix $\varepsilon>0$ and a symmetric, continuous, non-negative cost function $c\in\mathcal{C}(X\times X;[0,\infty))$ that induces a positive definite universal kernel $k_{c}(x,y)=\exp(-c(x,y)/\varepsilon)$.
We work with the space of signed Radon measures $\mathcal{M}(X)$, which is identified with the dual space of $\mathcal{C}(X)$ through the Riesz–Markov–Kakutani Theorem. We denote the subset of non-negative measures by $\mathcal{M}_{+}(X)$, the space of “balanced measures” $\nu\in\mathcal{M}(X)$ with $\int\mathbb{1}_{X}\,\mathrm{d}\nu=0$ by $\mathcal{M}_{0}(X)$, and Radon probability measures by $\mathcal{P}(X)$.
[…]
2 Preliminaries on entropic optimal transport
The entropic optimal transport problem, as defined in (5), is a constrained convex optimization problem whose dual formulation reads as
$\mathrm{OT}_{\varepsilon}(\mu,\nu)=\sup_{f,g\in\mathcal{C}(X)}\langle\mu,f\rangle+\langle\nu,g\rangle-\varepsilon\left\langle\mu\otimes\nu,\exp\left(\frac{1}{\varepsilon}(f\oplus g-c)\right)-1\right\rangle.$ (11)
Here the notation $f\oplus g$ stands for the function $(x,y)\mapsto f(x)+g(y)$ defined on $X\times X$. The dual problem admits maximizers ${f_{\mu,\nu}}$, ${g_{\mu,\nu}}$, which we call Schrödinger potentials, that satisfy
$\mathrm{OT}_{\varepsilon}(\mu,\nu)=\int_{X}{f_{\mu,\nu}}\,\mathrm{d}\mu+\int_{X}{g_{\mu,\nu}}\,\mathrm{d}\nu$ (12)
and are solutions of the Schrödinger system
$\begin{cases}{f_{\mu,\nu}}&=T_{\varepsilon}({g_{\mu,\nu}},\nu),\\ {g_{\mu,\nu}}&=T_{\varepsilon}({f_{\mu,\nu}},\mu),\end{cases}$ (13)
where $T_{\varepsilon}:C(X)\times\mathcal{P}(X)\to C(X)$ is defined by
$T_{\varepsilon}(f,\mu)(y)\coloneqq-\varepsilon\log\int_{X}\exp\left(\frac{1}{\varepsilon}(f(x)-c(x,y))\right)\,\mathrm{d}\mu(x).$ (14)
We will sometimes write the condition ${g_{\mu,\nu}}=T_{\varepsilon}({f_{\mu,\nu}},\mu)$ as
$\int_{X}\exp\left(\frac{1}{\varepsilon}({f_{\mu,\nu}}(x)+{g_{\mu,\nu}}(y)-c(x,y))\right)\,\mathrm{d}\mu(x)=1$ (15)
for every $y\in X$. The Schrödinger potentials are unique in the space $\mathcal{C}(X)\times\mathcal{C}(X)/\sim$ with the relation $(f,g)~{}\sim~{}(f~{}+~{}\lambda\mathbb{1}_{X},~{}g~{}-~{}\lambda\mathbb{1}_{X})$ for all $\lambda\in\mathbb{R}$. They inherit the modulus of continuity from the cost function $c$ [50, Lem. 3.1]. Uniqueness of the potentials up to constant shifts yields that we can always choose ${f_{\mu,\nu}}={g_{\nu,\mu}}$, and in particular, throughout this article we will adopt the convention that ${f_{\mu,\mu}}={g_{\mu,\mu}}$. The optimal transport plan $\pi$ in (5) is given by
$\pi=\exp\left(\frac{1}{\varepsilon}({f_{\mu,\nu}}\oplus{g_{\mu,\nu}}-c)\right)(\mu\otimes\nu).$ (16)
Note that the function ${f_{\mu,\nu}}\oplus{g_{\mu,\nu}}$ is unambiguous since constant shifts cancel out. Optimal $({f_{\mu,\nu}},{g_{\mu,\nu}})$ can be obtained numerically by iterating the fixed point equation (13), which corresponds to the celebrated Sinkhorn algorithm. We refer to [54] for an overview on computational optimal transport, and to [41, 49] for a detailed discussion on entropic optimal transport and related literature.
3 The Hessian of the Sinkhorn divergence
3.1 Statement of the expansion
Theorem 1.1 shows that the Sinkhorn divergence (6) is positive definite and convex in each of its input variables. However, as we show in Section 7.1, $S_{\varepsilon}$, or more precisely $\sqrt{S_{\varepsilon}}$, in general does not satisfy the triangle inequality. To construct the metric $\mathsf{d}_{S}$ announced in (9) in the introduction, we first characterize how $S_{\varepsilon}$ behaves close to the diagonal $\mu=\nu$. Specifically, we want to find an expansion of the form
$S_{\varepsilon}(\mu,\mu_{t})\sim t^{2}\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})$ (17)
as $t\to 0$ for a suitably differentiable path $(\mu_{t})_{t}$ in $\mathcal{P}(X)$ with $\mu_{0}=\mu$, $\dot{\mu}_{0}=\dot{\mu}$. The term $\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})$ will then serve as the metric tensor and it will be examined in detail in Section 4. By Theorem 1.1 there is a good hope that $\mathbf{g}_{\mu}$ will be a positive quadratic form on a meaningful tangent space to probability measures. The appropriate notion of tangent space will become apparent in Section 4. In this section we restrict to a simple class of perturbations.
Definition 3.1 (Vertical perturbation).
Let $\mu\in\mathcal{P}(X)$, $\delta\mu\in\mathcal{M}_{0}(X)$. We call the path
$(\mu_{t})_{t}=(\mu+t\delta\mu)_{t}$ (18)
a vertical perturbation of $\mu$ if there exists $\tau>0$ such that $\mu_{t}\in\mathcal{P}(X)$ for all $t\in(-\tau,\tau)$.
A measure $\delta\mu\in\mathcal{M}_{0}(X)$ induces a vertical perturbation of $\mu$ if and only if $\delta\mu\ll\mu$ and the Radon–Nikodym derivative $\frac{\mathrm{d}\delta\mu}{\mathrm{d}\mu}$ is uniformly bounded from above and below. The constraint $\langle\delta\mu,\mathbb{1}_{X}\rangle=0$ in the space of balanced measures $\mathcal{M}_{0}(X)$ comes from the preservation of mass.
Before stating the main result of this section, we need to introduce integral operators related to the transport kernel. Since we assume that $c\in\mathcal{C}(X\times X)$, the Schrödinger potentials ${f_{\mu,\nu}},{g_{\mu,\nu}}$ also belong to $\mathcal{C}(X)$. We define the self-transport kernel as follows:
$k_{\mu}(x,y):=\exp\left(\frac{1}{\varepsilon}\left({f_{\mu,\mu}}(x)+{f_{\mu,\mu}}(y)-c(x,y)\right)\right).$
Recall here that we adopt the convention ${f_{\mu,\mu}}={g_{\mu,\mu}}$ so that $k_{\mu}$ is unambiguous.
Definition 3.2.
For $\mu\in\mathcal{P}(X)$ we define the kernel operators $K_{\mu}$ and $H_{\mu}$ by
$K_{\mu}\colon\mathcal{C}(X)\to\mathcal{C}(X),\quad K_{\mu}[\phi](y)=\int_{X}\phi(x)k_{\mu}(x,y)\,\mathrm{d}\mu(x),$
$H_{\mu}\colon\mathcal{M}(X)\to\mathcal{C}(X),\quad H_{\mu}[\nu](y)=\int_{X}k_{\mu}(x,y)\,\mathrm{d}\nu(x).$
The operator $K_{\mu}$ can be extended to the domain $L^{1}(X,\mu)$. If $\phi\in L^{1}(X,\mu)$, then $K_{\mu}[\phi]=H_{\mu}[\phi\mu]$. The operator $H_{\mu}$ is the kernel mean embedding of measures into the Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}_{\mu}$ induced by $k_{\mu}$, and its domain can be extended to $\mathcal{H}_{\mu}^{*}$. This will be discussed in Section 4.
The space $\mathcal{M}_{0}(X)$ is dual to $\mathcal{C}(X)/\mathbb{R}$, the space of continuous functions identified up to shifts by constant functions (identified with $\mathbb{R}$). Hence the pairing $\langle\nu,\phi\rangle$ is well defined between $\nu\in\mathcal{M}_{0}(X)$ and $\phi\in\mathcal{C}(X)/\mathbb{R}$. The operator $K_{\mu}$ descends to a well-defined operator
$K_{\mu}\colon\mathcal{C}(X)/\mathbb{R}\to\mathcal{C}(X)/\mathbb{R}$
because $K_{\mu}[\mathbb{1}_{X}]=\mathbb{1}_{X}$ by (15). Using the canonical projection $\mathcal{C}(X)\to\mathcal{C}(X)/\mathbb{R}$ we can always move to the quotient space.
Theorem 3.3 (Hessian of the Sinkhorn divergence).
Let $\mu\in\mathcal{P}(X)$, $\delta\mu\in\mathcal{M}_{0}(X)$, such that $\mu_{t}=\mu+t\delta\mu$ is a vertical perturbation of $\mu$. Then
$\lim_{t\to 0}\frac{S_{\varepsilon}(\mu,\mu_{t})}{t^{2}}=\frac{\varepsilon}{2}\left\langle\delta\mu,(\operatorname{id}-K_{\mu}^{2})^{-1}H_{\mu}\left[\delta\mu\right]\right\rangle.$ (19)
The expression on the right-hand side is well defined, in the sense that $(\operatorname{id}-K_{\mu}^{2})^{-1}$ exists as an operator on $\mathcal{C}(X)/\mathbb{R}$, and thus the duality pairing is understood between $\mathcal{M}_{0}(X)$ and $\mathcal{C}(X)/\mathbb{R}$. In Section 4 we will take the formula on the right-hand side of (19) as a starting point and analyze for which class of perturbations $\delta\mu$ we can make sense of it.
Remark 3.4 (Discrete measures).
When $\mu=\sum_{i=1}^{n}a_{i}\delta_{x_{i}}$ is a discrete measure with $a_{i}>0$ and the vertical perturbation reads $\delta\mu=\sum_{i=1}^{n}b_{i}\delta_{x_{i}}$ with $\sum_{i}b_{i}=0$ the formula can be expanded as follows. With $f={f_{\mu,\mu}}$, we introduce the $n\times n$ matrices $K,H$ as
$K_{ij}=\exp\left(\frac{1}{\varepsilon}(f(x_{i})+f(x_{j})-c(x_{i},x_{j}))\right)a_{j},\quad H_{ij}=\exp\left(\frac{1}{\varepsilon}(f(x_{i})+f(x_{j})-c(x_{i},x_{j}))\right)$
for indices $i,j\in\{1,\ldots,n\}$. With $b=(b_{i})_{i}\in\mathbb{R}^{n}$ Theorem 3.3 reads
$\lim_{t\to 0}\frac{S_{\varepsilon}(\mu,\mu_{t})}{t^{2}}=\frac{\varepsilon}{2}b^{\top}(\operatorname{id}-K^{2})^{\dagger}Hb,$
where $(\operatorname{id}-K^{2})^{\dagger}$ denotes the Moore-Penrose pseudo-inverse of the matrix $\operatorname{id}-K^{2}$.
The proof of Theorem 3.3 relies on the differentiability of the Schrödinger potentials along the path $(\mu_{t})_{t}$ and on explicit formulas for the derivatives. We follow the line of the works [14, 16] which use the implicit function theorem. Note however that these works are in the more involved context of a multi-marginal optimal transport problem. As our context is a bit different and the proofs in Section 3.3 a bit simpler, we include them for the sake of completeness.
Remark 3.5 (Extending Theorem 3.3 to horizontal perturbations).
All steps in the proof of Theorem 3.3 can be carried out in a similar fashion when $X\subseteq\mathbb{R}^{d}$ is the closure of a bounded open set, $c\in\mathcal{C}^{3}(X\times X)$, and the path follows a horizontal perturbation
$\mu_{t}=\Psi(t,\cdot)_{\#}\mu$ (20)
for a continuous function $\Psi\colon(-\tau,\tau)\times X\to X$ admitting two continuous time derivatives. Then $\frac{\partial^{k}\mu_{t}}{\partial t^{k}}$ exist in $\mathcal{C}^{k}(X)^{*}$ for $k=1,2$ with
$\left\langle\frac{\partial^{k}\mu_{t}}{\partial t^{k}},\phi\right\rangle=\int_{X}\frac{\partial^{k}}{\partial t^{k}}\phi(\Psi(t,x))\,\mathrm{d}\mu(x)$
for $\phi\in\mathcal{C}^{k}(X)$. When $\Psi$ is the flow of a (possibly time-dependent) vector field $v$, then this equation with $k=1$ reads $\dot{\mu}_{t}=-\operatorname{div}(v\mu_{t})$. The time derivatives of the Schrödinger potentials we compute in Proposition 3.11 then even exist in $\mathcal{C}^{1}(X)/\mathbb{R}$. We return to this discussion in Remark 3.14.
3.2 Preliminary results on the operators $K_{\mu}$ and $H_{\mu}$
In this section we analyze the operators $K_{\mu}$ and $H_{\mu}$ which will appear naturally when differentiating the fixed point equation (13) that characterizes the Schrödinger potentials. We then compute the derivatives of the Schrödinger potentials, and finally prove Theorem 3.3 in Section 3.3.
Proposition 3.6.
Let $\mu\in\mathcal{P}(X)$. The operators $H_{\mu}\colon\mathcal{M}(X)\to\mathcal{C}(X)$ and $K_{\mu}\colon\mathcal{C}(X)\to\mathcal{C}(X)$ are continuous and compact.
Proof.
As $K_{\mu}[\phi]=H_{\mu}[\phi\mu]$, it is enough to prove the result for $H_{\mu}$. From [50, Lem. 3.1] the set $\left\{k_{\mu}(x,\cdot)\,\middle|\,x\in X\right\}$ is equi-continuous. This easily implies equi-continuity of the set $\left\{H_{\mu}[\nu]\,\middle|\,\nu\in\mathcal{M}(X),~{}\lVert\nu\rVert_{\mathrm{TV}}\leq 1\right\}$, where .$\lVert.\rVert_{\mathrm{TV}}$ denotes the total variation norm. Equi-boundedness follows from $\lVert H_{\mu}[\nu]\rVert_{\infty}\leq\lVert k_{\mu}\rVert_{\infty}\cdot\lVert\nu\rVert_{\mathrm{TV}}$. Thus compactness of $H_{\mu}$ follows thanks to the Arzelà–Ascoli Theorem. ∎
We recall that $\mathcal{C}(X)/\mathbb{R}$ is the quotient space of continuous functions by constant functions. We equip this space with the quotient norm
$\lVert\phi\rVert_{\mathcal{C}/\mathbb{R}}:=\inf_{\lambda\in\mathbb{R}}\lVert\phi-\lambda\cdot\mathbb{1}_{X}\rVert_{\infty}=\frac{1}{2}\left(\sup_{x\in X}\phi(x)-\inf_{y\in X}\phi(y)\right).$
Proposition 3.7.
Let $q:=1-\exp(-4\lVert c\rVert_{\infty}/\varepsilon)\in(0,1)$. Then $\lVert K_{\mu}\phi\rVert_{\mathcal{C}/\mathbb{R}}\leq q\cdot\lVert\phi\rVert_{\mathcal{C}/\mathbb{R}}$ for all $\phi\in\mathcal{C}(X)$.
Proof.
By [24, Thm. 1.2] the Schrödinger potentials ${f_{\mu,\nu}},{g_{\mu,\nu}}\in\mathcal{C}(X)$ can be chosen such that $\lVert{f_{\mu,\nu}}\rVert_{\infty},\lVert{g_{\mu,\nu}}\rVert_{\infty}\leq\frac{3}{2}\lVert c\rVert_{\infty}$. With $a=\exp(-4\lVert c\rVert_{\infty}/\varepsilon)$ and using $c\geq 0$, we see that $k_{\mu}(y,z)\geq a\cdot k_{\mu}(x,z)$ for any triplet $x,y,z\in X$. For any $x,y\in X$, this, together with (15), leads to
$K_{\mu}\phi(x)-K_{\mu}\phi(y)$
=$\qquad=\int_{\{\phi>0\}}\phi(z)\big{(}k_{\mu}(x,z)-k_{\mu}(y,z)\big{)}\,\mathrm{d}\mu(z)+\int_{\{\phi<0\}}-\phi(z)\big{(}-k_{\mu}(x,z)+k_{\mu}(y,z)\big{)}\,\mathrm{d}\mu(z)$
$\qquad\leq\int_{\{\phi>0\}}\phi(z)\left(1-a\right)k_{\mu}(x,z)\,\mathrm{d}\mu(z)+\int_{\{\phi<0\}}-\phi(z)\left(1-a\right)k_{\mu}(y,z)\,\mathrm{d}\mu(z)$
$\qquad\leq 2\left(1-a\right)\cdot\lVert\phi\rVert_{\infty}.$
The operator norm estimate with respect to the quotient norm .$\lVert.\rVert_{\mathcal{C}/\mathbb{R}}$ on domain and codomain follows by taking a supremum over $x$ and $y$ and considering the infimum over shifts $\phi-\lambda$ by constants $\lambda\in\mathbb{R}$. Finally $q=1-a$. ∎
From this estimate we prove that the inverse of the map $\operatorname{id}-K_{\mu}^{2}$, central in the expansion of the Sinkhorn divergence, actually exists. We use $\mathbb{B}(\mathbb{X})$ to denote the set of bounded linear operators on a normed space $\mathbb{X}$.
Theorem 3.8.
Let $\mu\in\mathcal{P}(X)$. Then $(\operatorname{id}-K_{\mu}^{2})^{-1}$ exists in $\mathbb{B}(\mathcal{C}(X)/\mathbb{R})$ and $(\operatorname{id}+K_{\mu})^{-1}$ exists in $\mathbb{B}(\mathcal{C}(X))$.
The result will be strengthened in Theorem 4.3 in the next Section.
Proof.
$\operatorname{id}-K_{\mu}^{2}$ and $\operatorname{id}+K_{\mu}$ are compact perturbations of the identity on, respectively, the Banach spaces $\mathcal{C}(X)/\mathbb{R}$ and $\mathcal{C}(X)$. To check that they are invertible, by the Fredholm alternative [12, Theorem 6.6] it is enough to verify that the equations $\phi=K_{\mu}^{2}[\phi]$ and $\phi=-K_{\mu}[\phi]$ only have the trivial solution $\phi=0$. In both cases Proposition 3.7 directly yields that $\lVert\phi\rVert_{\mathcal{C}/\mathbb{R}}=0$. In the case $\phi=K_{\mu}^{2}(\phi)$, this gives $\phi=0$ in $\mathcal{C}(X)/\mathbb{R}$. For the equation $\phi=-K_{\mu}(\phi)$, this yields that $\phi$ is constant, and from $K_{\mu}[\mathbb{1}_{X}]=\mathbb{1}_{X}$ we obtain $\phi=0$. ∎
Remark 3.9 (A probabilistic interpretation).
Note that $K_{\mu}$ can be interpreted as follows. For $\mu\in\mathcal{P}(X)$, we can consider a Markov chain on $X$ with equilibrium measure $\mu$, and whose transition kernel is given by $x\mapsto k_{\mu}(x,y)\,\mathrm{d}\mu(y)$. The operator $K_{\mu}$ is simply the dual Markov operator of this chain. Indeed, the entropic optimal transport coupling $\pi$ between $\mu$ and itself, given by (16), is the law of the pair of random variables $(x_{1},x_{2})$ following this Markov chain if the initial distribution is $\mu$.
As $k_{\mu}$ is symmetric, we see that the Markov chain is reversible. The spectrum of the operator $K_{\mu}$ is studied in more detail in Proposition 4.2 which builds on Proposition 3.7: all the eigenvalues belong to $[0,1]$, and the spectral gap of $\operatorname{id}-K_{\mu}$ is at least $\exp(-4\lVert c\rVert_{\infty}/\varepsilon)$.
3.3 Proof of the expansion
We are now ready for the proof of Theorem 3.3. As mentioned earlier, we will need to differentiate the Schrödinger potentials with respect to the input measures. Following [14, 16] this will be done with the help of the implicit function theorem. In the rest of this section, a vertical perturbation $\mu_{t}=\mu+t\delta\mu$ for $t\in(-\tau,\tau)$ is fixed. As we will use the fixed point equation (13) defining the optimal pair of potentials $(f_{\mu_{t},\mu_{s}},g_{\mu_{t},\mu_{s}})$, we first need to understand the derivatives of the operator $T_{\varepsilon}$ featured in it. Whereas $T_{\varepsilon}$ is a priori defined on $\mathcal{C}(X)\times\mathcal{P}(X)$, from $T_{\varepsilon}(f+\lambda\mathbb{1}_{X},\mu)=T_{\varepsilon}(f,\mu)-\lambda\mathbb{1}_{X}$ for any $f,\mu$ and $\lambda\in\mathbb{R}$, we see that $T_{\varepsilon}$ descends to an operator defined on $\mathcal{C}(X)/\mathbb{R}\times\mathcal{P}(X)$ and valued in $\mathcal{C}(X)/\mathbb{R}$.
Lemma 3.10.
Let the path $\mu_{t}=\mu+t\delta\mu$ for $t\in(-\tau,\tau)$ be a vertical perturbation of $\mu~{}\in~{}\mathcal{P}(X)$. Then the parametrized map $\mathcal{C}(X)\times(-\tau,\tau)\to\mathcal{C}(X),~{}(f,t)\mapsto T_{\varepsilon}(f,\mu_{t})$ built from (14) is continuously differentiable. It satisfies
$D_{1}T_{\varepsilon}({f_{\mu,\mu}},\mu)[\phi]=-K_{\mu}[\phi],\qquad\frac{\mathrm{d}}{\mathrm{d}t}\Big{|}_{t=0}T_{\varepsilon}({f_{\mu,\mu}},\mu_{t})=-\varepsilon H_{\mu}[\delta\mu].$
These formulas remain true when $\mathcal{C}(X)$ is replaced by $\mathcal{C}(X)/\mathbb{R}$.
Proof.
By the chain rule
$D_{1}T_{\varepsilon}(f,\mu)[\phi](y)=-\varepsilon\frac{\int\frac{1}{\varepsilon}\phi(x)\exp\left(\frac{1}{\varepsilon}(f(x)-c(x,y))\right)\,\mathrm{d}\mu(x)}{\int\exp\left(\frac{1}{\varepsilon}(f(x)-c(x,y))\right)\,\mathrm{d}\mu(x)}.$
As $T_{\varepsilon}({f_{\mu,\mu}},\mu)={f_{\mu,\mu}}$ the denominator is equal to $\exp\left(-\frac{1}{\varepsilon}{f_{\mu,\mu}}(y)\right)$, thus finding
$D_{1}T_{\varepsilon}({f_{\mu,\mu}},\mu)[\phi](y)=-\int_{X}\phi(x)\exp\left(\frac{1}{\varepsilon}({f_{\mu,\mu}}(y)+{f_{\mu,\mu}}(x)-c(x,y))\right)\,\mathrm{d}\mu(x)=-K_{\mu}[\phi](y).$
For the derivative in the second component, using $k(x,y)=\exp(\frac{1}{\varepsilon}(f(x)+f(y)-c(x,y)))$ we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\Big{|}_{t=0}T_{\varepsilon}(f,\mu_{t})(y)=\frac{\mathrm{d}}{\mathrm{d}t}\Big{|}_{t=0}\left(-\varepsilon\log\left\langle\mu_{t},k(\cdot,y)\right\rangle\right)=-\varepsilon\frac{\langle\delta\mu,k(\cdot,y)\rangle}{\langle\mu,k(\cdot,y)\rangle},$
which at $f={f_{\mu,\mu}}$, that is, $k=k_{\mu}$, simplifies to $-\varepsilon\langle\delta\mu,k_{\mu}(\cdot,y)\rangle=-\varepsilon H_{\mu}[\delta\mu](y)$. Lemma A.1 shows that the convergence is uniform in $y$.
The extension to the case where $f\in\mathcal{C}(X)/\mathbb{R}$ instead of $\mathcal{C}(X)$ is straightforward. ∎
Proposition 3.11.
Let the path $\mu_{t}=\mu+t\delta\mu$ for $t\in(-\tau,\tau)$ be a vertical perturbation of $\mu\in\mathcal{P}(X)$. Then the Schrödinger potentials $f_{t,s}=f_{\mu_{t},\mu_{s}}$ are continuously differentiable in $\mathcal{C}(X)/\mathbb{R}$ with respect to the time parameters in a neighborhood of $(t,s)=(0,0)$. Moreover, they satisfy
$\begin{split}&\frac{\partial f_{0,s}}{\partial s}\Big{|}_{s=0}=-\varepsilon(\operatorname{id}-K_{\mu}^{2})^{-1}H_{\mu}[\delta\mu]\in\mathcal{C}(X)/\mathbb{R},\\ &\frac{\partial f_{t,0}}{\partial t}\Big{|}_{t=0}=\varepsilon K_{\mu}(\operatorname{id}-K_{\mu}^{2})^{-1}H_{\mu}[\delta\mu]\in\mathcal{C}(X)/\mathbb{R},\\ &\frac{\partial f_{t,t}}{\partial t}\Big{|}_{t=0}=-\varepsilon(\operatorname{id}+K_{\mu})^{-1}H_{\mu}[\delta\mu]\in\mathcal{C}(X).\end{split}$ (21)
Remark 3.12.
The proofs of Lemma 3.10 and Proposition 3.11 extend directly to the case where difference quotients for the path $(\mu_{t})_{t}$ in $\mathcal{P}(X)\subseteq\mathcal{M}(X)$ converge weak-$\ast$ and the derivative is weak-$\ast$-continuous. We will discuss the extension to horizontal perturbations in Remark 3.14.
Proof of Proposition 3.11.
We use that $f_{t,s}$ and $g_{t,s}=f_{s,t}$ are characterized by the fixed point equation (13). Specifically, we have
$f_{t,s}=T_{\varepsilon}(g_{t,s},\mu_{s})=T_{\varepsilon}(T_{\varepsilon}(f_{t,s},\mu_{t}),\mu_{s}),$ (22)
for all $t$ and $s$. We will use the implicit function theorem [25, Thm. 10.2.1] applied to the map
$\tilde{T}\colon\mathcal{C}(X)/\mathbb{R}\times(-\tau,\tau)^{2}\to\mathcal{C}(X)/\mathbb{R},\quad(f,t,s)\mapsto(f-T_{\varepsilon}(T_{\varepsilon}(f,\mu_{t}),\mu_{s})).$
The condition $\tilde{T}(f,t,s)=0$ is equivalent to $f=f_{t,s}$ in $\mathcal{C}(X)/\mathbb{R}$ by (22). The assumptions of [25, Thm. 10.2.1] ask that the map $\tilde{T}$ is of class $C^{1}$ and that $D_{1}\tilde{T}$ is invertible at the point $(t,s)=(0,0)$, $f=f_{0,0}$. The former is a direct consequence of Lemma 3.10), while for the latter we have $D_{1}\tilde{T}(f_{0,0},0,0)=\operatorname{id}-K_{\mu}^{2}$ by Lemma 3.10 and this is invertible on $\mathcal{C}(X)/\mathbb{R}$ by Theorem 3.8.
We can thus apply [25, Thm. 10.2.1], which ensures that the solution $f=f_{t,s}$ solving $\tilde{T}(f_{t},t,s)=0$ is continuously differentiable in $\mathcal{C}(X)/\mathbb{R}$ with respect to the time parameters in a neighborhood of $(t,s)=(0,0)$, and the partial derivatives are given by
$\frac{\partial f_{t,0}}{\partial t}\Big{|}_{t=0}$ $=-(D_{1}\tilde{T}(f_{0,0},0))^{-1}\left(\frac{\partial\tilde{T}(f_{0,0},t,0)}{\partial t}\Big{|}_{t=0}\right),$
$\frac{\partial f_{0,s}}{\partial s}\Big{|}_{s=0}$ $=-(D_{1}\tilde{T}(f_{0,0},0))^{-1}\left(\frac{\partial\tilde{T}(f_{0,0},0,s)}{\partial s}\Big{|}_{s=0}\right).$
Following Lemma 3.10 we obtain
$\frac{\partial\tilde{T}(f_{0,0},t,0)}{\partial t}\Big{|}_{t=0}=-\varepsilon K_{\mu}H_{\mu}[\delta\mu],\qquad\frac{\partial\tilde{T}(f_{0,0},0,s)}{\partial s}\Big{|}_{s=0}=\varepsilon H_{\mu}[\delta\mu].$
This yields the first two derivatives.
To calculate the time derivative of $f_{t,t}$ we may drop the quotient and use the map $\tilde{T}^{\prime}(f,t):=f-T_{\varepsilon}(f,\mu_{t})$ defined on $\mathcal{C}(X)\times(-\tau,\tau)$ instead of $\tilde{T}$. Here $D_{1}\tilde{T}^{\prime}(f_{0,0},0)=\operatorname{id}+K_{\mu}$ whose invertibility in the space $\mathcal{C}(X)$ can be found once again in Theorem 3.8. Analogous arguments to the previous case give the result. ∎
Lemma 3.13.
Let $U\subseteq\mathbb{R}^{2}$ be open with $(0,0)\in U$ and let $F\in\mathcal{C}^{2}(U)$. Then
$\lim_{t\to 0}\frac{F(0,t)+F(t,0)-F(0,0)-F(t,t)}{t^{2}}=-\frac{\partial^{2}F}{\partial t\partial s}(0,0).$
Proof.
Immediate from the second-order Taylor expansion at $(t,s)=(0,0)$. ∎
This lemma provides the final tool to obtain the expansion of the Sinkhorn divergence:
Proof of Theorem 3.3.
Denote the Schrödinger potentials by $f_{t,s}:=f_{\mu_{t},\mu_{s}}$. Define $F(t,s):=\mathrm{OT}_{\varepsilon}(\mu_{t},\mu_{s})$. As this function is symmetric we have
$S_{\varepsilon}(\mu,\mu_{t})=\frac{F(0,t)+F(t,0)-F(0,0)-F(t,t)}{2}.$
So provided we prove that $F$ is of class $\mathcal{C}^{2}$ around $(0,0)$, Lemma 3.13 yields
$\lim_{t\to 0}\frac{S_{\varepsilon}(\mu,\mu_{t})}{t^{2}}=-\frac{1}{2}\frac{\partial^{2}F}{\partial t\partial s}(0,0).$
Taking one derivative of $F$ is standard. It can be found for instance in [26, Prop. 2]. Thus we get
$\frac{\partial F}{\partial t}(t,s)=\left\langle\frac{\partial\mu_{t}}{\partial t},f_{t,s}\right\rangle.$
For a vertical perturbation, we have $\frac{\partial^{2}\mu_{t}}{\partial t^{2}}=0$. The continuous differentiability of $f_{t,s}$ by the time parameters from Proposition 3.11 ensures that $F$ is of class $\mathcal{C}^{2}$ around $(0,0)$. Moreover, we see that
$\frac{\partial^{2}F}{\partial t\partial s}(0,0)=\left\langle\frac{\partial\mu_{t}}{\partial t}\Big{|}_{t=0},\frac{\partial f_{0,s}}{\partial s}\Big{|}_{s=0}\right\rangle.$
Note that on the right-hand side we have the derivative of the “first” Schrödinger potential when the “second” measure is changed. From Proposition 3.11 we obtain
$\frac{\partial f_{0,s}}{\partial s}\Big{|}_{s=0}=-\varepsilon(\operatorname{id}-K_{\mu}^{2})^{-1}H_{\mu}\left[\delta\mu\right]\in\mathcal{C}(X)/\mathbb{R}.$
We conclude that
$\lim_{t\to 0}\frac{S_{\varepsilon}(\mu,\mu_{t})}{t^{2}}=-\frac{1}{2}\frac{\partial^{2}F}{\partial t\partial s}(0,0)=-\frac{1}{2}\left\langle\frac{\partial\mu_{t}}{\partial t}\Big{|}_{t=0},\frac{\partial f_{0,s}}{\partial s}\Big{|}_{s=0}\right\rangle=\frac{\varepsilon}{2}\left\langle\delta\mu,(\operatorname{id}-K_{\mu}^{2})^{-1}H_{\mu}[\delta\mu]\right\rangle.\qed$
Remark 3.14.
We add to Remarks 3.5 and 3.12: When $X\subseteq\mathbb{R}^{d}$ is the closure of its interior and $c\in\mathcal{C}^{2}(X\times X)$, we can extend Proposition 3.11 to paths where $\dot{\mu}_{t}\in\mathcal{C}^{1}(X)^{*}$ exists as a distribution and the path $t\mapsto\dot{\mu}_{t}$ is weak-$\ast$-continuous in $\mathcal{C}^{1}(X)^{*}$ in a neighborhood of $t=0$ by replacing the space $\mathcal{C}(X)$ with $\mathcal{C}^{1}(X)$ everywhere. Lemma 3.10 also remains valid then, which can be seen by differentiating everything once in the $y$-coordinate. The proof for [26, Prop. 2] used in the proof of Theorem 3.3 directly extends to such paths as well. The only part which might fail is the smoothness assumption for Lemma 3.13. When $c\in\mathcal{C}^{3}(X\times X)$ and the path follows a horizontal perturbation (20), the function $F$ is of class $\mathcal{C}^{2}$ around $(0,0)$, as then $\frac{\partial^{2}\mu_{t}}{\partial t^{2}}$ is a second-order distribution and $t\mapsto f_{t,s}\in\mathcal{C}^{2}(X)/\mathbb{R}$ is continuous [14, Thm. 1.1]. This shows that Theorem 3.3 also holds true for horizontal perturbations.
4 The metric tensor and its regularity
In this section, we analyze the expression obtained as the second-order approximation of the Sinkhorn divergence in Theorem 3.3. We provide a way to extend the expression to a larger class of perturbations. Recall that $\mathcal{M}_{0}(X)$ is the set of balanced signed measures $\nu\in\mathcal{M}(X)$, such that $\langle\nu,\mathbb{1}_{X}\rangle=0$.
Definition 4.1.
For $\mu\in\mathcal{P}(X)$ and $\dot{\mu}_{1}$, $\dot{\mu}_{2}$ in $\mathcal{M}_{0}(X)$, we define the metric tensor $\mathbf{g}_{\mu}(\dot{\mu}_{1},\dot{\mu}_{2})$ as
$\mathbf{g}_{\mu}(\dot{\mu}_{1},\dot{\mu}_{2})=\frac{\varepsilon}{2}\left\langle\dot{\mu}_{1},(\operatorname{id}-K_{\mu}^{2})^{-1}H_{\mu}[\dot{\mu}_{2}]\right\rangle.$
We will see that $\mathbf{g}_{\mu}$ defines an inner product on $\mathcal{M}_{0}(X)$ whose completion is canonically identified with (a subset of) the dual of a Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}_{\mu}$. However, this RKHS, which can be thought of as a tangent space to the space of probability measures at $\mu$, depends on $\mu$. We thus do a global change of variables to embed the set of probability distributions $\mathcal{P}(X)$ into the unit sphere of another RKHS $\mathcal{H}_{c}$. This makes all tangent spaces subspaces of $\mathcal{H}_{c}$. Further, the metric tensor $\tilde{\mathbf{g}}_{\mu}$, expressed in these new coordinates, is equivalent to the norm on $\mathcal{H}_{c}$ (uniformly in $\mu$). This change of variables will be especially useful in Section 5, where we define our new geodesic distance, as it yields uniform coercivity of the metric tensor in these coordinates.
We first recall some results on RKHS before moving to the analysis of the completion of the space $\mathcal{M}_{0}(X)$ endowed with $\mathbf{g}_{\mu}$, and then to our global change of coordinates.
4.1 A primer on RKHS, and the operators $K_{\mu}$ and $H_{\mu}$ in the RKHS $\mathcal{H}_{\mu}$
Take $k:X\times X\to[0,\infty)$ a non-negative continuous and universal kernel on $X$. In the sequel we will consider $k_{c}=\exp(-c/\varepsilon)$ or $k_{\mu}=\exp(({f_{\mu,\mu}}\oplus{f_{\mu,\mu}}-c)/\varepsilon)$.
We recall that there exists a unique Reproducing Kernel Hilbert Space (RKHS) induced by the kernel $k$ [7, 52]. It is a subset of $\mathcal{C}(X)$ defined as the completion of the linear span of $\left\{k(x,\cdot)\,\middle|\,x\in X\right\}$ in the norm induced by the inner product
$\left\langle\sum_{i=1}^{n}a_{i}k(x_{i},\cdot),\sum_{j=1}^{m}b_{j}k(y_{j},\cdot)\right\rangle_{\mathcal{H}_{k}}:=\sum_{i=1}^{n}\sum_{j=1}^{m}a_{i}b_{j}k(x_{i},y_{j}).$ (23)
Here $n,m\in\mathbb{N}_{0}$, $a_{i},b_{j}\in\mathbb{R}$ and $x_{i},y_{j}\in X$. We write $\mathcal{H}_{k}$ for this Hilbert space, and use $\mathcal{H}_{c}$ for the kernel $k_{c}$ and $\mathcal{H}_{\mu}$ for the kernel $k_{\mu}$. By our assumptions on $c$ both $k_{c}$ and $k_{\mu}$ have the required properties (recall that a kernel $k$ is universal provided $\mathcal{H}_{k}$ is dense in $\mathcal{C}(X)$ in supremum norm). Moreover, when $k$ is bounded, the embedding of $\mathcal{H}_{k}$ into $\mathcal{C}(X)$ is continuous. The evaluation operator is continuous in $\mathcal{H}_{k}$: for any $\phi\in\mathcal{H}_{k}$ there holds $\phi(x)=\langle\phi,k(x,\cdot)\rangle_{\mathcal{H}_{k}}$. More details can be found in Appendix B.
If $\sigma\in\mathcal{H}_{k}^{*}$ is a bounded linear functional on $\mathcal{H}_{k}$, we denote by $H_{k}[\sigma]\in\mathcal{H}_{k}$ its representative given by the Riesz Theorem. By definition, for any $\phi\in\mathcal{H}_{k}$ it satisfies
$\langle\sigma,\phi\rangle=\langle H_{k}[\sigma],\phi\rangle_{\mathcal{H}_{k}}.$ (24)
Evaluating this formula on $\phi=k(y,\cdot)$ we see that the function $H_{k}[\sigma]$ is defined by
$H_{k}[\sigma](y)=\langle\sigma,k(y,\cdot)\rangle.$ (25)
In particular, for $k=k_{\mu}$ the notation is consistent: $H_{\mu}$, introduced in Definition 3.2, can be extended into an operator defined on $\mathcal{H}_{\mu}^{*}$ with values in $\mathcal{H}_{\mu}$. The space $\mathcal{H}_{\mu}^{*}$ contains $\mathcal{M}(X)$ as $\mathcal{H}_{\mu}\subset\mathcal{C}(X)$. In addition, we deduce that the operator $K_{\mu}$, defined by $H_{\mu}[\phi\mu]$ for any $\phi\in\mathcal{C}(X)$, actually takes values in $\mathcal{H}_{\mu}$. We also recall the identity
$\|H_{k}[\sigma]\|^{2}_{\mathcal{H}_{k}}=\iint_{X\times X}k(x,y)\,\mathrm{d}\sigma(x)\,\mathrm{d}\sigma(y)$
valid for all $\sigma\in\mathcal{M}(X)$, which is usually referred to as a Maximum Discrepancy Distance on the space $\mathcal{M}(X)$.
Proposition 4.2.
The operators $H_{\mu}\colon\mathcal{M}(X)\to\mathcal{H}_{\mu}$ and $K_{\mu}\colon\mathcal{C}(X)\to\mathcal{H}_{\mu}$ are compact. Furthermore $K_{\mu}\colon\mathcal{H}_{\mu}\to\mathcal{H}_{\mu}$ is self-adjoint. In particular, it has an eigendecomposition. The eigenvalue $1$ has multiplicity $1$ and all the other eigenvalues belong to $[0,q]$ with $q<1$ as in Proposition 3.7.
Proof.
Compactness of $H_{\mu}$ follows from Lemma B.2. Compactness of $K_{\mu}$ follows directly.
The self-adjointness of $K_{\mu}$ is straightforward, as for all $\phi,\psi\in\mathcal{H}_{\mu}$ it holds
$\langle K_{\mu}[\phi],\psi\rangle_{\mathcal{H}_{\mu}}=\langle H_{\mu}[\phi\mu],\psi\rangle_{\mathcal{H}_{\mu}}=\langle\phi,\psi\rangle_{L^{2}(X,\mu)}=\langle\phi,K_{\mu}[\psi]\rangle_{\mathcal{H}_{\mu}}.$ (26)
As $\mathcal{H}_{\mu}$ injects continuously into $\mathcal{C}(X)$, the operator $K_{\mu}$ is also compact from $\mathcal{H}_{\mu}$ into $\mathcal{H}_{\mu}$. Thus the spectral theorem [12, Theorem 6.11] applies, giving the eigendecomposition.
We now analyze the spectrum of $K_{\mu}$. The computation $\langle K_{\mu}[\phi],\psi\rangle_{\mathcal{H}_{\mu}}=\langle\phi,\psi\rangle_{L^{2}(X,\mu)}$ from (26) shows that $K_{\mu}$ is non-negative, so all its eigenvalues are non-negative. Now, if $\phi\in\mathcal{H}_{\mu}\setminus\{0\}$, $\lambda\in[0,\infty)$ with $K_{\mu}[\phi]=\lambda\phi$, we see that either $\lVert\phi\rVert_{\mathcal{C}(X)/\mathbb{R}}=0$ so that $\phi$ is constant and thus $\lambda=1$, or $\lVert K_{\mu}[\phi]\rVert_{\mathcal{C}(X)/\mathbb{R}}=\lvert\lambda\rvert\cdot\lVert\phi\rVert_{\mathcal{C}(X)/\mathbb{R}}$ so that $\lambda\in[0,q]$ by Proposition 3.7. ∎
Based on these observations we now prove a strengthening of Theorem 3.8. As $\mathbb{1}_{X}=H_{\mu}[\mu]\in\mathcal{H}_{\mu}$, we can consider $\mathcal{H}_{\mu}/\mathbb{R}$ the space $\mathcal{H}_{\mu}$ quotiented by constant functions, equipped with the quotient norm $\lVert\phi\rVert_{\mathcal{H}_{\mu}/\mathbb{R}}=\inf_{\lambda\in\mathbb{R}}\lVert\phi-\lambda\mathbb{1}_{X}\rVert_{\mathcal{H}_{\mu}}$. The Banach space $\mathcal{H}_{\mu}/\mathbb{R}$ is isometric to $\{\mathbb{1}_{X}\}^{\perp}\subseteq\mathcal{H}_{\mu}$ and thus canonically is a Hilbert space.
Theorem 4.3.
Let $m\geq 1$. On the Hilbert space $\mathcal{H}_{\mu}/\mathbb{R}$ equipped with the quotient norm the operators $(\operatorname{id}-K_{\mu}^{m})^{-1}$ are continuous and satisfy
$\operatorname{id}\leq(\operatorname{id}-K_{\mu}^{m})^{-1}\leq\frac{1}{1-q^{m}}\operatorname{id}$
in the partial ordering of self-adjoint operators, with $q<1$ as in Proposition 3.7. The same conclusion holds in the space $L^{2}(X,\mu)/\mathbb{R}$.
These operators appear in the metric tensor $\mathbf{g}_{\mu}$ for $m=2$ and in $\tilde{\mathbf{g}}_{\mu}$ for $m=1$.
Proof.
Note that $\mathcal{H}_{\mu}/\mathbb{R}$ is isometric to the subspace orthogonal to the eigenspace for the eigenvalue $1$. So in this space $K_{\mu}$ is self-adjoint and compact with its spectrum being contained in $[0,q]$. The estimate directly follows. Exactly the same arguments carry through to $L^{2}(X,\mu)/\mathbb{R}$. ∎
4.2 The metric tensor in the RKHS $\mathcal{H}_{\mu}$
Recall that perturbations integrate to zero against $\mathbb{1}_{X}$ because of mass conservation, something that we also need to take in account here. For the sake of clarity we state the following result without proof, which would work for any Banach space of functions on $X$ containing $\mathbb{1}_{X}$.
Lemma 4.4.
The dual of $\mathcal{H}_{\mu}/\mathbb{R}$ is $\mathcal{H}_{\mu,0}^{*}$ the space of bounded linear functionals $\sigma$ on $\mathcal{H}_{\mu}$ such that $\langle\sigma,\mathbb{1}_{X}\rangle=0$.
In particular, $H_{\mu}$ extends to an operator
$H_{\mu}\colon\mathcal{H}_{\mu,0}^{*}\to\mathcal{H}_{\mu}/\mathbb{R},$
which still corresponds to mapping $\sigma\in\mathcal{H}_{\mu,0}^{*}$ to its representative given by the Riesz Theorem: $\langle\sigma,\phi\rangle=\langle H_{\mu}[\sigma],\phi\rangle_{\mathcal{H}_{\mu}/\mathbb{R}}$ for any $\phi\in\mathcal{H}_{\mu}/\mathbb{R}$. From here we can rewrite the metric tensor as follows: If $\dot{\mu}_{1},\dot{\mu}_{2}\in\mathcal{M}_{0}(X)$, we see that $H_{\mu}[\dot{\mu}_{1}]$, $H_{\mu}[\dot{\mu}_{2}]\in\mathcal{H}_{\mu}/\mathbb{R}$. As $(\operatorname{id}-K_{\mu}^{2})^{-1}$ is a bounded operator on this space, we have
$\mathbf{g}_{\mu}(\dot{\mu}_{1},\dot{\mu}_{2})=\frac{\varepsilon}{2}\left\langle\dot{\mu}_{1},(\operatorname{id}-K_{\mu}^{2})^{-1}H_{\mu}[\dot{\mu}_{2}]\right\rangle=\frac{\varepsilon}{2}\left\langle H_{\mu}[\dot{\mu}_{1}],(\operatorname{id}-K_{\mu}^{2})^{-1}H_{\mu}[\dot{\mu}_{2}]\right\rangle_{\mathcal{H}_{\mu}/\mathbb{R}}.$ (27)
This formula directly characterizes the completion of the space $\mathcal{M}_{0}(X)$ endowed with $\mathbf{g}_{\mu}$.
Theorem 4.5.
On $\mathcal{M}_{0}(X)$, the quadratic form $\mathbf{g}_{\mu}$ is positive definite. The completion of $\mathcal{M}_{0}(X)$ with respect to this quadratic form coincides with $\mathcal{H}_{\mu,0}^{*}$, the space of bounded linear functionals $\sigma$ on $\mathcal{H}_{\mu}$ such that $\langle\sigma,\mathbb{1}_{X}\rangle=0$. Moreover, endowing the latter space with its operator norm, for any $\dot{\mu}\in\mathcal{H}_{\mu,0}^{*}$ one has
$\frac{\varepsilon}{2}\lVert\dot{\mu}\rVert^{2}_{\mathcal{H}_{\mu,0}^{*}}\leq\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})\leq\frac{\varepsilon}{2}\frac{1}{1-q^{2}}\lVert\dot{\mu}\rVert^{2}_{\mathcal{H}_{\mu,0}^{*}},$ (28)
with $q$ as in Proposition 3.7.
Proof.
We start from the formula (27), which is a rewriting of Definition 4.1 in the language of RKHS. Combining it with Theorem 4.3 we obtain
$\frac{\varepsilon}{2}\lVert H_{\mu}[\dot{\mu}]\rVert^{2}_{\mathcal{H}_{\mu}/\mathbb{R}}\leq\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})\leq\frac{\varepsilon}{2}\frac{1}{1-q^{2}}\lVert H_{\mu}[\dot{\mu}]\rVert^{2}_{\mathcal{H}_{\mu}/\mathbb{R}},$
at least for $\dot{\mu}\in\mathcal{M}_{0}(X)$. As $H_{\mu}$ is an isometry from $\mathcal{H}_{\mu,0}^{*}$ to the space $\mathcal{H}_{\mu}/\mathbb{R}$, we deduce that (28) is valid for $\dot{\mu}\in\mathcal{M}_{0}(X)$. Thus $\mathbf{g}_{\mu}$ is positive definite.
It remains to show that the completion of $\mathcal{M}_{0}(X)$ is $\mathcal{H}^{*}_{\mu,0}$. Given what we just proved, it is enough to prove that $\mathcal{M}_{0}(X)$ is dense in $\mathcal{H}^{*}_{\mu,0}$ for the norm topology. We use the isometry of the latter space to $\mathcal{H}_{\mu}/\mathbb{R}$ via $\mathcal{H}_{\mu}$. The set $H_{\mu}[\mathcal{M}(X)]$ is dense in $\mathcal{H}_{\mu}$, as discrete measures already yield all linear combinations of $\left\{k_{\mu}(x,\cdot)\,\middle|\,x\in X\right\}$. To conclude, we observe that $\mathcal{M}(X)=\mathbb{R}\cdot\mu\oplus\mathcal{M}_{0}(X)$ and $H_{\mu}[\mu]=\mathbb{1}_{X}$. ∎
Remark 4.6 (Class of admissible perturbations).
If $\dot{\mu}$ is a balanced measure, the norm in $\mathcal{H}_{\mu,0}^{*}$ reads
$\lVert\dot{\mu}\rVert^{2}_{\mathcal{H}_{\mu,0}^{*}}=\langle\dot{\mu},H_{\mu}[\dot{\mu}]\rangle=\iint_{X\times X}\exp\left(\frac{1}{\varepsilon}\left({f_{\mu,\mu}}(x)+{f_{\mu,\mu}}(y)-c(x,y)\right)\right)\,\mathrm{d}\dot{\mu}(x)\,\mathrm{d}\dot{\mu}(y).$
If $X$ is the closure of an open bounded set of $\mathbb{R}^{d}$ and $c$ is a $\mathcal{C}^{(m,m)}$ cost functions on $X\times X$, meaning that mixed partial derivatives up to order $m$ on $x$ and $y$ separately are continuous, then the Schrödinger potentials ${f_{\mu,\mu}}$ are themselves of class $\mathcal{C}^{m}$ on $X$ [14, Proposition 2.2]. Thus, from the equation above, we see that $\mathcal{H}_{\mu,0}^{*}$ contains all balanced distribution of order $m$. For instance, for any cost of class $\mathcal{C}^{2}$ (or more generally $\mathcal{C}^{(1,1)}$) any “vertical perturbation” $\dot{\mu}=\nu$ for $\nu\in\mathcal{M}_{0}(X)$ (here not necessarily with $\nu\ll\mu$) and any “horizontal perturbation” $\dot{\mu}=-\operatorname{div}(v\mu)$ for $v\in L^{1}(X,\mu;\mathbb{R}^{d})$ belong to $\mathcal{H}_{\mu,0}^{*}$.
4.3 The metric tensor in the RKHS $\mathcal{H}_{c}$
Although Theorem 4.5 exhibits a very neat structure, it has the main drawback that the ambient space of the tangent space $\mathcal{H}_{\mu,0}^{*}\sim\mathcal{H}_{\mu}/\mathbb{R}$ depends on the point $\mu$. The space $\mathcal{H}_{\mu}$ is however in a simple bijection with the RKHS $\mathcal{H}_{c}$, built on the kernel $k_{c}=\exp(-c/\varepsilon)$, as the next proposition shows. In the following we will abbreviate
$a_{\mu}:=\exp({f_{\mu,\mu}}/\varepsilon).$
Proposition 4.7.
The spaces $\mathcal{H}_{c}$ and $\mathcal{H}_{\mu}$ are isometric to one another under the map
$\iota_{\mu}\colon\mathcal{H}_{c}\to\mathcal{H}_{\mu},\quad\phi\mapsto\phi\cdot\exp\left(\frac{1}{\varepsilon}{f_{\mu,\mu}}\right)=\phi\cdot a_{\mu}.$
Proof.
We can directly verify $\exp({f_{\mu,\mu}}(y)/\varepsilon)k_{c}(x,y)=\exp(-{f_{\mu,\mu}}(x)/\varepsilon)k_{\mu}(x,y)$ and hence
$\left\lVert\exp\left(\frac{1}{\varepsilon}{f_{\mu,\mu}}\right)\sum_{i=1}^{n}b_{i}k_{c}(x_{i},\cdot)\right\rVert_{\mathcal{H}_{\mu}}^{2}$ $=\sum_{i,j=1}^{n}b_{i}b_{j}\exp\left(-\frac{1}{\varepsilon}{f_{\mu,\mu}}(x_{i})\right)\exp\left(-\frac{1}{\varepsilon}{f_{\mu,\mu}}(x_{j})\right)k_{\mu}(x_{i},x_{j})$
$=\sum_{i,j=1}^{n}b_{i}b_{j}k_{c}(x_{i},x_{j})=\left\lVert\sum_{i=1}^{n}b_{i}k_{c}(x_{i},\cdot)\right\rVert_{\mathcal{H}_{c}}^{2}.$
We consider the maps $A\colon\mathcal{P}(X)\to\mathcal{M}(X)$ and $B=H_{c}\circ A\colon\mathcal{P}(X)\to\mathcal{H}_{c}$ given by
$\begin{split}&A(\mu)=\exp\left(\frac{1}{\varepsilon}{f_{\mu,\mu}}\right)\mu,\\ &B(\mu)=H_{c}[A(\mu)]=\int_{X}\exp\left(\frac{1}{\varepsilon}({f_{\mu,\mu}}(y)-c(\cdot,y))\right)\,\mathrm{d}\mu(y)=\exp\left(-\frac{1}{\varepsilon}{f_{\mu,\mu}}\right).\end{split}$
Recall that the RKHS embeddings $H_{c}\colon\mathcal{H}_{c}^{*}\to\mathcal{H}_{c}$ and $H_{\mu}\colon\mathcal{H}_{\mu}^{*}\to\mathcal{H}_{\mu}$ are characterized by (24) and (25). From a path $(\mu_{t})_{t}\in\mathcal{P}(X)$ one obtains the paths $(H_{\mu}[\mu_{t}])_{t}\in\mathcal{H}_{\mu}$, $(\alpha_{t})_{t}\in\mathcal{M}_{+}(X)$, and $(\beta_{t})_{t}\in\mathcal{H}_{c}$ with $\alpha_{t}=A(\mu_{t})$, $\beta_{t}=B(\mu_{t})=H_{c}[A(\mu_{t})]$. The change of variables $\mu\leftrightarrow\alpha$ was already used by Feydy et al. in the proof of [26, Prop. 3]. Note that, for a fix measure $\nu$, the space $\mathcal{P}(X)$ embeds into $\mathcal{H}_{\nu}$ via $H_{\nu}$, and that the latter can be mapped into $\mathcal{H}_{c}$ with $\iota_{\nu}^{-1}$. However, this embedding $\mu\mapsto\iota_{\nu}^{-1}(H_{\nu}[\mu])=H_{c}[a_{\nu}\cdot\mu]$ of $\mathcal{P}(X)$ into $\mathcal{H}_{c}$ differs from $B:\mu\mapsto H_{c}[a_{\mu}\cdot\mu]$.
The measures $\alpha_{t}$ are naturally interpreted as linear functionals on the common space $\mathcal{H}_{c}$ (represented by $\beta_{t}$). One can hope that differentiability on this fixed space is well behaved. We will see that the metric tensor can be expressed in the $\beta$ variable and that it is equivalent to the squared norm on the space $\mathcal{H}_{c}$ in Theorem 4.11.
We first prove that $\mu\leftrightarrow\alpha$ and $\mu\leftrightarrow\beta$ are valid changes of variables in the sense that $A$ and $B$ are homeomorphisms onto their images. Recall that $\operatorname{P}_{1}$ stands for the projection map $(x,y)\mapsto x$.
Theorem 4.8.
The maps $A$ and $B$ are homemorphisms onto their images. Explicitly:
  1. 1.
    $A$ is a weak-$\ast$ homeomorphism onto $\mathcal{A}:=\left\{\alpha\in\mathcal{M}_{+}(X)\,\middle|\,\lVert H_{c}[\alpha]\rVert_{\mathcal{H}_{c}}=1\right\}$ with inverse $A^{-1}(\alpha)=(\operatorname{P}_{1})_{\#}(k_{c}\cdot\alpha\otimes\alpha)$. In particular, $\mathcal{A}$ is weak-$\ast$-compact.
  2. 2.
    $B$ is a weak-$\ast$-to-weak homeomorphism onto :=$\mathcal{B}~{}:=~{}\left\{\beta~{}\in~{}H_{c}[\mathcal{M}_{+}(X)]\,\middle|\,\lVert\beta\rVert_{\mathcal{H}_{c}}=1\right\}$. Weak convergence and norm convergence agree on $\mathcal{B}$, and the set $\mathcal{B}$ is weakly and norm compact in $\mathcal{H}_{c}$.
Note that $\mathcal{B}$ is the intersection of the convex cone $H_{c}[\mathcal{M}_{+}(X)]$ with the unit sphere. The proof for the weak compactness of $\mathcal{B}\subset\mathcal{H}_{c}$ relies heavily on the positivity constraint of the cone $H_{c}[\mathcal{M}_{+}(X)]$. In contrast, the weak closure of the whole unit sphere in infinite dimension is the unit ball, which is not norm compact.
Proof.
The map $A$ and its image: Denote $\alpha=A(\mu)=a_{\mu}\mu$. Note that the map $A$ is injective since $\mu$ can be reconstructed as $\mu=F(\alpha):=(\operatorname{P}_{1})_{\#}(k_{c}\cdot\alpha\otimes\alpha)$ thanks to (16). This also implies that the inverse map is weak-$\ast$-continuous: let $\alpha_{n}\to\alpha$ weak-$\ast$ with corresponding $\mu_{n}=(\operatorname{P}_{1})_{\#}(k_{c}\cdot\alpha_{n}\otimes\alpha_{n})$. Then $\alpha_{n}\otimes\alpha_{n}$ converges weak-$\ast$ to $\alpha\otimes\alpha$ (see e.g. [57, Lem. 7.3]). Weak convergence of $\mu_{n}$ to $\mu$ easily follows. To prove that the forward map is continuous, it suffices to use the continuity of ${f_{\mu,\mu}}$ as a function of $\mu$ in supremum norm from [14, Thm. 1.1].
We now characterize the image of $A$: Clearly $A(\mu)\geq 0$ and
$\lVert H_{c}[A(\mu)]\rVert_{\mathcal{H}_{c}}^{2}=\langle a_{\mu}\mu\otimes a_{\mu}\mu,k_{c}\rangle=\langle\mu\otimes\mu,k_{\mu}\rangle=1$
for all $\mu\in\mathcal{P}(X)$. The map $F\colon\mathcal{A}\to\mathcal{P}(X)$ is a left inverse of $A\colon\mathcal{P}(X)\to\mathcal{A}$. To show that $F$ is also a right inverse, we may equivalently show that $F$ is injective. We have $F(\alpha)=\mu$ if and only if $k_{c}\cdot\alpha\otimes\alpha\in\Pi(\mu,\mu)$. For every such $\alpha\in\mathcal{A}$ it must hold that $\alpha\ll\mu$. Thus $k_{c}\cdot\alpha\otimes\alpha=\exp(\phi\oplus\phi)k_{c}\cdot\mu\otimes\mu$ for $\phi=\log\frac{\mathrm{d}\alpha}{\mathrm{d}\mu}\colon X\to[-\infty,\infty)$. A transport plan of this form is necessarily optimal for the entropic optimal transport problem (1) (see [49, Thm. 2.1]). Hence $\phi={f_{\mu,\mu}}/\varepsilon$ and $\alpha=A(\mu)$. Finally, $\mathcal{A}$ is weak-$\ast$-compact as the weak-$\ast$-continuous image of the weak-$\ast$-compact set $\mathcal{P}(X)$.
The map $B$ and its image: The map $A$ is injective. The map $H_{c}$ is a bijection between $\mathcal{H}_{c}^{*}$ and $\mathcal{H}_{c}$ by definition, so it is injective. Thus $B$ is injective. It is straightforward that $\mathcal{B}$ is the image of $\mathcal{A}$ by $H_{c}$. Thus $\mathcal{B}$ is the image of $\mathcal{P}(X)$ by $B$.
As the injection of $\mathcal{H}_{c}$ into $\mathcal{C}(X)$ is continuous by (53), the map $H_{c}$ is continuous from $\mathcal{M}(X)$ endowed with the weak-$\ast$ topology into $\mathcal{H}_{c}$ endowed with the weak topology. Thus $B$ is weak-$\ast$-to-weak continuous and $\mathcal{B}$ is weakly compact as the image of the weak-$\ast$-compact set $\mathcal{A}$. As a continuous bijection between compact Hausdorff spaces, $B$ is a weak-$\ast$-to-weak homeomorphism.
Finally, if $(\beta_{n})_{n}$ is any sequence in $\mathcal{B}$ which converges weakly to a limit $\beta$ in $\mathcal{H}_{c}$, as $\mathcal{B}$ is weakly compact we must have $\beta\in\mathcal{B}$ so $\|\beta\|_{\mathcal{H}_{c}}=1=\lim_{n\to\infty}\|\beta_{n}\|_{\mathcal{H}_{c}}$. This implies norm convergence of the sequence $(\beta_{n})_{n}$. ∎
After the investigation of the topological properties of these changes of variables, we move on to the differential properties. For a path $(\mu_{t})_{t}$ in $\mathcal{P}(X)$, we denote $\alpha_{t}=A(\mu_{t})$, $\beta_{t}=B(\mu_{t})$. By definition this means $\langle\alpha_{t},\phi\rangle=\langle\beta_{t},\phi\rangle_{\mathcal{H}_{c}}$ for all $\phi\in\mathcal{H}_{c}$. We will express the metric tensor in the variable $\dot{\beta}_{t}:=\frac{\mathrm{d}}{\mathrm{d}t}\beta_{t}=\frac{\mathrm{d}}{\mathrm{d}t}\exp(-f_{\mu_{t},\mu_{t}}/\varepsilon)\in\mathcal{H}_{c}$. Let us first define the push-forward of the metric tensor before making a rigorous link to Definition 4.1 in Theorem 4.11.
Definition 4.9.
If $\mu\in\mathcal{P}(X)$ with $\beta=B(\mu)\in\mathcal{B}$, and $\dot{\beta}\in\mathcal{H}_{c}$ with $\langle\beta,\dot{\beta}\rangle_{\mathcal{H}_{c}}=0$, we define the quadratic form $\tilde{\mathbf{g}}_{\mu}(\dot{\beta},\dot{\beta})$ by
$\tilde{\mathbf{g}}_{\mu}(\dot{\beta},\dot{\beta})=\frac{\varepsilon}{2}\left(\lVert\dot{\beta}\rVert_{\mathcal{H}_{c}}^{2}+2\left\langle a_{\mu}\cdot\dot{\beta},(\operatorname{id}-K_{\mu})^{-1}[a_{\mu}\cdot\dot{\beta}]\right\rangle_{L^{2}(X,\mu)}\right).$
To explicitly link $\mathbf{g}$ and $\tilde{\mathbf{g}}$, we need to consider curves valued in $\mathcal{P}(X)$ and their corresponding curves in $\mathcal{B}$. The following definition fixes the convention about what we mean by a differentiable path.
Definition 4.10 (Differentiability of paths).
Let $\mathbb{X}$ be a vector space with a topology $\tau$. We say that a path $(x_{s})_{s\in I}$ in $\mathbb{X}$ is $\tau$-differentiable at $t$ when $t$ is an accumulation point of $I$ and difference quotients $(x_{s}-x_{t})/(s-t)$ converge in $\tau$ to some $\dot{x}_{t}$ as $s\to t$. When $\mathbb{X}$ is a normed vector space and $\tau$ is the norm topology, we drop the $\tau$ in the notation.
We define $\dot{\mu}_{t}$ by $\langle\dot{\mu}_{t},\phi\rangle=\frac{\mathrm{d}}{\mathrm{d}s}\big{|}_{s=t}\langle\mu_{s},\phi\rangle$ for $\phi\in\mathcal{H}_{\mu_{t}}$ whenever this is a bounded linear functional on $\mathcal{H}_{\mu_{t}}$, i.e. when the path $(\mu_{s})_{s}$, interpreted in the dual space $\mathcal{H}_{\mu_{t}}^{*}$, is weakly differentiable in $\mathcal{H}_{\mu_{t}}^{*}$ at $t$. Note that automatically $\dot{\mu}_{t}\in\mathcal{H}_{\mu_{t},0}^{*}$.
In the sequel, for paths $(\mu_{t})_{t}$, $(\alpha_{t})_{t}$, $(\beta_{t})_{t}$ we write $\mu=\mu_{0}$, $\dot{\mu}=\dot{\mu}_{0}$ and similarly $\alpha=\alpha_{0}$, $\dot{\alpha}=\dot{\alpha}_{0}$ and $\beta=\beta_{0}$, $\dot{\beta}=\dot{\beta}_{0}$.
Theorem 4.11.
Let $(\mu_{t})_{t}$ be a path in $\mathcal{P}(X)$, such that the associated path $(\beta_{t})_{t}=(B(\mu_{t}))_{t}$ is weakly differentiable in $\mathcal{H}_{c}$ at $t=0$. Then $\langle\dot{\beta},\beta\rangle_{\mathcal{H}_{c}}=0$, the path $(\mu_{t})_{t}$ is weakly differentiable in $\mathcal{H}_{\mu}^{*}$ at $t=0$, and
$\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})=\tilde{\mathbf{g}}_{\mu}(\dot{\beta},\dot{\beta}).$ (29)
To prove this theorem we first prove a preliminary lemma relating $\dot{\beta}$ and $\dot{\mu}$. Note that since $\beta_{t}=\exp(f_{\mu_{t},\mu_{t}}/\varepsilon)$, (30) is formally equivalent to the last equation of Proposition 3.11. The differences are the regularity assumptions and the functional spaces involved.
Lemma 4.12.
Let $(\mu_{t})_{t}$ be a path in $\mathcal{P}(X)$, such that the corresponding path $(\beta_{t})_{t}$ is continuous and weakly differentiable in $\mathcal{H}_{c}$ at $t=0$. Then $(\mu_{t})_{t}$ is weakly differentiable in $\mathcal{H}_{\mu}^{*}$ at $t=0$ with
$H_{\mu}[\dot{\mu}]=(\operatorname{id}+K_{\mu})[a_{\mu}\cdot\dot{\beta}].$ (30)
Explicitly, this means that for all $\phi\in\mathcal{H}_{\mu}$ it holds
$\frac{\mathrm{d}}{\mathrm{d}t}\Big{|}_{t=0}\left\langle\mu_{t},\phi\right\rangle=\left\langle(\operatorname{id}+K_{\mu})[a_{\mu}\dot{\beta}],\phi\right\rangle_{\mathcal{H}_{\mu}}=\left\langle\dot{\beta},a_{\mu}^{-1}\cdot(\operatorname{id}+K_{\mu})[\phi]\right\rangle_{\mathcal{H}_{c}}.$
Proof.
By Theorem 4.8, the paths $(\mu_{t})_{t}$ and $(\alpha_{t})_{t}$ are weak-$\ast$-continuous at $t=0$.
Recall that we have $\mu_{t}=\alpha_{t}\exp(-f_{\mu_{t},\mu_{t}}/\varepsilon)$ and $\beta_{t}=\exp(-f_{\mu_{t},\mu_{t}}/\varepsilon)$, thus $\mu_{t}=\alpha_{t}\beta_{t}$. We will apply Leibniz’s rule to this formula, but we rederive it given the low regularity of the objects involved. Indeed, as measures on $X$, we can write
$\frac{\mu_{t}-\mu}{t}=\frac{\alpha_{t}\beta_{t}-\alpha\beta}{t}=\frac{\alpha_{t}-\alpha}{t}\beta+\alpha_{t}\frac{\beta_{t}-\beta}{t}.$
We fix $\phi\in\mathcal{H}_{\mu}$ and pair it with the expression above. Using $\beta=a_{\mu}^{-1}$, we find
$\begin{split}\left\langle\frac{\mu_{t}-\mu}{t},\phi\right\rangle&=\left\langle\frac{\alpha_{t}-\alpha}{t},a_{\mu}^{-1}\phi\right\rangle+\left\langle\frac{\beta_{t}-\beta}{t},\alpha_{t}\phi\right\rangle\\ &=\left\langle\frac{\beta_{t}-\beta}{t},a_{\mu}^{-1}\phi\right\rangle_{\mathcal{H}_{c}}+\left\langle\frac{\beta_{t}-\beta}{t},H_{c}[\alpha_{t}\phi]\right\rangle_{\mathcal{H}_{c}}.\end{split}$ (31)
For the first summand we use that $a_{\mu}^{-1}\phi\in\mathcal{H}_{c}$ and the weak differentiability of $(\beta_{t})$. For the second summand we use that $H_{c}[\phi\cdot\alpha_{t}]$ converges in norm in $\mathcal{H}_{c}$ by Lemma B.2, while $(\beta_{t})$ is weakly differentiable. We thus find
$\frac{\mathrm{d}}{\mathrm{d}t}\Big{|}_{t=0}\left\langle\mu_{t},\phi\right\rangle=\left\langle\dot{\beta},a_{\mu}^{-1}\phi\right\rangle_{\mathcal{H}_{c}}+\left\langle\dot{\beta},H_{c}[\alpha\phi]\right\rangle_{\mathcal{H}_{c}}.$
The proof is concluded by rewriting this expression. From Proposition 4.7 we have $\langle\dot{\beta},a_{\mu}^{-1}\phi\rangle_{\mathcal{H}_{c}}=\langle a_{\mu}\dot{\beta},\phi\rangle_{\mathcal{H}_{\mu}}$, while for the second term
$\left\langle\dot{\beta},H_{c}[\alpha\phi]\right\rangle_{\mathcal{H}_{c}}=\left\langle a_{\mu}\dot{\beta},a_{\mu}H_{c}[a_{\mu}\mu\phi]\right\rangle_{\mathcal{H}_{\mu}}=\left\langle a_{\mu}\dot{\beta},K_{\mu}[\phi]\right\rangle_{\mathcal{H}_{\mu}}.$
The conclusion follows as $K_{\mu}$ is self-adjoint on $\mathcal{H}_{\mu}$ by Proposition 4.2. ∎
Proof of Theorem 4.11.
The equation $\langle\dot{\beta},\beta\rangle_{\mathcal{H}_{c}}=0$ is directly obtained by differentiating the equation $\|\beta_{t}\|_{\mathcal{H}_{c}}^{2}=1$ in time.
We then plug the formula for $\dot{\mu}$ from Lemma 4.12 into the metric tensor
$\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})=\frac{\varepsilon}{2}\left\langle\dot{\mu},(\operatorname{id}-K_{\mu}^{2})^{-1}H_{\mu}[\dot{\mu}]\right\rangle=\frac{\varepsilon}{2}\left\langle(\operatorname{id}+K_{\mu})[a_{\mu}\cdot\dot{\beta}],(\operatorname{id}-K_{\mu}^{2})^{-1}(\operatorname{id}+K_{\mu})[a_{\mu}\cdot\dot{\beta}]\right\rangle_{\mathcal{H}_{\mu}}.$
The operator $(\operatorname{id}+K_{\mu})$ is self-adjoint on $\mathcal{H}_{\mu}$. Furthermore, $(\operatorname{id}+K_{\mu})(\operatorname{id}-K_{\mu}^{2})^{-1}(\operatorname{id}+K_{\mu})=(\operatorname{id}+K_{\mu})(\operatorname{id}-K_{\mu})^{-1}=\operatorname{id}+2K_{\mu}(\operatorname{id}-K_{\mu})^{-1}$. To conclude, use $\langle\phi,K_{\mu}[\psi]\rangle_{\mathcal{H}_{\mu}}=\langle\phi,\psi\rangle_{L^{2}(X,\mu)}$ and Proposition 4.7. ∎
We finish this section by studying the continuity properties of the metric tensor $\tilde{\mathbf{g}}$, and by proving that it can be uniformly compared to the norm of $\mathcal{H}_{c}$.
Proposition 4.13.
The metric tensor $\tilde{\mathbf{g}}_{\mu}(\dot{\beta},\dot{\beta})$ is jointly continuous (resp. lower semicontinuous) with respect to weak-$\ast$ convergence of $\mu$ in $\mathcal{P}(X)$ and norm convergence (resp. weak convergence) of $\dot{\beta}$ in $\mathcal{H}_{c}$.
Proof.
The first part $\frac{\varepsilon}{2}\lVert\dot{\beta}\rVert_{\mathcal{H}_{c}}^{2}$ of the metric tensor $\tilde{\mathbf{g}}_{\mu}$ does not depend on $\mu$ and is of course continuous (resp. lower semicontinuous) with respect to norm convergence (resp. weak convergence) of $\dot{\beta}$ in $\mathcal{H}_{c}$.
For the second part, we actually claim that the map
$\left\{(\mu,\phi)\in\mathcal{P}(X)\times\mathcal{H}_{c}\,\middle|\,\langle B(\mu),\phi\rangle_{\mathcal{H}_{c}}=0\right\}\to\mathbb{R},\quad(\mu,\phi)\mapsto\left\langle a_{\mu}\cdot\phi,(\operatorname{id}-K_{\mu})^{-1}[a_{\mu}\cdot\phi]\right\rangle_{L^{2}(X,\mu)}$
is jointly continuous with respect to weak-$\ast$ convergence on $\mathcal{P}(X)$ and weak convergence on $\mathcal{H}_{c}$. Indeed, let $\mu_{n}\to\mu$ weak-$\ast$ in $\mathcal{P}(X)$, $\phi_{n}\to\phi$ weakly in $\mathcal{H}_{c}$, and $\langle B(\mu_{n}),\phi_{n}\rangle=0$. In particular, we have uniform convergence $f_{n}=f_{\mu_{n},\mu_{n}}\to{f_{\mu,\mu}}$, hence $a_{\mu_{n}}\to a_{\mu}$, and by Lemma B.1 also $\phi_{n}\to\phi$ uniformly. Let us write
$u_{n}=(\operatorname{id}-K_{\mu_{n}})^{-1}[a_{\mu_{n}}\cdot\phi_{n}]\quad\Leftrightarrow\quad u_{n}=a_{\mu_{n}}\cdot\phi_{n}+K_{\mu_{n}}u_{n}.$
We show that the sequence $u_{n}$ converges to $u=(\operatorname{id}-K_{\mu})^{-1}[a_{\mu}\cdot\phi]$ in $\mathcal{C}(X)/\mathbb{R}$. Indeed, the uniform convergence of $a_{\mu_{n}}\cdot\phi_{n}$ and the operator norm estimate in Proposition 3.7 show that the functions $u_{n}$ are uniformly bounded in $\mathcal{C}(X)/\mathbb{R}$. By [24, Thm. 1.2] the potentials $f_{\mu_{n},\mu_{n}}$ are equi-bounded. They are also equi-continuous, as by [50, Lem. 3.1] they have the same modulus of continuity as the cost function $c$. Thus it is clear that the functions $K_{\mu_{n}}u_{n}$ are equi-continuous and equi-bounded. Hence, up to extraction of a subsequence, $K_{\mu_{n}}u_{n}$ converges uniformly to a limit. Now also $u_{n}=a_{\mu_{n}}\cdot\phi_{n}+K_{\mu_{n}}u_{n}$ converges uniformly to a limit $u$. Once we have this information, we can identify this limit with $(\operatorname{id}-K_{\mu})^{-1}[a_{\mu}\cdot\phi]$, because
$u_{n}=a_{\mu_{n}}\cdot\phi_{n}+K_{\mu_{n}}u_{n}\to a_{\mu}\cdot\phi+K_{\mu}u=u.$
Once we have this convergence,
$\left\langle a_{\mu_{n}}\cdot\phi_{n},(\operatorname{id}-K_{\mu_{n}})^{-1}[a_{\mu_{n}}\cdot\phi_{n}]\right\rangle_{L^{2}(X,\mu_{n})}=\int_{X}a_{\mu_{n}}\phi_{n}u_{n}\,\mathrm{d}\mu_{n},$
and the integrand $a_{\mu_{n}}\phi_{n}u_{n}$ converges uniformly to $a_{\mu}\phi u$, while the measure $\mu_{n}$ converges weak-$\ast$ to $\mu$. It is easy to pass to the limit using Lemma A.2. ∎
Proposition 4.14 (The metric tensor is equivalent to the norm on $\mathcal{H}_{c}$).
Let $\mu\in\mathcal{P}(X)$, $\dot{\beta}\in\mathcal{H}_{c}$ with $\langle B(\mu),\dot{\beta}\rangle_{\mathcal{H}_{c}}=0$. It holds that
$\frac{\varepsilon}{2}\lVert\dot{\beta}\rVert_{\mathcal{H}_{c}}^{2}\leq\tilde{\mathbf{g}}_{\mu}(\dot{\beta},\dot{\beta})\leq C\frac{\varepsilon}{2}\lVert\dot{\beta}\rVert_{\mathcal{H}_{c}}^{2}$
for $C=1+2\exp\left(\frac{11}{2\varepsilon}\lVert c\rVert_{\infty}\right)$.
Proof.
From Theorem 4.3, on the space $L^{2}(X,\mu)/\mathbb{R}$ we have $\operatorname{id}\leq(\operatorname{id}-K_{\mu})^{-1}\leq\frac{1}{1-q}\operatorname{id}$ with $q=1-\exp(-4\lVert c\rVert_{\infty}/\varepsilon)$ as in Proposition 3.7 (and in particular $(\operatorname{id}-K_{\mu})^{-1}$ is non-negative). Thus we get
$\frac{\varepsilon}{2}\lVert\dot{\beta}\rVert_{\mathcal{H}_{c}}^{2}\leq\tilde{\mathbf{g}}_{\mu}(\dot{\beta},\dot{\beta})\leq\frac{\varepsilon}{2}\left(\lVert\dot{\beta}\rVert_{\mathcal{H}_{c}}^{2}+\frac{2}{1-q}\lVert a_{\mu}\cdot\dot{\beta}\rVert^{2}_{L^{2}(X,\mu)/\mathbb{R}}\right).$
We then control the $L^{2}(X,\mu)/\mathbb{R}$ norm by the supremum norm, using that $\lVert\dot{\beta}\rVert_{\infty}\leq\lVert\dot{\beta}\rVert_{\mathcal{H}_{c}}$ as the cost function is non-negative. Thus
$\frac{\varepsilon}{2}\lVert\dot{\beta}\rVert_{\mathcal{H}_{c}}^{2}\leq\tilde{\mathbf{g}}_{\mu}(\dot{\beta},\dot{\beta})\leq\frac{\varepsilon}{2}\left(\lVert\dot{\beta}\rVert_{\mathcal{H}_{c}}^{2}+\frac{2}{1-q}\|a_{\mu}\|_{\infty}\lVert\dot{\beta}\rVert^{2}_{\mathcal{H}_{c}}\right).$
Eventually, using [24, Thm. 1.2] we have $\lVert a_{\mu}\rVert_{\infty}\leq\exp(\lVert{f_{\mu,\mu}}\rVert_{\infty}/\varepsilon)\leq\exp(\frac{3}{2}\lVert c\rVert_{\infty}/\epsilon)$, and together with the definition of $q$ we conclude. ∎
5 The induced path metric
5.1 Definition and metrization of the weak-$\ast$ topology
Definition 5.1.
  1. 1.
    We call a path $(\mu_{t})_{t\in[0,1]}$ admissible when the corresponding path $(\beta_{t})_{t\in(0,1)}=(B(\mu_{t}))_{t\in(0,1)}$ belongs to the Sobolev space $\mathcal{H}^{1}((0,1);\mathcal{H}_{c})$ (see Appendix C). We denote the set of admissible paths as $P$ and write $P(\mu_{0},\mu_{1})$ for the subset of $P$ with fixed start and end points $\mu_{0},\mu_{1}\in\mathcal{P}(X)$.
  2. 2.
    The energy of a path $(\mu_{t})_{t}\in P$ is defined as
    $E((\mu_{t})_{t}):=\int_{0}^{1}\tilde{\mathbf{g}}_{\mu_{t}}(\dot{\beta}_{t},\dot{\beta}_{t})\,\mathrm{d}t=\int_{0}^{1}\mathbf{g}_{\mu_{t}}(\dot{\mu}_{t},\dot{\mu}_{t})\,\mathrm{d}t.$
    (These two integrals indeed coincide if $(\mu_{t})_{t}\in P$ thanks to Theorem 4.11.)
  3. 3.
    For $\mu_{0},\mu_{1}\in\mathcal{P}(X)$ we define
    $\mathsf{d}_{S}(\mu_{0},\mu_{1}):=\left(\inf_{(\mu_{t})_{t}\in P(\mu_{0},\mu_{1})}E((\mu_{t})_{t})\right)^{\frac{1}{2}}.$ (32)
Theorem 4.8 tells us that $P$ corresponds to $\mathcal{H}^{1}((0,1);\mathcal{B})$ and the paths $(\mu_{t})_{t}$ are weak-$\ast$-continuous. If the cost function is of class $\mathcal{C}^{2}$, then we are allowed to move measures by vertical (18) or horizontal (20) perturbations: this was already noticed in Remark 4.6. Lemma B.2 then shows that the corresponding path $(\beta_{t})_{t}$ is differentiable (in norm). Intuitively, combinations of vertical and horizontal perturbations would seem sufficiently rich to describe $\dot{\mu}$ for all relevant trajectories, but we extend admissible paths to $P$ to guarantee completeness. We now show that $\mathsf{d}_{S}$ is a distance (Theorem 5.2) which is geodesic (Theorem 5.3).
Theorem 5.2.
The function $\mathsf{d}_{S}$ is a metric on $\mathcal{P}(X)$ that metrizes the weak-$\ast$ topology. Further, for all $\mu_{0},\mu_{1}\in\mathcal{P}(X)$, $\beta_{0}=B(\mu_{0})$, $\beta_{1}=B(\mu_{1})$ there holds
$\sqrt{\frac{\varepsilon}{2}}\lVert\beta_{0}-\beta_{1}\rVert_{\mathcal{H}_{c}}\leq\mathsf{d}_{S}(\mu_{0},\mu_{1})\leq\sqrt{\frac{\varepsilon}{2}\pi C}\lVert\beta_{0}-\beta_{1}\rVert_{\mathcal{H}_{c}}$
with $C=1+2\exp\left(\frac{11}{2\varepsilon}\lVert c\rVert_{\infty}\right)$ as in Proposition 4.14.
Proof.
It is easy to see that $\mathsf{d}_{S}(\mu_{0},\mu_{1})=\mathsf{d}_{S}(\mu_{1},\mu_{0})\geq 0$. We show the triangle inequality and the relation to the $\mathcal{H}_{c}$ norm. The latter also implies finiteness, positive definiteness, and the metrization of the weak-$\ast$ topology.
The proof of the triangle inequality is analogous to that of [11, Thm. 1]. Let $\mu,\nu,\eta\in\mathcal{P}(X)$, $(\mu_{t})_{t}\in P(\mu,\nu)$, $(\nu_{t})_{t}\in P(\nu,\eta)$ and denote $\beta^{1}_{t}=B(\mu_{t})$, $\beta^{2}_{t}=B(\nu_{t})$. Assume that $\mathsf{d}_{S}(\mu,\nu)>0$ and $\mathsf{d}_{S}(\nu,\eta)>0$. Let $\tau\in(0,1)$ and set $\gamma_{t}=\mu_{t/\tau}$ for $t\in[0,\tau]$, $\gamma_{t}=\nu_{\frac{t-\tau}{1-\tau}}$ for $t\in[\tau,1]$. Using Theorem C.2 and the weak-$\ast$ continuity of $(\gamma_{t})_{t}$, one can see that $(\gamma_{t})_{t}\in P(\mu,\eta)$ and by simple rescaling there holds
$E((\gamma_{t})_{t})$ $=\frac{1}{\tau}\int_{0}^{1}\tilde{\mathbf{g}}_{\mu_{t}}(\dot{\beta}^{1}_{t},\dot{\beta}^{1}_{t})\,\mathrm{d}t+\frac{1}{1-\tau}\int_{0}^{1}\tilde{\mathbf{g}}_{\nu_{t}}(\dot{\beta}^{2}_{t},\dot{\beta}^{2}_{t})\,\mathrm{d}t$
$=\frac{1}{\tau}E((\mu_{t})_{t})+\frac{1}{1-\tau}E((\nu_{t})_{t}).$
Setting $\tau={E((\mu_{t})_{t})^{\frac{1}{2}}}/{(E((\mu_{t})_{t})^{\frac{1}{2}}+E((\nu_{t})_{t})^{\frac{1}{2}})}$ gives
$\frac{1}{\tau}E((\mu_{t})_{t})+\frac{1}{1-\tau}E((\nu_{t})_{t})=(E((\mu_{t})_{t})^{\frac{1}{2}}+E((\nu_{t})_{t})^{\frac{1}{2}})^{2}.$
Taking the infinum over $(\mu_{t})_{t}\in P(\mu,\nu)$ and $(\nu_{t})_{t}\in P(\nu,\eta)$ yields the triangle inequality for $\mathsf{d}_{S}$.
To show the equivalence to the $\mathcal{H}_{c}$-norm, we use that $\frac{\varepsilon}{2}\lVert\dot{\beta}_{t}\rVert_{\mathcal{H}_{c}}^{2}\leq\tilde{\mathbf{g}}_{\mu_{t}}(\dot{\beta}_{t},\dot{\beta}_{t})\leq C\frac{\varepsilon}{2}\lVert\dot{\beta}_{t}\rVert_{\mathcal{H}_{c}}^{2}$ from Proposition 4.14.
For the lower bound, let $(\mu_{t})_{t}\in P(\mu_{0},\mu_{1})$ and denote $\beta_{t}=B(\mu_{t})$. Then Jensen’s inequality, the norm estimate from Theorem C.1, and the fundamental theorem of calculus in $\mathcal{H}^{1}((0,1);\mathcal{H}_{c})$ from Theorem C.2 show
$\frac{2}{\varepsilon}\int_{0}^{1}\tilde{\mathbf{g}}_{\mu_{t}}(\dot{\beta}_{t},\dot{\beta}_{t})\,\mathrm{d}t\geq\int_{0}^{1}\lVert\dot{\beta}_{t}\rVert_{\mathcal{H}_{c}}^{2}\,\mathrm{d}t\geq\left\lVert\int_{0}^{1}\dot{\beta}_{t}\,\mathrm{d}t\right\rVert_{\mathcal{H}_{c}}^{2}=\lVert\beta_{1}-\beta_{0}\rVert_{\mathcal{H}_{c}}^{2}.$
This gives the lower bound when we take the infimum over all curves.
For the upper bound, let $(\beta_{t})_{t}$ be the arc connecting $\beta_{0}$ to $\beta_{1}$ with $\lVert\beta_{t}\rVert_{\mathcal{H}_{c}}=1$ parametrized by arc length of $(\beta_{t})_{t}$ in $\mathcal{H}_{c}$. By Theorem 4.8, this is an admissible path. A basic geometry argument in the two-dimensional Hilbert space $\operatorname{span}\{\beta_{0},\beta_{1}\}$ yields
$\left(\int_{0}^{1}\lVert\dot{\beta}_{t}\rVert_{\mathcal{H}_{c}}^{2}\,\mathrm{d}t\right)^{\frac{1}{2}}=\int_{0}^{1}\lVert\dot{\beta}_{t}\rVert_{\mathcal{H}_{c}}\,\mathrm{d}t=2\arcsin\left(\frac{1}{2}\lVert\beta_{1}-\beta_{0}\rVert_{\mathcal{H}_{c}}\right),$
which is bounded by $\frac{\pi}{2}\lVert\beta_{1}-\beta_{0}\rVert_{\mathcal{H}_{c}}$.
Finally, for $\mu,\mu_{n}\in\mathcal{P}(X)$, the established estimate and Theorem 4.8 show that
$-.\displaystyle\mathsf{d}_{S}(\mu_{n},\mu)\to 0\quad\Leftrightarrow\quad\lVert B(\mu_{n})-B(\mu)\rVert_{\mathcal{H}_{c}}\to 0\quad\Leftrightarrow\quad\mu_{n}\to\mu\textnormal{ weak-$\ast$.}$- ∗ .
5.2 Existence of geodesics
Given the definition of $\mathsf{d}_{S}$ as infimization over the energy of paths, it is natural to ask whether geodesics in the sense of energy minimizing paths exist. The set of admissible paths $P$ has been chosen carefully to obtain existence of geodesics for $\mathsf{d}_{S}$.
Theorem 5.3.
The variational problem (32) defining $\mathsf{d}_{S}$ has a minimizer. Hence there exist Riemannian geodesics for the metric $\mathsf{d}_{S}$.
The proof will follow the direct method of calculus of variations. Indeed, we first show lower semicontinuity of $E$ on $\mathcal{H}^{1}((0,1);\mathcal{H}_{c})$ by rewriting $E((\mu_{t})_{t})$ as a supremum of weakly lower semicontinuous problems in Lemma 5.5. Using that boundedness in $E$ implies boundedness in the Sobolev norm, we then find that the weak cluster points of any minimizing sequence for the least energy problem (32) are minimizers, hence geodesic. This geodesic belongs to $P$, as $\mathcal{H}^{1}((0,1);\mathcal{B})$ is weakly closed.
Remark 5.4.
We can obtain the existence of geodesics for $\mathsf{d}_{S}$ in the metric sense in a much simpler way by showing that in $(\mathcal{P}(X),\mathsf{d}_{S})$ all pairs of points almost admit a midpoint (see [11, Def. 3.6]) and applying the Hopf–Rinow Theorem [11, Lem. B.3]. For any $\mu_{0},\mu_{1}\in\mathcal{P}(X)$ this would give a path $(\mu_{t})_{t}$ in $\mathcal{P}(X)$, such that for all $t,s\in[0,1]$ it holds
$\mathsf{d}_{S}(\mu_{t},\mu_{s})=\lvert t-s\rvert\cdot\mathsf{d}_{S}(\mu_{0},\mu_{1}).$
However, it is not immediately clear whether the path $(\mu_{t})_{t}$ belongs to $P(\mu_{0},\mu_{1})$ and whether $\mathsf{d}_{S}(\mu_{0},\mu_{1})=E((\mu_{t})_{t})$, i.e. whether $(\mu_{t})_{t}$ is also a geodesic in the Riemannian sense as a minimizer of the energy in (32).
Lemma 5.5.
Let $(\mu_{t})_{t}\in P$ and define $\bm{\mu}\in\mathcal{M}([0,1]\times X)$ via $\langle\bm{\mu},\psi\rangle=\iint\psi(t,\cdot)\,\mathrm{d}\mu_{t}\,\mathrm{d}t$ for any $\psi\in\mathcal{C}([0,1]\times X)$. The energy of the path $(\mu_{t})_{t}$ rewrites as
$E((\mu_{t})_{t})=\frac{\varepsilon}{2}\int_{0}^{1}\lVert\dot{\beta}_{t}\rVert_{\mathcal{H}_{c}}^{2}\,\mathrm{d}t+2\varepsilon\cdot\sup_{\psi}\Bigg{\{}\int_{0}^{1}\langle H_{c}[\psi(t,\cdot)A(\mu_{t})],\dot{\beta}_{t}\rangle_{\mathcal{H}_{c}}\,\mathrm{d}t\\ -\frac{1}{2}\int_{0}^{1}\langle\psi(t,\cdot),(\operatorname{id}-K_{\mu_{t}})\psi(t,\cdot)\rangle_{L^{2}(X,\mu_{t})}\,\mathrm{d}t\Bigg{\}},$ (33)
where the supremum in $\psi$ is taken over any space of functions dense in $L^{2}([0,1]\times X,\bm{\mu})$, e.g. the jointly continuous functions over $[0,1]\times X$.
Proof.
It is a well-known fact in convex analysis that if $T$ is a symmetric, coercive and continuous operator on a Hilbert space $\mathcal{H}$ it holds
$\frac{1}{2}\langle x,T^{-1}x\rangle_{\mathcal{H}}=\sup_{y}\,\langle x,y\rangle_{\mathcal{H}}-\frac{1}{2}\langle y,Ty\rangle_{\mathcal{H}}$ (34)
for all $x\in\mathcal{H}$. This is Fenchel-Legendre duality, or could also be seen as a corollary of the Lax-Milgram theorem [12, Corollary 5.8]. Continuity of $T$ allows the restriction of the supremum to a dense subset of $\mathcal{H}$. We apply (34) to
[…]
  • the operator $T[\psi](t,y)=(\operatorname{id}-K_{\mu_{t}})[\psi(t,\cdot)](y)$,
  • the point $x=(a_{\mu_{t}}\dot{\beta}_{t})_{t}\in\mathcal{H}$.
The operator $T$ is coercive and continuous on $\mathcal{H}$ by Theorem 4.3. In this notation
$E((\mu_{t})_{t})=\int_{0}^{1}\tilde{\mathbf{g}}_{\mu_{t}}(\dot{\beta}_{t},\dot{\beta}_{t})\,\mathrm{d}t=\frac{\varepsilon}{2}\int_{0}^{1}\lVert\dot{\beta}_{t}\rVert^{2}_{\mathcal{H}_{c}}\,\mathrm{d}t+\varepsilon\langle x,T^{-1}x\rangle_{\mathcal{H}}.$
We then use that $\langle a_{\mu_{t}}\dot{\beta}_{t},\psi(t,\cdot)\rangle_{L^{2}(X,\mu_{t})}=\langle\psi(t,\cdot)a_{\mu_{t}}\mu_{t},\dot{\beta}_{t}\rangle=\langle H_{c}[\psi(t,\cdot)A(\mu_{t})],\dot{\beta}_{t}\rangle_{\mathcal{H}_{c}}$. With $\langle\dot{\beta}_{t},\beta_{t}\rangle=0$ and $(\operatorname{id}-K_{\mu_{t}})[\mathbb{1}_{X}]=0$, one sees that the expression on the right-hand side of (33) gives the same value on functions whose difference is constant in space for almost all times $t$. Hence the supremum remains the same when taken on $\mathcal{H}$ or $L^{2}([0,1]\times X;\bm{\mu})$. ∎
Proof of Theorem 5.3.
We apply the direct method of calculus of variations. Let $(\mu^{n}_{t})_{t}\in P(\mu_{0},\mu_{1})$ with corresponding $(\beta^{n}_{t})_{t}\in\mathcal{H}^{1}((0,1);\mathcal{H}_{c})$ be such that
$\mathsf{d}_{S}(\mu_{0},\mu_{1})^{2}=\lim_{n\to\infty}\int_{0}^{1}\tilde{\mathbf{g}}_{\mu^{n}_{t}}(\dot{\beta}^{n}_{t},\dot{\beta}^{n}_{t})\,\mathrm{d}t.$
From $\lVert\beta_{t}^{n}\rVert_{\mathcal{H}_{c}}=1$ and Proposition 4.14 we see that the sequence of paths $(\beta^{n}_{t})_{t}$ must be bounded in $\mathcal{H}^{1}((0,1);\mathcal{H}_{c})$. Hence, up to taking a subsequence, $(\beta^{n}_{t})_{t}$ converges weakly in $\mathcal{H}^{1}((0,1);\mathcal{H}_{c})$ to some limit path $(\beta_{t})_{t}$. It remains to check that $\beta_{t}\in\mathcal{B}$ for all $t\in(0,1)$ and that the corresponding path of $\mu_{t}=B^{-1}(\beta_{t})$ satisfies $E((\mu_{t})_{t})=\mathsf{d}_{S}(\mu_{0},\mu_{1})^{2}$.
Note that the sequence $(\beta^{n}_{t})$ is uniformly equi-continuous: For any $n\in\mathbb{N}$, $s<t\in[0,1]$ set $v=(\beta^{n}_{t}-\beta^{n}_{s})/\lVert\beta^{n}_{t}-\beta^{n}_{s}\rVert_{\mathcal{H}_{c}}$. Then Theorem C.2 and the Cauchy–Schwarz inequality give
$\lVert\beta^{n}_{t}-\beta^{n}_{s}\rVert_{\mathcal{H}_{c}}=\int_{s}^{t}\langle\dot{\beta}^{n}_{\tau},v\rangle\,\mathrm{d}\tau=\langle(\dot{\beta}^{n}_{\tau})_{\tau},v\cdot\mathbb{1}_{[s,t]}\rangle_{L^{2}([0,1];\mathcal{H}_{c})}\leq\lVert(\beta^{n}_{\tau})_{\tau}\rVert_{\mathcal{H}^{1}((0,1);\mathcal{H}_{c})}\cdot\sqrt{|t-s|}.$
Thus, again up to taking a subsequence, we can assume that $(\beta^{n}_{t})$ converges uniformly, hence the limit satisfies $\beta_{t}\in\mathcal{B}$ for all $t\in(0,1)$.
We now estimate the energy. By lower semicontinuity of the norm for weak convergence, we have directly
$\int_{0}^{1}\lVert\dot{\beta}_{t}\rVert_{\mathcal{H}_{c}}^{2}\,\mathrm{d}t\leq\liminf_{n\to+\infty}\int_{0}^{1}\lVert\dot{\beta}^{n}_{t}\rVert_{\mathcal{H}_{c}}^{2}\,\mathrm{d}t.$
Thus by Lemma 5.5, using $\alpha^{n}_{t}=A(\mu^{n}_{t})=a_{\mu^{n}_{t}}\mu^{n}_{t}$ and rewriting the operator $K_{\mu^{n}_{t}}$ using $k_{\mu^{n}_{t}}=k_{c}\cdot a_{\mu^{n}_{t}}\otimes a_{\mu^{n}_{t}}$, to conclude the proof it suffices to show that for all $\psi\in\mathcal{C}([0,1]\times X)$ we have
$\liminf_{n\to\infty}\int_{0}^{1}\left\langle H_{c}[\psi(t,\cdot)\alpha^{n}_{t}],\dot{\beta}^{n}_{t}\right\rangle_{\mathcal{H}_{c}}\mathrm{d}t-\frac{1}{2}\int_{0}^{1}\int_{X}\psi(t,x)^{2}\,\mathrm{d}\mu^{n}_{t}(x)\,\mathrm{d}t$
$\qquad\qquad+\frac{1}{2}\int_{0}^{1}\iint_{X\times X}k_{c}(x,y)\psi(t,x)\psi(t,y)\,\mathrm{d}\alpha^{n}_{t}(x)\,\mathrm{d}\alpha^{n}_{t}(y)\,\mathrm{d}t$
$\qquad\geq\int_{0}^{1}\left\langle H_{c}[\psi(t,\cdot)\alpha_{t}],\dot{\beta}_{t}\right\rangle_{\mathcal{H}_{c}}\mathrm{d}t-\frac{1}{2}\int_{0}^{1}\int_{X}\psi(t,x)^{2}\,\mathrm{d}\mu_{t}(x)\,\mathrm{d}t$
$\qquad\qquad+\frac{1}{2}\int_{0}^{1}\iint_{X\times X}k_{c}(x,y)\psi(t,x)\psi(t,y)\,\mathrm{d}\alpha_{t}(x)\,\mathrm{d}\alpha_{t}(y)\,\mathrm{d}t.$
As $\beta^{n}_{t}$ converges to $\beta_{t}$, we know from Theorem 4.8 that $\mu^{n}_{t}\to\mu_{t}$, $\alpha^{n}_{t}\to\alpha_{t}$ weak-$\ast$ in $\mathcal{M}(X)$. In addition, we can bound the mass of $\alpha^{n}_{t}$ and the supremum of $\psi(t,\cdot)$ uniformly in $t$ and $n$, and so the second and third integral are easy to pass to the limit. For the first one note that
$(H_{c}[\psi(t,\cdot)\alpha^{n}_{t}])_{t}\to(H_{c}[\psi(t,\cdot)\alpha_{t}])_{t}.$
The convergence is pointwise in $t$ thanks to Lemma B.2, and also in norm in $L^{2}([0,1];\mathcal{H}_{c})$ thanks to the Lebesgue dominated convergence theorem. Since the first integral rewrites as
$\int_{0}^{1}\left\langle H_{c}[\psi(t,\cdot)\alpha^{n}_{t}],\dot{\beta}^{n}_{t}\right\rangle_{\mathcal{H}_{c}}\,\mathrm{d}t=\left\langle(H_{c}[\psi(t,\cdot)\alpha^{n}_{t}])_{t},(\dot{\beta}^{n}_{t})_{t}\right\rangle_{L^{2}([0,1];\mathcal{H}_{c})},$
convergence to the desired limit holds as a pairing between weakly and strongly convergent sequences in $L^{2}([0,1];\mathcal{H}_{c})$ (see Lemma A.2). ∎
5.3 Mean estimate for the squared Euclidean distance
In this section $c(x,y)=\lVert x-y\rVert^{2}$ is the squared Euclidean distance and we work with measures in $\mathbb{R}^{d}$ with bounded support. For such a measure $\mu$, we write $m=\langle\mu,\operatorname{id}\rangle\in\mathbb{R}^{d}$ for its mean, and $\bar{\mu}=(\operatorname{id}-m)_{\#}\mu$ for the centered measure. As recalled in the introduction, it is well understood that for classical optimal transport the cost $\mathrm{OT}_{0}$ between $\mu_{0}$ and $\mu_{1}$ decomposes as
$\mathrm{OT}_{0}(\mu_{0},\mu_{1})=\lVert m_{1}-m_{0}\rVert^{2}+\mathrm{OT}_{0}(\bar{\mu}_{0},\bar{\mu}_{1}).$
The same proof carries through for the Sinkhorn divergence (Proposition 5.6). We then extend this orthogonal decomposition result to the metric tensor in Proposition 5.7 and to the distance $\mathsf{d}_{S}$ in Theorem 5.9, which in particular implies that translations $\mu_{t}=(\operatorname{id}+t\cdot v)_{\#}\mu_{0}$ for $v\in\mathbb{R}^{d}$ are geodesics for $\mathsf{d}_{S}$.
Proposition 5.6.
On $\mathbb{R}^{d}$ let $c(x,y)=\lVert x-y\rVert^{2}$. Let $\mu_{0},\mu_{1}\in\mathcal{P}(\mathbb{R}^{d})$ with bounded support and means $m_{0},m_{1}\in\mathbb{R}^{d}$. Their centered versions $\bar{\mu}_{0},\bar{\mu}_{1}$ satisfy
$S_{\varepsilon}(\mu_{0},\mu_{1})=\lVert m_{1}-m_{0}\rVert^{2}+S_{\varepsilon}(\bar{\mu}_{0},\bar{\mu}_{1}).$
Proof.
The proof is basically the same as for plain optimal transport, where $\varepsilon=0$. For the squared Euclidean distance, optimal couplings for $\mathrm{OT}_{\varepsilon}$ remain optimal under translations in the following sense: for all $\pi\in\Pi(\mu_{0},\mu_{1})$ it holds $(\operatorname{id}-m_{0},\operatorname{id}-m_{1})_{\#}\pi\in\Pi(\bar{\mu}_{0},\bar{\mu}_{1})$ and
$\langle(\operatorname{id}-m_{0},\operatorname{id}-m_{1})_{\#}\pi,c\rangle$ $=\int_{X\times X}\lVert x-m_{0}-(y-m_{1})\rVert^{2}\,\mathrm{d}\pi(x,y)$
$=\int_{X\times X}\lVert x-y\rVert^{2}-2\langle x-y,m_{0}-m_{1}\rangle+\lVert m_{1}-m_{0}\rVert^{2}\,\mathrm{d}\pi(x,y)$
$=\langle\pi,c\rangle-\lVert m_{1}-m_{0}\rVert^{2}.$
The $\operatorname{KL}$-divergence is invariant under such shifts, i.e.
$\operatorname{KL}((\operatorname{id}-m_{0},\operatorname{id}-m_{1})_{\#}\pi\,|\,(\operatorname{id}-m_{0})_{\#}\mu_{0}\otimes(\operatorname{id}-m_{1})_{\#}\mu_{1})=\operatorname{KL}(\pi\,|\,\mu_{0}\otimes\mu_{1}).$
Therefore, minimizing over $\pi$ yields $\mathrm{OT}_{\varepsilon}(\mu_{0},\mu_{1})=\lVert m_{1}-m_{0}\rVert^{2}+\mathrm{OT}_{\varepsilon}(\bar{\mu}_{0},\bar{\mu}_{1})$. Analogously one sees that $\mathrm{OT}_{\varepsilon}(\mu_{i},\mu_{i})=\mathrm{OT}_{\varepsilon}(\bar{\mu}_{i},\bar{\mu}_{i})$ for $i=0,1$. From the definition of $S_{\varepsilon}$ in (6) we obtain the result. ∎
We now extend this decomposition to the level of the metric tensor. If $v\in\mathbb{R}^{d}$, the uniform translation of a measure $\mu$ by $v$ corresponds to the derivative $\dot{\mu}=-\operatorname{div}(v\mu)\in\mathcal{H}_{\mu,0}^{*}$: this is the continuity equation. For $\mu\in\mathcal{P}(X)$ let $\mathcal{S}_{\mu}:=\left\{-\operatorname{div}(v\mu)\,\middle|\,v\in\mathbb{R}^{d}\right\}\subseteq\mathcal{H}_{\mu,0}^{*}$ be the set of constant shift perturbations of $\mu$. Denote the orthogonal projection onto $\mathcal{S}_{\mu}$ with respect to the inner product $\mathbf{g}_{\mu}$ by
$P_{\mathcal{S}_{\mu}}\colon\mathcal{H}_{\mu,0}^{*}\to\mathcal{S}_{\mu}.$
Proposition 5.7.
Let $X\subseteq\mathbb{R}^{d}$ be compact, $c(x,y)=\lVert x-y\rVert^{2}$. Let $\mu\in\mathcal{P}(X)$ and let $\dot{\mu}\in\mathcal{H}_{\mu,0}^{*}$. Then
  1. 1.
    the mean $\dot{m}=\langle\dot{\mu},\operatorname{id}\rangle$ exists in $\mathbb{R}^{d}$ and is characterized by $P_{\mathcal{S}_{\mu}}(\dot{\mu})=-\operatorname{div}(\dot{m}\mu)$,
  2. 2.
    there holds $\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})=\lVert\dot{m}\rVert^{2}+\mathbf{g}_{\mu}(\dot{\mu}+\operatorname{div}(\dot{m}\mu),\dot{\mu}+\operatorname{div}(\dot{m}\mu))$.
Proof.
For $v\in\mathbb{R}^{d}$ denote $\dot{\mu}^{v}=-\operatorname{div}(v\mu)$. As a distribution of order one it belongs to $\mathcal{H}_{\mu}^{*}$ (see Remark 4.6). We compute
$\begin{split}H_{\mu}[\dot{\mu}^{v}](x)&=\int_{X}\langle v,\nabla_{y}k_{\mu}(x,y)\rangle\,\mathrm{d}\mu(y)\\ &=\frac{1}{\varepsilon}\int_{X}\langle v,\nabla{f_{\mu,\mu}}(y)-2y+2x\rangle k_{\mu}(x,y)\,\mathrm{d}\mu(y)\\ &=\frac{1}{\varepsilon}K_{\mu}[\langle v,\nabla{f_{\mu,\mu}}-2\operatorname{id}\rangle](x)+\frac{2}{\varepsilon}\langle v,x\rangle.\end{split}$ (35)
Differentiating the relation ${f_{\mu,\mu}}(y)=T_{\varepsilon}({f_{\mu,\mu}},\mu)(y)$ in $y$ yields
$2y-\nabla{f_{\mu,\mu}}(y)=\int_{X}k_{\mu}(y,z)2z\,\mathrm{d}\mu(z)=2K_{\mu}[\operatorname{id}](y).$ (36)
Thus $\langle v,\nabla{f_{\mu,\mu}}-2\operatorname{id}\rangle=-2K_{\mu}[\langle v,\operatorname{id}\rangle]$. Applying this in (35), we deduce the identity $\frac{\varepsilon}{2}(\operatorname{id}-K_{\mu}^{2})^{-1}H_{\mu}[\dot{\mu}^{v}]=\langle v,\operatorname{id}\rangle$ in $\mathcal{H}_{\mu}/\mathbb{R}$. In particular, the function $\langle v,\operatorname{id}\rangle$ belongs to $\mathcal{H}_{\mu}/\mathbb{R}$ for any $v\in\mathbb{R}^{d}$. Since every coordinate function $x\mapsto x_{i}$ belongs to $\mathcal{H}_{\mu}/\mathbb{R}$, the mean $\dot{m}=\langle\dot{\mu},\operatorname{id}\rangle$ is well defined as an element of $\mathbb{R}^{d}$.
Given the expression of the metric tensor (Definition 4.1), we see that
$\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu}^{v})=\langle\dot{\mu},\langle v,\operatorname{id}\rangle\rangle=\langle v,\dot{m}\rangle=\mathbf{g}_{\mu}(\dot{\mu}^{\dot{m}},\dot{\mu}^{v})$
Therefore $\dot{\mu}^{\dot{m}}$ is the orthogonal projection of $\dot{\mu}$ to $\mathcal{S}_{\mu}$ with respect to the metric tensor $\mathbf{g}_{\mu}$. The second point of the proposition is a direct consequence of the Pythagorean identity for the inner product $\mathbf{g}_{\mu}$. ∎
Remark 5.8.
For vertical (18) or horizontal (20) perturbations, for which the metric tensor is linked to the Sinkhorn divergence via Theorem 3.3 or Remark 3.14, the proof of Proposition 5.7 can be done directly as an application of Proposition 5.6.
The orthogonal decomposition at the level of the metric tensor from Proposition 5.7 can be integrated and extended to the level of the distance $\mathsf{d}_{S}$.
Theorem 5.9.
On $\mathbb{R}^{d}$ let $c(x,y)=\lVert x-y\rVert^{2}$. Let $\mu_{0},\mu_{1}\in\mathcal{P}(\mathbb{R}^{d})$ with bounded support and means $m_{0},m_{1}\in\mathbb{R}^{d}$ and centered measures $\bar{\mu}_{0}$, $\bar{\mu}_{1}$. Then
$\mathsf{d}_{S}(\mu_{0},\mu_{1})^{2}=\lVert m_{1}-m_{0}\rVert^{2}+\mathsf{d}_{S}(\bar{\mu}_{0},\bar{\mu}_{1})^{2}.$
In particular $\mathsf{d}_{S}(\mu_{0},\mu_{1})\geq\lVert m_{1}-m_{0}\rVert$ with equality if and only if $\mu_{1}$ is a translation of $\mu_{0}$. In this case $\mu_{t}=(\operatorname{id}+t(m_{1}-m_{0}))_{\#}\mu_{0}$ is the unique geodesic between them.
Before diving into the proof of the theorem, we need an additional technical lemma. Whereas Proposition 5.7 is concerned with what is happening at the level of the metric tensor, we need to check additionally that the map $\mu\mapsto\bar{\mu}=(\operatorname{id}-m)_{\#}\mu$ with $m=\langle\mu,\operatorname{id}\rangle$ preserves the class of admissible curves.
Lemma 5.10.
Let $(\mu_{t})_{t}\in P$ with support in some bounded set and write $m_{t}$ and $\bar{\mu}_{t}$ for the mean and centered part of $\mu_{t}$. Then $(\bar{\mu}_{t})\in P$ (with support in some possibly larger bounded set) and $(m_{t})\in\mathcal{H}^{1}((0,1);\mathbb{R}^{d})$.
Proof.
We start with a preliminary remark. We proved in Proposition 5.7 that for any $v\in\mathbb{R}^{d}$ and any $\mu$ we have $\langle v,\operatorname{id}\rangle=\frac{\varepsilon}{2}(\operatorname{id}-K_{\mu}^{2})^{-1}H_{\mu}[-\operatorname{div}(\mu v)]$. Thus with the help of Theorem 4.3 and Remark 4.6 we have $\langle v,\operatorname{id}\rangle\in\mathcal{H}_{\mu}$ and
$\|\langle v,\operatorname{id}\rangle\|_{\mathcal{H}_{\mu}}\leq C\lVert v\rVert$ (37)
for a constant $C$ which only depends on the radius of the support of $\mu\in\mathcal{P}(X)$.
We first prove that $m_{t}\in\mathcal{H}^{1}([0,1],\mathbb{R}^{d})$. It is enough to prove that $\langle v,m_{t}\rangle\in\mathcal{H}^{1}([0,1],\mathbb{R})$ for any $v\in\mathbb{R}^{d}$. Recall from (31) that for any $\phi\in\mathcal{H}_{\mu_{s}}$ and any $t,s$,
$\langle\mu_{t}-\mu_{s},\phi\rangle=\left\langle\beta_{t}-\beta_{s},a_{\mu_{s}}^{-1}\phi\right\rangle_{\mathcal{H}_{c}}+\left\langle\beta_{t}-\beta_{s},H_{c}[\alpha_{t}\phi]\right\rangle_{\mathcal{H}_{c}}.$
For $\phi=\langle v,\operatorname{id}\rangle$, with the help of the Cauchy–Schwarz inequality and using that multiplication by $a_{\mu_{s}}^{-1}$ is an isometry from $\mathcal{H}_{\mu_{s}}$ to $\mathcal{H}_{c}$, this gives
$|\langle v,m_{t}-m_{s}\rangle|\leq\lVert\beta_{t}-\beta_{s}\rVert_{\mathcal{H}_{c}}\left(\lVert\langle v,\operatorname{id}\rangle\rVert_{\mathcal{H}_{\mu_{s}}}+\lVert H_{c}[\alpha_{t}\langle v,\operatorname{id}\rangle]\rVert_{\mathcal{H}_{c}}\right).$
Given that $(\beta_{t})_{t}\in\mathcal{H}([0,1];\mathcal{H}_{c})$, we need only to prove that $\lVert\phi\rVert_{\mathcal{H}_{\mu_{s}}}+\lVert H_{c}[\alpha_{t}\phi]\rVert_{\mathcal{H}_{c}}$ can be bounded independently on $t,s$ to conclude with Theorem C.2. This is a consequence of (37) for the first term while for the second term it is enough to notice that $\alpha_{t}$ is bounded in total variation uniformly in $t$.
We now proceed to prove that $\bar{\mu}_{t}\in P$. By uniqueness of the Schrödinger potentials we see that $f_{\bar{\mu}_{t},\bar{\mu}_{t}}(x)=f_{\mu_{t},\mu_{t}}(x+m_{t})$. Writing $\bar{\beta}_{t}=B(\bar{\mu}_{t})$, this gives $\bar{\beta}_{t}(x)=\beta_{t}(x+m_{t})$. Thus for any $t,s$,
$\lVert\bar{\beta}_{t}-\bar{\beta}_{s}\rVert_{\mathcal{H}_{c}}$ $=\lVert\beta_{t}(\cdot+m_{t})-\beta_{s}(\cdot+m_{s})\rVert_{\mathcal{H}_{c}}$
$\leq\lVert\beta_{t}(\cdot+m_{t})-\beta_{s}(\cdot+m_{t})\rVert_{\mathcal{H}_{c}}+\lVert\beta_{s}(\cdot+m_{t})-\beta_{s}(\cdot+m_{s})\rVert_{\mathcal{H}_{c}}.$
Using that the norm on $\mathcal{H}_{c}$ is invariant by translation because so is the kernel, the first term coincides with $\lVert\beta_{t}-\beta_{s}\rVert_{\mathcal{H}_{c}}$. For the second term, we can write
$\beta_{s}(x-m_{t})-\beta_{s}(x-m_{s})=\int_{s}^{t}\dot{m}_{r}\cdot\nabla\beta_{s}(x-m_{r})\,\mathrm{d}r,$
and so we only need to check that $\nabla\beta_{s}(\cdot-m_{r})\in\mathcal{H}_{c}$ with a controlled norm to justify that the equality above holds as equality of functions in $\mathcal{H}_{c}$. Recalling $\beta=\exp(-{f_{\mu,\mu}}/\varepsilon)$ and recalling that we have already computed $\nabla{f_{\mu,\mu}}$ in (36), we find
$\dot{m}_{r}\cdot\nabla\beta_{s}=\frac{2}{\varepsilon}\left(K_{\mu_{s}}[\langle\dot{m}_{r},\operatorname{id}\rangle]-\langle\dot{m}_{r},\operatorname{id}\rangle\right)\beta_{s}.$
Using invariance by translation of the norm in $\mathcal{H}_{c}$, and that multiplication by $\beta_{s}=a_{\mu_{s}}^{-1}$ is an isometry between $\mathcal{H}_{\mu_{s}}$ and $\mathcal{H}_{c}$,
$\lVert\dot{m}_{r}\cdot\nabla\beta_{s}(\cdot-m_{s})\rVert_{\mathcal{H}_{c}}=\lVert\dot{m}_{r}\cdot\nabla\beta_{s}\rVert_{\mathcal{H}_{c}}\leq\frac{2}{\varepsilon}\left(\lVert K_{\mu_{s}}[\langle\dot{m}_{r},\operatorname{id}\rangle]\rVert_{\mathcal{H}_{\mu_{s}}}+\lVert\langle\dot{m}_{r},\operatorname{id}\rangle\rVert_{\mathcal{H}_{\mu_{s}}}\right).$
We use that $K_{\mu_{s}}$ is a contraction (see Proposition 4.2), and again (37) to conclude that the right-hand side can be bounded by $C\cdot\lVert\dot{m}_{r}\rVert$ for some $C>0$ independently on $s$. Thus
$\lVert\bar{\beta}_{t}-\bar{\beta}_{s}\rVert_{\mathcal{H}_{c}}\leq\lVert\beta_{t}-\beta_{s}\rVert_{\mathcal{H}_{c}}+C\int_{t}^{s}\lVert\dot{m}_{r}\rVert\,\mathrm{d}r.$
From Theorem C.2 we conclude that $(\bar{\beta}_{t})_{t}\in\mathcal{H}^{1}([0,1];\mathcal{H}_{c})$ so that $(\bar{\mu}_{t})_{t}\in P$. ∎
Proof of Theorem 5.9.
If $(\mu_{t})_{t}\in P(\mu_{0},\mu_{1})$, then by Lemma 5.10 $(B(\bar{\mu}_{t}))_{t}\in P$ and $(m_{t})_{t}$ exists in $\mathcal{H}^{1}((0,1);\mathbb{R}^{d})$. Writing $\dot{\bar{\mu}}_{t}$ for the derivative of the curve $t\mapsto\bar{\mu}_{t}$ and $\bar{\dot{\mu}}_{t}$ for the translation of the perturbation $\dot{\mu}_{t}$ by $-m_{t}$, it easy to check that $\dot{\bar{\mu}}_{t}=\bar{\dot{\mu}}_{t}+\operatorname{div}(\dot{m}_{t}\bar{\mu}_{t})$ holds in $\mathcal{H}^{*}_{\mu_{t},0}$ for any $t$ when all three derivatives exists. In particular, translation invariance of the metric tensor yields
$\mathbf{g}_{\mu_{t}}(\dot{\mu}+\operatorname{div}(\dot{m}\mu),\dot{\mu}_{t}+\operatorname{div}(\dot{m}_{t}\mu_{t}))=\mathbf{g}_{\bar{\mu}_{t}}(\bar{\dot{\mu}}_{t}+\operatorname{div}(\dot{m}_{t}\bar{\mu}_{t}),\bar{\dot{\mu}}_{t}+\operatorname{div}(\dot{m}_{t}\bar{\mu}_{t}))=\mathbf{g}_{\bar{\mu}_{t}}(\dot{\bar{\mu}}_{t},\dot{\bar{\mu}}_{t}).$
From Proposition 5.7 we deduce that
$\int_{0}^{1}\mathbf{g}_{\mu_{t}}(\dot{\mu}_{t},\dot{\mu}_{t})\,\mathrm{d}t$ $=\int_{0}^{1}\lVert\dot{m}_{t}\rVert^{2}\,\mathrm{d}t+\int_{0}^{1}\mathbf{g}_{\mu_{t}}(\dot{\mu}+\operatorname{div}(\dot{m}\mu),\dot{\mu}_{t}+\operatorname{div}(\dot{m}_{t}\mu_{t}))\,\mathrm{d}t$
$=\int_{0}^{1}\lVert\dot{m}_{t}\rVert^{2}\,\mathrm{d}t+\int_{0}^{1}\mathbf{g}_{\bar{\mu}_{t}}(\dot{\bar{\mu}}_{t},\dot{\bar{\mu}}_{t})\,\mathrm{d}t.$
Conversely, from a curve of centered measures $(\bar{\mu}_{t})_{t}$ and a path $(m_{t})_{t}\in\mathcal{H}^{1}([0,1],\mathbb{R}^{d})$ of means we can reconstruct $\mu_{t}=(\operatorname{id}+m_{t})_{\#}\bar{\mu}_{t}$. We get the decomposition of the distance by minimizing over $(\mu_{t})_{t}$.
The inequality follows, and if we have equality, then necessarily $(\bar{\mu}_{t})_{t}$ is a constant curve and $(m_{t})_{t}$ is a geodesic in $\mathbb{R}^{d}$. ∎
Corollary 5.11 (Energy of a travelling Dirac measure).
Let $c(x,y)=\lVert x-y\rVert^{2}$. Let $(x_{t})_{t}\in\mathcal{H}^{1}((0,1);\mathbb{R}^{d})$ and set $\mu_{t}=\delta_{x_{t}}$. Then
$E((\mu_{t})_{t})=\int_{0}^{1}\lVert\dot{x}_{t}\rVert^{2}\,\mathrm{d}t.$
In particular, $\mathsf{d}_{S}(\delta_{x_{0}},\delta_{x_{1}})=\lVert x_{0}-x_{1}\rVert$ agrees with the metric on $\mathbb{R}^{d}$ and the unique geodesic is $(\delta_{y_{t}})_{t\in[0,1]}$ with $y_{t}=(1-t)x_{0}+tx_{1}$.
6 Examples in the metric $\mathsf{d}_{S}$
6.1 Explicit formula for Gaussian measures
Figure 1: Representation of the Wasserstein and the (Gaussian-restricted) Sinkhorn-based geodesics for one-dimensional centered Gaussian distributions parametrized by their variance $v_{t}$, $t\in[0,1]$. In the right panel a translation was added for better visibility, cf. Theorem 5.9. For comparison we further show the $\varepsilon$-Schrödinger bridge.
Throughout this section we fix $c(x,y)=\lvert x-y\rvert^{2}$ as the squared Euclidean distance on $\mathbb{R}$. We derive the metric tensor as well as an explicit formula for the geodesic restriction of the distance $\mathsf{d}_{S}$ to the family of Gaussian measures in dimension one. By virtue of Theorem 5.9 we can restrict ourselves to centered measures. Strictly speaking, our theory a priori only applies to compactly supported measures, but we can still evaluate the metric tensor on Gaussians. We parametrize Gaussian measures by their variance $v$ rather than their standard deviation $\sigma=\sqrt{v}$. Writing $\mathcal{N}(0,v)$ for the centered Gaussian measure with variance $v$, we recall from [27, 61] that
$\mathrm{OT}_{0}(\mathcal{N}(0,v_{0}),\mathcal{N}(0,v_{1}))=\left(\sqrt{v_{1}}-\sqrt{v_{0}}\right)^{2},$
which corresponds to the unregularized metric tensor $\mathbf{g}^{0}_{\mathcal{N}(0,v)}(\dot{v},\dot{v})=\frac{1}{4v}\dot{v}^{2}$. As discussed in the introduction, page 1, geodesics with respect to $\mathsf{d}_{S}$ differ from a $\varepsilon$-Schrödinger bridge, for which an explicit formula for Gaussian measures can be found in [13]: the result is displayed in Figure 1. With our metric we obtain the following:
Theorem 6.1.
Consider a path $(v_{t})_{t}$ in $(0,\infty)$ that is differentiable around $t=0$ and denote by $\dot{v}_{t}$ its derivative at $t$. Identifying the Gaussian distribution $\mathcal{N}(0,v_{t})$ with its variance $v_{t}$ and writing $v=v_{0}$, $\dot{v}=\dot{v}_{0}$, the metric tensor reads
$\mathbf{g}_{v}(\dot{v},\dot{v})=g(v)\dot{v}^{2}\quad\text{with}\quad g(v)=\frac{1}{4\sqrt{\frac{\varepsilon^{2}}{16}+v^{2}}}.$ (38)
Proof.
We identify the Gaussian $\mu_{t}=\mathcal{N}(0,v_{t})$ with its variance $v_{t}$. As shown in [44, Cor. 1], the entropic optimal transport cost between 1-dimensional Gaussian distributions with variance $v_{t},v_{s}$ is explicitly given by
$\mathrm{OT}_{\varepsilon}(v_{t},v_{s})=v_{t}+v_{s}-\frac{\varepsilon}{2}\left(\kappa_{ts}-\log\kappa_{ts}+\log 2-2\right),\qquad\kappa_{ts}:=1+\sqrt{1+\frac{16}{\varepsilon^{2}}v_{t}v_{s}}.$ (39)
Let $v_{0}=v$, $\dot{v}_{0}=\dot{v}$. Let us introduce $f(x)=1+\sqrt{1+16\varepsilon^{-2}x}$, so that $\kappa_{ts}=f(v_{t}v_{s})$. We have
$\partial_{t}\kappa_{ts}=\dot{v}_{t}v_{s}f^{\prime}(v_{t}v_{s}),\qquad\partial_{s}\partial_{t}\kappa_{ts}=\dot{v}_{s}\dot{v}_{t}f^{\prime}(v_{t}v_{s})+\dot{v}_{s}\dot{v}_{t}v_{s}v_{t}f^{\prime\prime}(v_{t}v_{s}).$
As in the proof of Theorem 3.3, we now compute $\mathbf{g}_{v}(\dot{v},\dot{v})=-\frac{1}{2}\partial_{t}|_{t=0}\partial_{s}|_{s=0}\mathrm{OT}_{\varepsilon}(v_{t},v_{s})$. To this end, observe that
$\partial_{t}\partial_{s}\mathrm{OT}_{\varepsilon}(v_{t},v_{s})=\frac{\varepsilon}{2}\left(-\partial_{t}\partial_{s}\kappa_{ts}+\frac{\partial_{t}\partial_{s}\kappa_{ts}}{\kappa_{ts}}-\frac{\partial_{t}\kappa_{ts}\partial_{s}\kappa_{ts}}{\kappa_{ts}^{2}}\right).$
In the new notation this gives
$\partial_{t}|_{t=0}\partial_{s}|_{s=0}\mathrm{OT}_{\varepsilon}(v_{t},v_{s})=\frac{\varepsilon}{2}\dot{v}^{2}\left((f^{\prime}(v^{2})+v^{2}f^{\prime\prime}(v^{2}))\left(\frac{1}{f(v^{2})}-1\right)-v^{2}\frac{f^{\prime}(v^{2})^{2}}{f(v^{2})^{2}}\right).$
The metric tensor can thus be rewritten as
$\mathbf{g}_{v}(\dot{v},\dot{v})=g(v)\dot{v}^{2},\qquad g(v)=-\frac{\varepsilon}{4}\left\{v^{2}\left(f^{\prime\prime}(v^{2})\frac{1-f(v^{2})}{f(v^{2})}-\frac{f^{\prime}(v^{2})^{2}}{f(v^{2})^{2}}\right)+f^{\prime}(v^{2})\frac{1-f(v^{2})}{f(v^{2})}\right\}.$
It remains to simplify the expression for $g$. To this end, let us compute the first and second derivatives of $f$:
$f^{\prime}(x)=\frac{8}{\varepsilon^{2}}\left(1+\frac{16}{\varepsilon^{2}}x\right)^{-1/2},\qquad f^{\prime\prime}(x)=-\frac{64}{\varepsilon^{4}}\left(1+\frac{16}{\varepsilon^{2}}x\right)^{-3/2}.$
To simplify notations introduce the variable $u=1+16\varepsilon^{-2}v^{2}$ in such a way that
$f(v^{2})=1+\sqrt{u},\quad f^{\prime}(v^{2})=\frac{8}{\varepsilon^{2}\sqrt{u}},\quad f^{\prime\prime}(v^{2})=-\frac{64}{\varepsilon^{4}u^{3/2}},$
and in particular
$\frac{1-f(v^{2})}{f(v^{2})}=-\frac{\sqrt{u}}{1+\sqrt{u}}.$
Plugging this into the expression for $g$ yields
$\begin{split}g(v)&=\frac{\varepsilon}{4}\left\{-v^{2}\left(\frac{64}{\varepsilon^{4}}\frac{1}{u(1+\sqrt{u})}-\frac{64}{\varepsilon^{4}}\frac{1}{u(1+\sqrt{u})^{2}}\right)+\frac{8}{\varepsilon^{2}}\frac{1}{1+\sqrt{u}}\right\}\\ &=-\frac{16v^{2}}{\varepsilon^{3}}\frac{1}{\sqrt{u}(1+\sqrt{u})^{2}}+\frac{2}{\varepsilon}\frac{1}{1+\sqrt{u}}.\end{split}$
To lighten the notation further, use $\frac{16}{\varepsilon^{2}}v^{2}=u-1$ to get
$g(v)=-\frac{1}{\varepsilon}\frac{u-1}{\sqrt{u}(1+\sqrt{u})^{2}}+\frac{2}{\varepsilon}\frac{1}{1+\sqrt{u}}=\frac{1}{\varepsilon}\frac{1+2\sqrt{u}+u}{\sqrt{u}(1+\sqrt{u})^{2}}=\frac{1}{\varepsilon\sqrt{u}}.$
The claim follows when unpacking $u=1+\frac{16}{\varepsilon^{2}}v^{2}$. ∎
We now turn to the metric on $(0,\infty)$ induced by the tensor $\mathbf{g}$ of Theorem 6.1. Said differently, we look at the geodesic restriction of $\mathsf{d}_{S}$ to Gaussian measures. At this point we do not know whether the set of Gaussian measures is in fact geodesically convex.
Theorem 6.2.
Let $F$ be the antiderivative of the function $x\mapsto(1+x^{2})^{-1/4}$. Let $v_{0},v_{1}\geqslant 0$. Then the geodesic restriction $\widehat{\mathsf{d}_{S}}$ of $\mathsf{d}_{S}$ to Gaussians can be expressed as
$\widehat{\mathsf{d}_{S}}(v_{0},v_{1})=\frac{\sqrt{\varepsilon}}{4}\left|F\left(\frac{4}{\varepsilon}v_{1}\right)-F\left(\frac{4}{\varepsilon}v_{0}\right)\right|.$
The geodesic $(v_{t})_{t\in[0,1]}$ joining $v_{0}$ to $v_{1}$ is given by
$v_{t}=\frac{\varepsilon}{4}F^{-1}\left[(1-t)F\left(\frac{4}{\varepsilon}v_{0}\right)+tF\left(\frac{4}{\varepsilon}v_{1}\right)\right]$
and the map $t\mapsto v_{t}$ is convex.
Proof.
Given the metric tensor (38) on $(0,\infty)$, geodesics correspond to minimizers of the energy
$E((v_{t})_{t})=\int_{0}^{1}g(v_{t})\dot{v}_{t}^{2}\,\mathrm{d}t.$
It is easy to see that geodesics must be monotone and since $g$ is smooth on the interval $[v_{0},v_{1}]$ we can use standard tools from smooth calculus of variations on one-dimensional curves to obtain the minimizer. Let us call $q_{t}=\dot{v}_{t}$, and $L(v,q)=g(v)q^{2}$ the Lagrangian. Recall that the Euler–Lagrange equation is given by
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{\partial L}{\partial q}(v_{t},\dot{v_{t}})\right)=\frac{\partial L}{\partial v}(v_{t},\dot{v}_{t}).$
Expanding it yields
$g(v_{t})\ddot{v}_{t}=-\frac{g^{\prime}(v_{t})}{2}\dot{v}_{t}^{2}.$
A simple inspection reveals that $g$ is non-negative and decreasing. Thus $g^{\prime}\leqslant 0$ and the equation tells us that $\ddot{v}\geqslant 0$, which is equivalent to the fact that $(v_{t})_{t}$ is convex in time. Note that if $v_{0}\neq v_{1}$, then $\dot{v}$ cannot vanish: if it were the case, by the unique solvability of the above ODE, $v$ should be a constant function. Thus $(v_{t})_{t}$ is monotone, and without loss of generality we focus on $v_{0}\leqslant v_{1}$. In this case $(v_{t})_{t}$ is increasing in time. For solutions of the above ODE one has that $C=g(v_{t})\dot{v}_{t}^{2}$ is constant in time (as usual for geodesics) and non-negative and thus $v$ satisfies the first-order equation
$\dot{v}_{t}=\sqrt{\frac{C}{g(v_{t})}}=2\sqrt{C}\left(\frac{\varepsilon^{2}}{16}+v_{t}^{2}\right)^{1/4}=\sqrt{C\varepsilon}\left(1+\left(\frac{4v_{t}}{\varepsilon}\right)^{2}\right)^{1/4},$
This equation is separable and thus can be integrated. Recalling the definition of $F$ above, the ODE yields
$\frac{\mathrm{d}}{\mathrm{d}t}F\left(\frac{4}{\varepsilon}v_{t}\right)=4\sqrt{\frac{C}{\varepsilon}},$ (40)
and thus integrating between 00 and $1$ gives
$C=\frac{\varepsilon}{16}\left[F\left(\frac{4}{\varepsilon}v_{1}\right)-F\left(\frac{4}{\varepsilon}v_{0}\right)\right]^{2}.$
Now $\widehat{\mathsf{d}_{S}}(v_{0},v_{1})=\sqrt{C}$ and the formula for $v_{t}$ follows from (40). ∎
6.2 The two-point space
In this section, we analyze the Riemannian tensor $\mathbf{g}_{\mu}$ and the metric $\mathsf{d}_{S}$ for a two-point space $X=\{x_{1},x_{2}\}$ where probability measures are fully specified by the mass $m\in[0,1]$ located at $x_{1}$ and the metric is fully characterized by the distance $d(x_{1},x_{2})$. The precise result is stated in Proposition 6.3 below. When $d(x_{1},x_{2})$ is much smaller than the blur scale $\sqrt{\varepsilon}$ we find that $\mathsf{d}_{S}(\delta_{x_{1}},\delta_{x_{2}})\approx d(x_{1},x_{2})$, which is the same as one would get on an interval $[x_{1},x_{2}]\subset\mathbb{R}$ of length $d(x_{1},x_{2})$. Thus, it seems that the metric $\mathsf{d}_{S}$ may ‘gloss over’ the discreteness of $X$ at very small length scales. Conversely, for $d(x_{1},x_{2})\gg\sqrt{\varepsilon}$ the diameter of the space $(\mathcal{P}(X),\mathsf{d}_{S})$ grows at least like $\exp(d(x_{1},x_{2})^{2}/(2\varepsilon))$. As $d(x_{1},x_{2})$ increases, the blur in the entropic optimal transport underlying the construction of $\mathsf{d}_{S}$ decreases. It is well known that $\mathcal{P}(X)$ equipped with the quadratic Monge–Kantorovich distance is not a length space [43, Remark 2.1]. Thus this exponential divergence of the diameter is consistent with the conjecture that $\mathsf{d}_{S}$ converges to the unregularized quadratic Monge–Kantorovich distance as $\varepsilon\to 0$. See Section 8 for a slightly longer discussion and some related references.
Proposition 6.3.
Let $X=\{x_{1},x_{2}\}$. For a parameter $r\in(0,\infty)$ let $d(x_{1},x_{2})=r$, $c=d^{2}$.
  1. 1.
    For $m\in(0,1)$, $\mu=m\delta_{x_{1}}+(1-m)\delta_{x_{2}}$, $\dot{m}\in\mathbb{R}$, $\dot{\mu}=\dot{m}(\delta_{x_{1}}-\delta_{x_{2}})$, denote by $p$ the amount of mass moving from $x_{1}$ to $x_{2}$ in the optimal self-transport plan from $\mu$ to itself. Then the metric tensor is given by
    $\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})=\frac{\varepsilon}{2}\dot{m}^{2}\frac{m(1-m)-p}{p(2m(1-m)-p)}.$
  2. 2.
    For $m\colon[0,1]\to[0,1]$ continuously differentiable set $\mu_{t}=m(t)\delta_{x_{1}}+(1-m(t))\delta_{x_{2}}$ for $t\in[0,1]$. Then, as $r\to 0$,
    $E((\mu_{t})_{t})=r^{2}\cdot\int_{0}^{1}\dot{m}_{t}^{2}\,\mathrm{d}t+\mathcal{O}\left(\frac{1}{\varepsilon}\lVert\dot{m}\rVert_{\infty}^{2}r^{4}\right).$ (41)
  3. 3.
    Let $C>0$. In the setting of point 1, when $\sqrt{m(1-m)}\geq C\cdot\exp(-2r^{2}/\varepsilon)$ it holds
    $\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})=\frac{\varepsilon}{2}\dot{m}^{2}\cdot\left(\frac{\exp(r^{2}/\varepsilon)}{2\sqrt{m(1-m)}}+\mathcal{O}((m(1-m))^{-1}))\right)$ (42)
    with the constant in the $\mathcal{O}$-term only depending on $C$.
In particular, for small distances, $\dot{m}=-1$ in (41) leads to $E(((1-t)\delta_{x_{1}}+t\delta_{x_{2}})_{t})=r^{2}+\mathcal{O}\left(\frac{1}{\varepsilon}r^{4}\right)$.
Proof.
Let $m\in(0,1)$, $\mu=m\delta_{x_{1}}+(1-m)\delta_{x_{2}}$, $M=\sqrt{m(1-m)}$. We compute the optimal self-transport plan $\pi$ and obtain $k_{\mu}$ as the density of $\pi$ with respect to $\mu\otimes\mu$. Since $X$ only contains two points, we may use the matrix vector notation of the discrete setting. To avoid ambiguity, we denote the entry-wise multiplication of two matrices $A,B$ by $A\odot B$. Denoting $a_{\mu}=\exp({f_{\mu,\mu}}/\varepsilon)$ as a vector $a=(a_{1},a_{2})^{\mathrm{T}}$, we see that $\pi=k_{\mu}\odot\mu\otimes\mu=(a\cdot a^{\mathrm{T}})\odot k_{c}\odot\mu\otimes\mu$. The set of admissible transport plans is described entirely by one parameter $p$ describing the exchange of mass between the two points. With $\exp(-2d(x_{1},x_{2})^{2}/\varepsilon)=:b\in(0,1)$,
$\pi=\begin{pmatrix}m-p&p\\ p&(1-m)-p\end{pmatrix},\qquad k_{c}\odot\mu\otimes\mu=\begin{pmatrix}m^{2}&M^{2}\sqrt{b}\\ M^{2}\sqrt{b}&(1-m)^{2}\end{pmatrix}.$
We have $p^{2}/((m-p)(1-m-p))=(a_{1}a_{2}M^{2}\sqrt{b})^{2}/(a_{1}^{2}m^{2}a_{2}^{2}(1-m)^{2})=b.$ This is a quadratic equation whose solution is
$p=\frac{1}{2(1-b)}\left[-b+\sqrt{b^{2}+4M^{2}b(1-b)}\right]$ (43)
We can express the matrix representing $K_{\mu}$ as a scaling of $\pi$:
$K_{\mu}$ $=\begin{pmatrix}m^{-1}&0\\ 0&(1-m)^{-1}\end{pmatrix}\cdot\pi=\begin{pmatrix}\frac{m-p}{m}&\frac{p}{m}\\ \frac{p}{1-m}&\frac{1-m-p}{1-m}\end{pmatrix}.$
Without suprise $(1,1)^{\mathrm{T}}$ is an eigenvector associated to the eigenvalue $\lambda_{1}=1$. We know that the other eigenvector of $K_{\mu}$ should be orthogonal to $(1,1)^{\mathrm{T}}$ in the inner product of $L^{2}(X,\mu)$, so it must be $(1-m,-m)^{\mathrm{T}}$. The corresponding eigenvalue is computed via
$K_{\mu}\begin{pmatrix}1-m\\ -m\end{pmatrix}=(1-{p}/{M^{2}})\begin{pmatrix}1-m\\ -m\end{pmatrix}.$ (44)
In particular, on this eigenspace $(\operatorname{id}-K_{\mu}^{2})^{-1}$ acts as multiplication by $(1-\lambda_{2}^{2})^{-1}$ with $\lambda_{2}=1-p/M^{2}$. To compute the metric tensor $\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})=\langle\dot{\mu},(\operatorname{id}-K_{\mu}^{2})^{-1}H_{\mu}[\dot{\mu}]\rangle$ for some vertical perturbation $\dot{\mu}=\dot{m}\cdot(1,-1)^{\mathrm{T}}$, we need to compute $H_{\mu}[\dot{\mu}]=K_{\mu}\left[{\mathrm{d}\dot{\mu}}/{\mathrm{d}\mu}\right]$. As ${\mathrm{d}\dot{\mu}}/{\mathrm{d}\mu}=\dot{m}/M^{2}(1-m,-m)^{\mathrm{T}}$, by the previous computation
$H_{\mu}\begin{pmatrix}1\\ -1\end{pmatrix}=\frac{1-{p}/{M^{2}}}{M^{2}}\begin{pmatrix}1-m\\ -m\end{pmatrix}.$ (45)
The formulas (44) and (45) lead to
$\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})=\frac{\varepsilon}{2}\frac{\dot{m}^{2}\lambda_{2}}{M^{2}(1-\lambda_{2}^{2})}=\frac{\varepsilon}{2}\dot{m}^{2}\frac{M^{2}-p}{p(2M^{2}-p)}.$ (46)
Small $r$ asymptotic: The small parameter in the expansion (43) is $\tau=2d(x_{1},x_{2})^{2}/\varepsilon$, so that we have $(1-b)=\tau+\mathcal{O}(\tau^{2})$. A second-order Taylor expansion of the square root leads to $p=M^{2}-\tau M^{4}+\mathcal{O}(\tau^{2})$, which gives $\lambda_{2}=\tau M^{2}+\mathcal{O}(\tau^{2})$.
Thus $(1-\lambda_{2}^{2})^{-1}=1+\mathcal{O}(\tau^{2})$ and (46) gives
$\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})=\frac{\varepsilon}{2}(\tau\dot{m}^{2}+\mathcal{O}(\tau^{2}\dot{m}^{2}))=\dot{m}^{2}\cdot d(x_{1},x_{2})^{2}+\mathcal{O}\left(\frac{1}{\varepsilon}\dot{m}^{2}d(x_{1},x_{2})^{4}\right).$
Big $r$ asymptotic: We set $\varrho:=\sqrt{b}=\exp(-d(x_{1},x_{2})^{2}/\varepsilon)$. Provided that $M\geq C\varrho$ for some $C>0$, rewriting (43) gives
$p=\frac{1}{2(1-\varrho^{2})}\left(-\varrho^{2}+2M\varrho\sqrt{1+\varrho^{2}(1-4M^{2})/(4M^{2})}\right)=M\varrho+\mathcal{O}(\varrho^{2})$
with the constant in the $\mathcal{O}$-term only depending on $C$. This gives
$\lambda_{2}=1-\frac{p}{M^{2}}=1-\frac{\varrho}{M}+\mathcal{O}((\varrho/M)^{2}$
and using (46) it leads to
$\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})$ $=\frac{\varepsilon}{2}\dot{m}^{2}\cdot\left(\frac{1}{2M\varrho}+\mathcal{O}(1/M^{2})\right)=\frac{\varepsilon}{2}\dot{m}^{2}\cdot\left(\frac{\exp(r^{2}/\varepsilon)}{2M}+\mathcal{O}(1/M^{2}))\right).\qed$
For small $M/\varrho$ one can use analogous methods to find $p=M^{2}+\mathcal{O}(M^{4}/\varrho^{2})$, which leads to $\lambda_{2}=\mathcal{O}(1/\varrho^{2})$ and $\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})=\mathcal{O}(\varepsilon\dot{m}^{2}/\varrho^{2})$.
7 Negative results on the Sinkhorn divergence and the metric tensor
7.1 $\sqrt{S_{\varepsilon}}$ violates the triangle inequality
One of the main goals of this paper is to find a “Riemannian” distance associated with the Sinkhorn divergence $S_{\varepsilon}$. The reader may wonder why one could not directly use $S_{\varepsilon}$ as a distance, since it enjoys several nice features, as recalled in Theorem 1.1. However, despite being positive definite, already for the squared Euclidean distance as the cost function the Sinkhorn divergence, or rather $\sqrt{S_{\varepsilon}}$ (since $S_{\varepsilon}$ converges to the square of the Monge-Kantorovich distance), is not itself a distance! Although this fact is well known in the community, we have not been able to find an explicit statement of this property. For this reason, we provide an elementary proof, which will actually exclude any convex power of $\sqrt{S_{\varepsilon}}$.
Theorem 7.1.
On $\mathbb{R}$ let $c(x,y)=\lvert x-y\rvert^{2}$. Then $S_{\varepsilon}^{\alpha}$ fails to satisfy the triangle inequality for any $\alpha\geq\frac{1}{2}$.
The case $0<\alpha<\frac{1}{2}$ is unsuitable for our purposes too, as $S_{\varepsilon}^{\alpha}(\delta_{x},\delta_{y})=\lVert x-y\rVert^{2\alpha}$ is a concave function of the Euclidean distance. Hence if $S_{\varepsilon}^{\alpha}$ were a metric at all, it would induce a Finslerian rather than Riemannian geometry.
Proof.
The idea is to look at the Sinkhorn divergence between 1-dimensional Gaussian distributions, for which an explicit formula is available as a direct consequence of (39) (see [44, Cor. 1]). Indeed, given $\mu_{i}=\mathcal{N}(m_{i},v_{i})$ with $m_{i}\in\mathbb{R}$, $v_{i}\geq 0$ (for $i=0,1$), the Sinkhorn divergence between them is explicitly given by
$S_{\varepsilon}(\mu_{0},\mu_{1})=\|m_{0}-m_{1}\|^{2}+\frac{\varepsilon}{4}\bigg{(}\kappa_{00}-2\kappa_{01}+\kappa_{11}+\log\Big{(}\frac{\kappa_{01}^{2}}{\kappa_{00}\kappa_{11}}\Big{)}\bigg{)},$ (47)
where $\kappa_{ij}:=1+(1+16\varepsilon^{-2}v_{i}v_{j})^{1/2}$, for $i,j\in\{0,1\}$, as in (39).
It is readily verified that $S_{\varepsilon}^{\alpha}$ cannot be a metric for any $\alpha>1/2$, as it is sufficient to consider $\mu_{i}=\delta_{m_{i}}$ and observe that $S_{\varepsilon}^{\alpha}(\mu_{0},\mu_{1})=\|m_{0}-m_{1}\|^{2\alpha}$ does not satisfy the triangle inequality.
Finally, $\sqrt{S_{\varepsilon}}$ is not a distance either. To see why this is the case, choose $\mu_{0}=\delta_{0}$, $\mu_{1}=\mathcal{N}(0,\frac{\varepsilon}{4}x)$ and $\mu_{2}=\mathcal{N}(0,\frac{\varepsilon}{4}v)$, for a fixed $v>0$: we claim that
$\sqrt{S_{\varepsilon}(\mu_{0},\mu_{2})}>\sqrt{S_{\varepsilon}(\mu_{0},\mu_{1})}+\sqrt{S_{\varepsilon}(\mu_{1},\mu_{2})}$ (48)
when $x$ is sufficiently small and $v$ sufficiently large. To prove the claim, observe that
$14\varepsilon^{-1}S_{\varepsilon}(\mu_{0},\mu_{2})=\sqrt{1+v^{2}}-\log(1+\sqrt{1+v^{2}})+\log 2-$ 1 .
To lighten the notation, let us introduce
$\begin{split}\zeta(x):=4\varepsilon^{-1}S_{\varepsilon}(\mu_{0},\mu_{1})&=\sqrt{1+x^{2}}-\log(1+\sqrt{1+x^{2}})+\log 2-1\\ \xi(x):=4\varepsilon^{-1}S_{\varepsilon}(\mu_{1},\mu_{2})&=\sqrt{1+v^{2}}-2\sqrt{1+vx}+2\log(1+\sqrt{1+vx})\\ &\qquad-\log(1+\sqrt{1+v^{2}})-\log 2+1+\zeta(x).\end{split}$
Squaring both sides, and after algebraic manipulations, (48) is thus equivalent to
$\Psi(x):=2\log(1+\sqrt{1+vx})-2\sqrt{1+vx}-2\log 2+2+2\zeta(x)+2\sqrt{\zeta(x)\xi(x)}<0$
for $x$ sufficiently close to 0. To prove this inequality, let us compute
$\begin{split}\Psi^{\prime}(x)=&-\frac{v}{1+\sqrt{1+vx}}+\frac{2x}{1+\sqrt{1+x^{2}}}+\left(\frac{\xi(x)}{\zeta(x)}\right)^{1/2}\frac{x}{1+\sqrt{1+x^{2}}}\\ &+\left(\frac{\zeta(x)}{\xi(x)}\right)^{1/2}\left(-\frac{v}{1+\sqrt{1+vx}}+\frac{x}{1+\sqrt{1+x^{2}}}\right)\end{split}$
and note that $\zeta(0)=0$, so that $\Psi(0)=0$ and
$\lim_{x\to 0}\Psi^{\prime}(x)=-\frac{v}{2}+\frac{\sqrt{\xi(0)}}{2}\lim_{x\to 0}\frac{x}{\sqrt{\zeta(x)}}$
with $\xi(0)=\sqrt{1+v^{2}}-1+\log 2-\log(1+\sqrt{1+v^{2}})>0$. It is now sufficient to observe that $\zeta^{\prime}(0)=0$ and $\zeta^{\prime\prime}(0)=1/2$: by a Taylor expansion of $\zeta(x)$ up to order 2 we finally get
$\Psi^{\prime}(0)=-\frac{v}{2}+(\sqrt{1+v^{2}}-1+\log 2-\log(1+\sqrt{1+v^{2}}))^{1/2},$
which is negative provided $v$ is sufficiently large. We have thus shown that $\Psi(0)=0$ and $\Psi^{\prime}(0)<0$, hence the conclusion. ∎
Remark 7.2.
One can also use the measures $\mu_{r}=\frac{1}{2}\delta_{r}+\frac{1}{2}\delta_{-r}$ from Section 7.2 to show that the square root of the Sinkhorn divergence does not satisfy the triangle inequality: Choose $r=\sqrt{\varepsilon}$. Then Lemma 7.5 gives
$\frac{1}{\varepsilon}S_{\varepsilon}(\mu_{0},\mu_{r})=1-\frac{1}{2}\log 2+\frac{1}{2}\log\left(1+\frac{1}{e^{4}}\right),$
$\frac{1}{\varepsilon}S_{\varepsilon}(\mu_{r},\mu_{2r})=1-\log\left(1+\frac{1}{e^{8}}\right)+\frac{1}{2}\log\left(1+\frac{1}{e^{4}}\right)+\frac{1}{2}\log\left(1+\frac{1}{e^{16}}\right),$
$\frac{1}{\varepsilon}S_{\varepsilon}(\mu_{0},\mu_{2r})=4-\frac{1}{2}\log 2+\frac{1}{2}\log\left(1+\frac{1}{e^{16}}\right).$
This yields $\frac{1}{\sqrt{\varepsilon}}\cdot\left(\sqrt{S_{\varepsilon}(\mu_{0},\mu_{2r})}-\sqrt{S_{\varepsilon}(\mu_{0},\mu_{r})}-\sqrt{S_{\varepsilon}(\mu_{r},\mu_{2r})}\right)\approx 0.093>0$, which breaks the triangle inequality.
Figure 2: Gap in the triangle inequality for $\sqrt{S_{\varepsilon}}$ on the measures $\mu_{r}=\frac{1}{2}\delta_{r}+\frac{1}{2}\delta_{-r}$, discussed in Remark 7.2, for $\varepsilon=1$. Left: The gap in the triangle inequality when going from $\mu_{s}$ to $\mu_{r}$ via $\mu_{1}$. The triangle inequality breaks in the blue region, where $s\ll 1=\sqrt{\varepsilon}\ll r$. Right: Gap for $s=0$ and varying intermediate position $0\leq t\leq r$ for different values of $r$.
Remark 7.3.
We observe that in the proof of Theorem 7.1 and in Figure 2, the triangle inequality for $\sqrt{S_{\varepsilon}}$ breaks when the initial and intermediate measures are concentrated below the blur scale and the target measure is more spread out.
7.2 The Sinkhorn divergence is not jointly convex
The second negative result about the Sinkhorn divergence deals with joint convexity. The Sinkhorn divergence $S_{\varepsilon}$ interpolates between squared MMD and Wasserstein distances which are both jointly convex functions of their inputs, and moreover $S_{\varepsilon}$ is coordinate-wise convex as recalled in Theorem 1.1. Thus it is natural to investigate the convexity properties of $S_{\varepsilon}$ as a function on $\mathcal{P}(X)\times\mathcal{P}(X)$, and we show that the answer is negative.
Throughout this section let $c(x,y)=\lvert x-y\rvert^{2}$ be the quadratic cost.
Theorem 7.4.
Let $X\subseteq\mathbb{R}$ be the closure of an open set with $\operatorname{diam}(X)>\sqrt{2\varepsilon}$. Then the Sinkhorn divergence $S_{\varepsilon}$ is not jointly convex over $\mathcal{P}(X)\times\mathcal{P}(X)$.
We first examine the measures that will form the counterexample.
Lemma 7.5.
For $r\in[0,\infty)$ set $\mu_{r}:=\frac{1}{2}\delta_{r}+\frac{1}{2}\delta_{-r}$. Then for $r,s\in[0,\infty)$ it holds
$\frac{1}{\varepsilon}\mathrm{OT}_{\varepsilon}(\mu_{r},\mu_{s})=\log 2-\log(k_{c}(r,s)+k_{c}(r,-s)),$
$\frac{1}{\varepsilon}S_{\varepsilon}(\mu_{r},\mu_{s})=-\log(k_{c}(r,s)+k_{c}(r,-s))+\frac{1}{2}\log(1+k_{c}(r,-r))+\frac{1}{2}\log(1+k_{c}(s,-s)).$
Proof.
We first compute the Schrödinger potentials for $(\mu_{r},\mu_{s})$. As $\mu_{r}$ and $\mu_{s}$ are symmetric around 00, we have $f_{\mu_{r},\mu_{s}}(r)=f_{\mu_{r},\mu_{s}}(-r)=:f(r)$, $g_{\mu_{r},\mu_{s}}(s)=g_{\mu_{r},\mu_{s}}(-s)=:g(s)$. The Schrödinger system (13) then reduces to
$\exp(-f/\varepsilon)=\frac{1}{2}k_{c}(r,s)\exp(g/\varepsilon)+\frac{1}{2}k_{c}(r,-s)\exp(g/\varepsilon),$
i.e. $\frac{1}{\varepsilon}(f\oplus g)=\log 2-\log(k_{c}(r,s)+\log(k_{c}(r,-s))$. Now the claim follows from (12). ∎
With these explicit expressions at hand, the proof of Theorem 7.4 boils down to an appropriate choice of parameters.
Proof of Theorem 7.4.
Use $\mu_{r}:=\frac{1}{2}\delta_{r}+\frac{1}{2}\delta_{-r}$ as in Lemma 7.5. For lighter notation, we assume that ±$\pm r\in X$ and we set $\kappa=k_{c}(r,-r)$. We compute the Taylor expansion of $S_{\varepsilon}(\mu_{r},\mu_{r+t})$ for $r\in[0,\infty)$. With the explicit formula for the Sinkhorn divergence from Lemma 7.5 and the second-order Taylor expansions
$\begin{split}&k_{c}(r,r+x)=1-\frac{1}{\varepsilon}x^{2}+\mathcal{O}(x^{3}),\quad k_{c}(r,-(r+x))=\kappa-\kappa\frac{4r}{\varepsilon}x-\kappa\left(\frac{1}{\varepsilon}-\frac{8r^{2}}{\varepsilon^{2}}\right)x^{2}+\mathcal{O}(x^{3}),\\ &\log(1+\kappa+x)=\log(1+\kappa)+\frac{1}{1+\kappa}x-\frac{1}{2(1+\kappa)^{2}}x^{2}+\mathcal{O}(x^{3})\end{split}$
we obtain
$\begin{split}\frac{1}{\varepsilon}S_{\varepsilon}(\mu_{r},\mu_{r+t})&=-\log\left(1+\kappa-\frac{1}{\varepsilon}t^{2}-\kappa\left(\frac{4r}{\varepsilon}t+\left(\frac{1}{\varepsilon}-\frac{8r^{2}}{\varepsilon^{2}}\right)t^{2}\right)+\mathcal{O}(t^{3})\right)+\frac{1}{2}\log(1+\kappa)\\ &\quad+\frac{1}{2}\log\left(1+\kappa-\kappa\left(\frac{4r}{\varepsilon}\cdot 2t+\left(\frac{1}{\varepsilon}-\frac{8r^{2}}{\varepsilon^{2}}\right)\cdot 4t^{2}\right)+\mathcal{O}(t^{3})\right)\\ &=\frac{1}{\varepsilon}t^{2}+t^{2}\frac{\kappa}{1+\kappa}\left(\frac{1}{1+\kappa}\frac{8r^{2}}{\varepsilon^{2}}-\frac{2}{\varepsilon}\right)+\mathcal{O}(t^{3}).\end{split}$ (49)
As $\kappa\leq 1$, when $r>\sqrt{\varepsilon/2}$ this is greater than $\frac{1}{\varepsilon}S_{\varepsilon}(\delta_{\pm r},\delta_{\pm(r+t)})=\frac{1}{\varepsilon}t^{2}$ for small $t$. Thus in this case $S_{\varepsilon}(\mu_{r},\mu_{r+t})>\frac{1}{2}S_{\varepsilon}(\delta_{r},\delta_{r+t})+\frac{1}{2}S_{\varepsilon}(\delta_{-r},\delta_{-(r+t)})$, disproving convexity.
When $\operatorname{diam}(X)>\sqrt{2\varepsilon}$ we can always translate $X$ to contain points ±$\pm r$ with $r>\sqrt{\varepsilon/2}$, hence disproving joint convexity of $S_{\varepsilon}$. ∎
The non-convexity of $S_{\varepsilon}$ descends to the level of the metric tensor.
Corollary 7.6.
Let $X\subseteq\mathbb{R}$ be the closure of an open set with $\operatorname{diam}(X)>\sqrt{2\varepsilon}$. Then the map $(\mu,\dot{\mu})\mapsto\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})$ is not jointly convex on $\left\{(\mu,\dot{\mu})\in\mathcal{P}(X)\times\mathcal{C}^{1}(X)^{*}\,\middle|\,\langle\dot{\mu},\mathbb{1}_{X}\rangle=0\right\}$.
Proof.
We follow the proof of Theorem 7.4. As discussed in Remark 3.14, Theorem 3.3 also applies to the path given by $(\mu_{r+t})_{t}$. Denote $\mu=\mu_{r}$, $\dot{\mu}=\frac{1}{2}\delta_{r}^{\prime}-\frac{1}{2}\delta_{-r}^{\prime}$. Theorem 3.3 and (49) give
$\mathbf{g}_{\mu}(\dot{\mu},\dot{\mu})=\lim_{t\to 0}\frac{S_{\varepsilon}(\mu_{r},\mu_{r+t})}{t^{2}}=1+\frac{\kappa}{1+\kappa}\left(\frac{1}{1+\kappa}\frac{8r^{2}}{\varepsilon}-2\right),$ (50)
which is greater than $\mathbf{g}_{\delta_{\pm r}}(\pm\delta_{\pm r}^{\prime},\pm\delta_{\pm r}^{\prime})=1$ when $r>\sqrt{2\varepsilon}$. ∎
As discussed in the introduction, if the metric tensor were jointly convex, then existence of geodesic (Theorem 5.3) could be obtained with arguments of convex analysis. Corollary 7.6 shows that some more involved arguments as given in Section 5.2 are necessary. Joint convexity of the metric tensor $\tilde{\mathbf{g}}_{\mu}(\dot{\beta},\dot{\beta})$ in $\mu$ and $\dot{\beta}$ is ill defined, because the compatibility assumption $\langle B(\mu),\dot{\beta}\rangle=0$ can fail at a convex combination.
Figure 3: Numerically computed values of the metric tensor of two point masses moving apart with $\mu=m\delta_{-r}+(1-m)\delta_{r}$, $\dot{\mu}=-m\delta_{-r}^{\prime}+(1-m)\delta_{r}^{\prime}$ for different values of $\varepsilon$. This is the linear interpolation between a point mass at $-r$ moving to the left and a point mass at $r$ moving to the right ($m=1$ and $m=0$ respectively). Left: fixed $r=1$, varying mass $m$. Joint convexity of the metric tensor fails, as the graphs sometimes exceed the value one. Right: fixed $m=\frac{1}{2}$, varying position $r$. This is the term (50) computed in the proof of Corollary 7.6, which only depends on ${r}/{\sqrt{\varepsilon}}$. In both plots, the black line shows $\lVert v\rVert^{2}_{L^{2}(X,\mu)}=1$ from the dynamical formulation of Wasserstein-2 optimal transport.
Remark 7.7.
Although very explicit, the previous computations do not shed light on the reason why $S_{\varepsilon}$ fails to be jointly convex. Instead, we present also a more abstract argument, which is originally how we proved Theorem 7.4. It proceeds by contradiction, by combining three results already present in the literature:
  1. (a)
    A very insightful remark of Savaré and Sodini [59], which says that $\mathrm{OT}:=\mathrm{OT}_{0}$ (i.e. the unregularized transport cost induced by $c$ from (1)) is the largest convex and weak-$\ast$-lower semi-continuous function which coincides with $c$ on Dirac masses. Since the set of couplings between two Dirac masses contains only one element, with $c(x,y)=|x-y|^{2}$ we have $S_{\varepsilon}(\delta_{x},\delta_{y})=c(x,y)$ for any $x,y\in X$. As a consequence, if $S_{\varepsilon}$ were jointly convex, as it is in addition weak-$\ast$-continuous, then $S_{\varepsilon}(\mu_{0},\mu_{1})\leq\mathrm{OT}(\mu_{0},\mu_{1})$ for all $\mu_{0},\mu_{1}\in\mathcal{P}_{2}(\mathbb{R})$. So rather than negating the convexity of $S_{\varepsilon}$, it is enough to find measures $\mu_{0},\mu_{1}$ such that $S_{\varepsilon}(\mu_{0},\mu_{1})>\mathrm{OT}(\mu_{0},\mu_{1})$.
  2. (b)
    We then use a small-$\varepsilon$ asymptotics of $S_{\varepsilon}$ [20, 19]. Introducing the Fisher information of a measure $\mu$ w.r.t. the Lebesgue measure as
    $I(\mu)\coloneqq 4\int_{\mathbb{R}}\left|\nabla\sqrt{\rho}\right|^{2}\,\mathrm{d}x,\qquad\textrm{if }\mu=\rho\,\mathrm{d}x\textrm{ and }\sqrt{\rho}\in H^{1}(\mathbb{R},\,\mathrm{d}x),$
    and +$\infty$+ ∞ otherwise, this result reads as follows: Let $\mu_{0},\mu_{1}\in\mathcal{P}(\mathbb{R})$ have bounded densities and supports and denote by $(\mu_{t})_{t\in[0,1]}$ the Wasserstein-2 geodesic connecting them. If $I(\mu_{0})$, $I(\mu_{1})$, and $\int_{0}^{1}I(\mu_{t})\,\mathrm{d}t$ are finite, then it holds
    $S_{\varepsilon}(\mu_{0},\mu_{1})-\mathrm{OT}(\mu_{0},\mu_{1})=\frac{\varepsilon^{2}}{8}\left(\int_{0}^{1}I(\mu_{t})\,\mathrm{d}t-\frac{I(\mu_{0})+I(\mu_{1})}{2}\right)+o(\varepsilon^{2}).$ (51)
    Thus, given (a), it is enough to find measures $\mu_{0}$, $\mu_{1}$ such that
    $\int_{0}^{1}I(\mu_{t})\,\mathrm{d}t>\frac{I(\mu_{0})+I(\mu_{1})}{2},$ (52)
    and this will disprove joint convexity of $S_{\varepsilon}$ for $\varepsilon$ small enough.
  3. (c)
    Eventually, we use the property that the Fisher information is not geodesically convex [17]. Specifically, by [17, Ex. 2 in Sec. 3.2] there exists a Wasserstein geodesic $(\tilde{\mu}_{t})_{t\in[0,1]}$ on $\mathbb{R}$ such that the function $t\mapsto I(\tilde{\mu}_{t})$ is not convex. This implies that we can find $t_{0},t_{1}$ such that
    $\frac{1}{t_{1}-t_{0}}\int_{t_{0}}^{t_{1}}I(\tilde{\mu}_{t})\,\mathrm{d}t>\frac{I(\tilde{\mu}_{t_{0}})+I(\tilde{\mu}_{t_{1}})}{2},$
    because if the equality were in the other direction for any $t_{0},t_{1}$, it would imply convexity of $t\mapsto I(\tilde{\mu}_{t})$. Defining $\mu_{t}:=\tilde{\mu}_{t_{0}+t(t_{1}-t_{0})}$ for $t\in[0,1]$, we obtain (52) and conclude.
A scaling argument then extends the lack of convexity to any $\varepsilon>0$. Indeed, it suffices to define $L:\mathbb{R}\to\mathbb{R}$, $L_{\varepsilon}(x)\coloneqq\varepsilon^{-1/2}x$ and observe that $S_{\varepsilon}(\mu,\nu)=\varepsilon S_{1}(L_{\#}\mu,L_{\#}\nu)$. This can be checked with the Schrödinger system (13), see also [46]. So that non-convexity of $S_{\varepsilon}$ for one $\varepsilon>0$ is equivalent to non-convexity for any $\varepsilon>0$.
At the technical level, however, there is a slight gap in the argument we have just outlined, since (51) requires $\mu_{0},\mu_{1}$ to have bounded support, while the counterexample built by Carrillo and Slepčev [17, Ex. 2 in Sec. 3.2] has unbounded support. This difficulty can be overcome by a standard truncation procedure, but the resulting argument would be much heavier to read than the straightforward computation of Proposition 7.5.
8 Outlook
In this article we have introduced a Riemannian distance associated with entropic optimal transport and established its most fundamental properties, such as its connection to RKHS spaces, the existence of geodesics, its metrization of the weak-$\ast$ topology, as well as some instructive examples, such as translations, Gaussians, and the two-point space. This Riemannian perspective provides further theoretical underpinning for the widespread application of the Sinkhorn divergence in machine learning. After this foundation has been established, there are various more advanced intriguing directions for further study that we comment now.
Convergence when $\varepsilon\to 0$. One immediate question is, as one sends $\varepsilon$ to 00, whether the limit of $\mathsf{d}_{S}$ is the quadratic Monge–Kantorovich distance, and if there is a corresponding metric convergence in the sense of Gromov–Hausdorff.
Homogeneization when the metric space $X$ is refined. The new Sinkhorn distance is a geodesic distance on any compact metric space $(X,d)$ when $c=d^{2}$ induces a universal kernel, in particular also in cases where $(X,d)$ is not a length space itself, such as the two-point space. This is in stark contrast to the quadratic Monge–Kantorovich distance, for which there are no geodesics when $X$ is discrete. A distance on probability measures over discrete graphs that resembles the Monge–Kantorovich distance has been introduced in [43] with non-trivial homogenization results given, for instance, in [32, 33]. For the distance $\mathsf{d}_{S}$ for measures on the two-point space, a consequence of Proposition 6.3 is that $\mathsf{d}_{S}(\delta_{x_{1}},\delta_{x_{2}})\approx d(x_{1},x_{2})$ when $d(x_{1},x_{2})\ll\sqrt{\varepsilon}$. Based on this one may wonder, whether the metric space $(\mathcal{P}(X),\mathsf{d}_{S})$ can be approximated by $(\mathcal{P}(X_{n}),\mathsf{d}_{S})$ where $(X_{n})_{n\in\mathbb{N}}$ is a sequence of increasingly fine discrete approximations of $X$ with respect to the ground metric $d$. In other words, does $\mathsf{d}_{S}$ over $\mathcal{P}(X)$ arise as the homogenization of itself? In combination with the previous question one might even ask, whether it is possible to decrease $\varepsilon$ to zero at a suitable rate, while $X_{n}$ is refined, such that the joint limit becomes the quadratic Monge–Kantorovich distance over $\mathcal{P}(X)$.
Numerical approximation. A more practical concern is the efficient numerical approximation of $\mathsf{d}_{S}$ and its geodesics. As the Riemannian tensor of $\mathsf{d}_{S}$ was chosen to be the Hessian of the Sinkhorn divergence $S_{\varepsilon}$, one might conjecture that
$\mathsf{d}_{S}(\mu,\nu)^{2}=\lim_{N\to\infty}\inf\left\{N\cdot\sum_{n=0}^{N-1}S_{\varepsilon}(\rho_{n},\rho_{n+1})\middle|\rho_{0},\ldots,\rho_{N}\in\mathcal{P}(X),\,\rho_{0}=\mu,\,\rho_{N}=\nu\right\}$
in the spirit of time-discrete variational geodesic calculus [56]. For this it seems that one requires a more uniform version of the expansion of Theorem 3.3.
Sample complexity. From a statistical perspective an interesting question is whether $\mathsf{d}_{S}$ inherits the favourable sample complexity of the Sinkhorn divergence [28, 46].
Linearization. For the quadratic Monge–Kantorovich distance local approximation via linearization has been established as a useful tool for data analysis applications [63] which reduces the computational complexity and allows to combine the geometry of optimal transport with standard data analysis tools on Hilbert spaces. This might also be of great practical relevance for the new metric $\mathsf{d}_{S}$ and thus one should look for a good approximation of the corresponding logarithmic map.
Gradient flows. Gradient flow with respect to the Monge-Kantorovich distance is a well-studied topic with important consequences at the theoretical and practical level. It would be important to understand how the gradient flows with respect to the metric $\mathsf{d}_{S}$ behave. They should be recovered as a scaling limit if ones tries to use the minimizing movemement scheme for gradient flows, but with the Sinkhorn divergence $S_{\varepsilon}$ in place of the more classical squared Monge-Kantorovich distance, for a fixed $\varepsilon>0$.
Of course this list is far from complete and in particular more questions will emerge over time, as our understanding of the new distance deepens.
Acknowledgements
The research leading to this work started during a visit of BS to HL and LT in the Department of Decision Sciences of Bocconi University, which is warmly acknowledged for the hospitality. HL and LT acknowledge the support of the MUR-Prin 2022-202244A7YL “Gradient Flows and Non-Smooth Geometric Structures with Applications to Optimization and Machine Learning”, funded by the European Union - Next Generation EU. LT is a also member of INdAM and the GNAMPA group. GM gratefully acknowledges funding of the DFG within CRC1456-A04. BS gratefully acknowledges funding of the DFG within the Emmy Noether programme and in project CRC1456-A03.
Appendix A Some results of functional analysis
Lemma A.1.
Let $Y$ be a Banach space, X a compact topological space. For $n\in\mathbb{N}$ let $\omega_{n}\in Y^{*}$ with $\omega_{n}\to\omega\in Y^{*}$ in the weak-$\ast$ topology. For $x\in X$ let $k_{x}\in Y$, such that $x\mapsto k_{x}$ is continuous with respect to the norm-topology on $Y$. Then
$\sup_{x\in X}\lvert\langle\omega_{n},k_{x}\rangle-\langle\omega,k_{x}\rangle\rvert\to 0,$
so $x\mapsto\langle\omega_{n},k_{x}\rangle$ converges uniformly on $X$ as $n\to\infty$.
We usually apply this with $k_{x}=k(x,\cdot)$ for some suitable $k\colon X\times X\to\mathbb{R}$ and $Y=\mathcal{M}(X)$.
Proof.
Let $\phi_{n}(x)=\langle\omega_{n},k_{x}\rangle$, $\phi(x)=\langle\omega,k_{x}\rangle$. By assumption $\phi,\phi_{n}\in\mathcal{C}(X)$, $\phi_{n}\to\phi$ pointwise. The functions $\phi_{n}$ are equi-continuous, as for any $n\in\mathbb{N}$
$\lvert\phi_{n}(x)-\phi_{n}(y)\rvert\leq\lVert\omega_{n}\rVert_{Y^{*}}\cdot\lVert k_{x}-k_{y}\rVert_{Y},$
the weak-$\ast$-convergent sequence $(\omega_{n})_{n}$ is bounded, and the map $x\mapsto k_{x}$ is uniformly continuous. Uniform convergence now follows from the Arzelà–Ascoli Theorem. ∎
Lemma A.2 ([12, Proposition 3.5 (iv)]).
Let $Y$ be a Banach space. For $n\in\mathbb{N}$ let $\omega_{n}\in Y^{*}$ with $\omega_{n}\to\omega\in Y^{*}$ in the weak-$\ast$ topology. Let $y_{n}\in Y$ with $y_{n}\to y\in Y$ in norm. Then
$\lim_{n\to\infty}\langle\omega_{n},y_{n}\rangle=\langle\omega,y\rangle.$
Appendix B Additional results on Reproducing Kernel Hilbert Spaces
Recall the construction of the RKHS $\mathcal{H}_{k}$ corresponding to a universal kernel $k\in\mathcal{C}(X\times X)$ from Section 4.1 on a compact Hausdorff space $X$. The symmetric bilinear form (23) is positive definite, because every universal kernel is already positive definite. A detailed analysis of these two conditions can be found in [60] and [47]. For all $\phi\in\mathcal{H}_{k}$ and all $x\in X$ holds $\phi(x)=\langle\phi,k(x,\cdot)\rangle_{\mathcal{H}_{k}}$. Using the Cauchy–Schwarz inequality we find
$\lVert\phi\rVert_{\infty}=\sup_{x\in X}\lvert\langle\phi,k(x,\cdot)\rangle_{\mathcal{H}_{k}}\rvert\leq\lVert\phi\rVert_{\mathcal{H}_{k}}\cdot\sup_{x\in X}k(x,x)$ (53)
for any $\phi\in\mathcal{H}_{k}$. Hence the injection of $\mathcal{H}_{k}$ into $\mathcal{C}(X)$ is continuous. Even weak convergence in the RKHS is quite a strong concept.
Lemma B.1.
Weak convergence in $\mathcal{H}_{k}$ implies uniform convergence.
Proof.
Let $\phi_{n}\in\mathcal{H}_{k}$ converge weakly to $\phi\in\mathcal{H}_{k}$. Then $\phi_{n}(x)=\langle\phi_{n},k(x,\cdot)\rangle_{\mathcal{H}_{k}}\to\langle\phi,k(x,\cdot)\rangle_{\mathcal{H}_{k}}=\phi(x)$ for all $x\in X$. Uniform convergence follows from Lemma A.1
The embedding $\mathcal{M}(X)\to\mathcal{H}_{k}^{*}$, restricting the integration map of a measure to functions in $\mathcal{H}_{k}$, is injective because $\mathcal{H}_{k}\subseteq\mathcal{C}(X)$ is dense and is a continuous linear operator by (53). By definition the Riesz map $H\colon\mathcal{H}_{k}^{*}\to\mathcal{H}_{k}$ given by (24) is an isometry. For a signed measure $\nu\in\mathcal{M}(X)$ it holds $H_{k}[\nu](x)=\int{k(x,\cdot)}\,\mathrm{d}{\nu}$ for all $x\in X$ and
$\lVert H_{k}[\nu]\rVert_{\mathcal{H}_{k}}^{2}=\langle\nu,H_{k}[\nu]\rangle=\iint_{X\times X}k(x,y)\,\mathrm{d}\nu(x)\,\mathrm{d}\nu(y).$
Lemma B.2.
The operator $H_{k}\colon\mathcal{M}(X)\to\mathcal{H}_{k}$ is weak-$\ast$-to-norm continuous. In particular, $H_{k}\colon\mathcal{M}(X)\to\mathcal{H}_{k}$ is a norm-to-norm compact operator and $H_{k}$ restricted to $\mathcal{P}(X)$ is a weak-$\ast$-to-norm homeomorphism onto its image.
Proof.
We prove weak-$\ast$-to-norm continuity. The other claims follow from the Banach–Alaoglu theorem and the fact that a continuous injection between compact Hausdorff spaces is a homeomorphism onto its image. Let $\nu_{n}\to\nu$ weak-$\ast$ in $\mathcal{M}(X)$. Since $\mathcal{H}_{k}\subseteq\mathcal{C}(X)$, we directly obtain $H_{k}[\nu_{n}]\to H_{k}[\nu]$ weakly in $\mathcal{H}_{k}$, which by Lemma B.1 implies uniform convergence of $H_{k}[\nu_{n}]$ to $H_{k}[\nu]$. Thus Lemma A.2 shows that
$\lVert H_{k}[\nu_{n}]-H_{k}[\nu]\rVert_{\mathcal{H}_{k}}^{2}=\langle\nu_{n}-\nu,H_{k}[\nu_{n}]-H_{k}[\nu]\rangle\to\langle 0,0\rangle=0.$
When $X\subseteq\mathbb{R}^{d}$ is the closure of its interior and $k\in\mathcal{C}^{2}(X\times X)$, Lemma B.2 remains true for the kernel mean embedding of distributions of order one by replacing $\mathcal{M}(X)$ with $\mathcal{C}^{1}(X)^{*}$ everywhere.
Appendix C The Sobolev space of Hilbert space valued functions
We briefly summarize a few results from [37, Sec. 3.2]. Let $\mathcal{H}$ be a Hilbert space. We write $L^{2}((0,1);\mathcal{H})$ for the space of measurable functions $u\colon(0,1)\to\mathcal{H}$ for which $\lVert u(\cdot)\rVert_{\mathcal{H}}$ is 2-integrable, up to equality almost everywhere. It carries the obvious inner product. We say that a function $u\in L^{2}((0,1);\mathcal{H})$ has the distributional derivative $v\in L^{2}((0,1);\mathcal{H})$ and write $v=u^{\prime}$ if for all $\phi\in\mathcal{C}_{\textup{c}}^{\infty}((0,1);\mathbb{R})$ it holds
$\int_{0}^{1}u(t)\phi^{\prime}(t)\,\mathrm{d}t=-\int_{0}^{1}v(t)\phi(t)\,\mathrm{d}t$
as Bochner integrals. The first Sobolev space is given by
$\mathcal{H}^{1}((0,1);\mathcal{H})=\left\{u\colon(0,1)\to\mathcal{H}\,\middle|\,u\textup{ has a distr.~{}derivative }u^{\prime}\textup{ and }u,u^{\prime}\in L^{2}((0,1);\mathcal{H})\right\}.$
It is a Hilbert space with the inner product $\langle u_{1},u_{2}\rangle_{\mathcal{H}^{1}}=\langle u_{1},u_{2}\rangle_{L^{2}}+\langle u_{1}^{\prime},u_{2}^{\prime}\rangle_{L^{2}}$.
Theorem C.1 ([37, Thm. 2.5]).
A measurable function $u\colon(0,1)\to\mathcal{H}$ is Bochner integrable if and only if $\lVert u(\cdot)\rVert_{\mathcal{H}}$ is Lebesgue-integrable. In this case $\int_{0}^{1}u(t)\,\mathrm{d}t\in\mathcal{H}$ and it holds
$\left\lVert\int_{0}^{1}u(t)\,\mathrm{d}t\right\rVert_{\mathcal{H}}\leq\int_{0}^{1}\lVert u(t)\rVert_{\mathcal{H}}\,\mathrm{d}t.$
Theorem C.2 ([37, Thm. 3.7, Prop. 3.8, Thm.3.13]).
Let $u\in L^{2}((0,1);\mathcal{H})$. Then $u\in\mathcal{H}^{1}((0,1);\mathcal{H})$ if and only if one of the following conditions is satisfied:
  1. 1.
    $u\in\mathrm{AC}^{2}((0,1);\mathcal{H})$, i.e. there exists a representation of $u$ and a function $m\in L^{2}((0,1);\mathbb{R})$, such that $\lVert u(t)-u(s)\rVert_{\mathcal{H}}\leq\lvert\int_{s}^{t}m(\tau)\,\mathrm{d}\tau\rvert$ for all $t,s\in(0,1)$.
  2. 2.
    $u$ is differentiable a.e., $u^{\prime}\in L^{2}((0,1);\mathcal{H})$ and there is a $t_{0}\in(0,1)$, such that for almost all $t$ it holds
    $u(t)=u(t_{0})+\int_{t_{0}}^{t}u^{\prime}(\tau)\,\mathrm{d}\tau.$
If $u\in\mathcal{H}^{1}((0,1);\mathcal{H})$ then $m(\tau)=\lVert u^{\prime}(\tau)\rVert_{\mathcal{H}}$ is a possible choice in 1.
As a consequence of absolute continuity, every $u\in\mathcal{H}^{1}((0,1);\mathcal{H})$ continuously extends to the boundary points $\{0,1\}$ of the interval $(0,1)$.